### v1

# Convergent transformations of visual representation in brains and models

Abstract
A fundamental question in cognitive neuroscience is what shapes visual perception: the external world’s structure or the brain’s internal architecture. Although some perceptual variability can be traced to individual differences, brain responses to naturalistic stimuli evoke similar activity patterns across individuals, suggesting a convergent representational principle. Here, we test if this stimulus-driven convergence follows a common trajectory across people and deep neural networks (DNNs) during its transformation from sensory to high-level internal representations. We introduce a unified framework that traces representational flow by combining inter-subject similarity with alignment to model hierarchies. Applying this framework to three independent fMRI datasets of visual scene perception, we reveal a cortex-wide network, conserved across individuals, organized into two pathways: a medial-ventral stream for scene structure and a lateral-dorsal stream tuned for social and biological content. This functional organization is captured by the hierarchies of vision DNNs but not language models, reinforcing the specificity of the visual-to-semantic transformation. These findings show a convergent computational solution for visual encoding in both human and artificial vision, driven by the structure of the external world.
1 Introduction
Visual perception arises from the brain’s ability to internalize statistical regularities in the environment, transforming raw sensory input into structured, meaningful experience. Yet whether perception primarily reflects these external patterns or is shaped by the intrinsic computational architecture of the brain remains a central issue in cognitive science. This longstanding debate traces from philosophical perspectives like Plato’s allegory of the cave and theories of direct realism 1, which posit that perception is a direct reflection of the external world, to modern constructivist theories, which frame perception not as a passive imprint but as an active, model-based inference process 2, 3. A stimulus-driven view predicts that individuals will form convergent neural representations aligned with environmental structure4. Conversely, if internal constraints dominate, representations may diverge, reflecting idiosyncratic predictive models 5, 6, 7. Understanding the balance between these external and internal constraints is therefore foundational to any theory of cortical representation.
Empirical evidence supports both sides of the debate. Substantial individual variability has been observed in low-level perceptual processes such as motion detection, contrast sensitivity, and color perception, linked to differences in cortical anatomy and function 8, 9, 10. In contrast, studies using naturalistic paradigms show that when different individuals process identical stimuli, they display similar activity patterns, indicating a shared representational organization across brains 11, 12, 13, 14. This suggests that despite anatomical and experiential variability, common transformation pathways reliably lead neural computation toward convergent representational outcomes.
The principle of stimulus-driven convergence extends beyond the human brain to deep neural networks (DNNs). A substantial body of research has established a hierarchical correspondence between DNN layers and the processing stages of the visual system—a relationship that holds across diverse sensory modalities15, 16, 17, 18, 19. This alignment makes DNNs particularly effective probes for how sensory input is transformed into abstract, conceptual representations20, 21, 22. By providing an explicit, layer-by-layer mapping from low-level features to high-level semantics, DNNs offer a powerful model system for isolating the transformations that occur along the cortical hierarchy. Thus, characterizing a brain region’s alignment across the DNN hierarchy provides a functional fingerprint of its abstraction level and enables a data-driven approach for tracing information flow through the cortex.
Here, we investigated whether the transformation of visual information into abstract meaning unfolds along a shared, systematic trajectory across individuals and whether this trajectory is reflected in the representational dynamics of DNNs. To address this, we introduce a unified, data-driven framework that combines inter-subject alignment to identify cortical areas and networks with shared geometry, model-based alignment to map the hierarchy, and decomposition techniques to interpret the stimuli features driving the correspondence (see Fig. 1). We applied this framework to the massive-scale Natural Scenes Dataset (NSD) 23 and further validated the generalization of our findings in two independent fMRI datasets–BOLD5000 24 and THINGS-fMRI 25–with different image-viewing tasks.
Our analysis applied to scene perception revealed a cortex-wide representational network that is shared across individuals and organized into two principal streams with distinct functional roles. One stream follows a medial-ventral trajectory dedicated to encoding visually grounded scene structure, aligning with the classical ventral pathway 26, 27, 28. The other traces a lateral-dorsal route, selectively tuned to biological and social content, offering empirical support and a data-driven refinement for recent proposals of a ‘third visual pathway’ for social perception 29, 30. This convergence across brains and artificial systems provides a principled framework for mapping how stimulus-driven information propagates through cortical systems, revealing the functional architecture that supports the transformation of sensory information into abstract representations.
Figure 1: A unified framework for tracing representational pathways. (A) Feature Extraction. For each image stimulus, we extracted equivalent representations from brain activity and deep neural networks. Single-trial fMRI responses were aggregated within HCP-MMP cortical parcels 31 to create vector representations of the brain’s response. Concurrently, layer-wise activations were extracted from pre-trained vision and language models to obtain representations across the full model hierarchies. Both brain and model vectors were used to compute Representational Dissimilarity Matrices (RDMs) based on their pairwise correlations. (B) Representational Alignment. Representational Similarity Analysis (RSA) 32 was used to compare RDMs, quantifying the alignment between brain parcels and model layers. (C) Representational Connectivity. Inter-parcel RSA was used to construct a cortical network based on shared representational geometry. Analyzing this network revealed the hierarchical flow of information and identified key representational hubs. (D) Shared Dimensions. Within these hubs, Kernel Multi-view Canonical Correlation Analysis (KMCCA) 33 was used to isolate a subspace of representational dimensions shared across all participants. This final step links the data-driven representational axes to specific properties of the scene content, allowing for a functional interpretation of the information driving the alignment across the network.
2 Results
2.1 Inter-subject convergence in stimulus representation
A central prediction of stimulus-driven perception is that different individuals processing the same stimuli should converge on a common representational geometry 13, 32. For the natural image-viewing tasks studied here, this implies that stimulus structure should drive representational alignment across observers in relevant cortical areas–from early visual cortex through high-level visual regions, and potentially into frontal areas implicated in visually guided decisions 34, 35. While inter-subject measures in naturalistic fMRI are often interpreted as ceiling estimates of signal-to-noise 23, 36, 37, we instead treated high inter-subject RSA (IS-RSA) as direct evidence that a region encodes stimulus information in a format shared across individuals. Identifying this cortical map of shared representational geometry thus provides a foundation for characterizing the spatial organization of stimulus-driven representations and serves as a starting point for examining how visual information is transformed across distinct brain regions.
To evaluate this prediction, we focused our analysis on the Natural Scenes Dataset (NSD) 23. This dataset comprises 7T fMRI recordings in which eight participants viewed thousands of naturalistic scenes images in a memory task, each presented up to three times across sessions spanning a full year. For each stimulus, we formed a beta vector for each HCP-MMP parcel by concatenating the single-trial responses of its voxels 31. We then computed a parcel-wise Representational Dissimilarity Matrix (RDM) for each subject and used Representational Similarity Analysis (RSA) to quantify the geometry alignment of these neural patterns 32. To isolate the stable, perception-driven component of the response from potential confounds like session timing, we computed inter-subject alignment (IS-RSA) using a ‘shifted-repetition’ matching. This method compares RDMs between subjects using responses to the same images but from different trial repetitions (e.g., the first repetition for one subject vs. the second for another), ensuring alignment is driven by stimulus content alone (see Supplementary Fig. A2 for a schematic; main pipeline in Fig. 1). Statistical significance was then assessed against a stimulus-shuffled null distribution (10,000 repetitions) to identify parcels with a significant shared geometry (see Methods).
IS-RSA captured significant inter-subject alignment throughout the visual hierarchy (Figs. 2a–b; see Supplementary Fig. D1 for full parcel-level statistics). We observed an alignment gradient that peaked in early visual cortex and extended into high-level ventral occipito-temporal and dorsal occipito-parietal regions–the main areas forming the visual system hierarchy34, 38, 35. These regions are also linked to high-level functions like scene perception and place memory 27, 39. Weaker, yet significant, alignment was also found in prefrontal cortex (PFC) areas typically linked to visual attention and task-dependent semantic processing and anterior areas associated to 40, 41. This spatial pattern is consistent with prior analyses of the NSD dataset using signal-to-noise and decoding methods23, 42, 43. Although some hemispheric asymmetries in alignment strength were present (Fig. 2b), the representational profiles of symmetric parcels in the left and right hemispheres were highly correlated, suggesting that both hemispheres encode a similar geometric structure (Supplementary Fig. A1). These results support the hypothesis that scene statistics impose a shared representational organization from low-level visual areas to higher-order association cortex.
Three anatomically clustered groups of parcels showed a consistent alignment across participants (Fig. 2b). The first is early visual cortex (V1–V4), region associated with processing low-level visual features 44. The second is a ventral hub spanning key regions for processing scene context and spatial layout (parcels: VMV1-3, PHA1-3) 27, 45, 28. The third is a dorsal hub in the lateral occipitotemporal cortex (LOTC), encompassing the MT+ complex and the temporoparietal junction (parcels: MT, MST, FST, V4t, TPOJ2-3). These regions, which constitute an extension of the dorsal stream beyond classic spatial processing, are typically linked to motion integration, social cognition, and multimodal processing 12, 30, 46, 39. In the following sections, we use these three hubs—early visual, ventral, and LOTC—as reference points for our subsequent analyses.
Figure 2: Cortical distribution of representational alignment. (A) Inter-subject alignment (RSA, Pearson $r$) computed per parcel and grouped by macro-anatomical clusters 47. Box-plots show the ten clusters with the highest mean alignment in the NSD sample ($N=8$; symmetric HCP-MMP atlas 31). Red lines mark the parcel-wise null distribution (mean ± s.d.; 10 000 label permutations). All displayed parcels exceed chance (two-tailed, FDR-corrected) at $p<0.001$, except EC and Pfm ($p<0.01$), 31pd ($p<0.05$) and 23d/31pv (n.s.). Peak clusters include early visual cortex (V1–V4), the ventral hub (VMV1–3, PHA1–3) and the LOTC hub (V4t, MT, MST, FST, TPOJ2–3). Full-atlas results appear in Supplementary Fig. D1. (B–D) Surface maps of inter-subject (B), vision-model (C) and language-model (D) alignment. RSA values are averaged across subjects and—where applicable—across models. (E–F) Parcel-wise relation between inter-subject alignment and model–brain alignment for vision (E) and language (F). Vision parcels follow a power-law fit ($R^{2}=0.94$, 95% bootstrap CI); language parcels cluster near zero except for those in the LOTC hub. (G) Mean alignment by modality within the three hubs. Paired $t$-tests ($t(7)$, two-tailed, FDR-corrected) show higher vision-model alignment in early-visual and ventral hubs, and higher language-model alignment in the LOTC hub (all $p<0.0001$).
2.2 Model–Brain Alignment Mirrors Inter-Subject Shared Geometry
Our previous analysis revealed a shared representational geometry across individuals; however, the nature of the information underlying this alignment remains unspecified. To determine whether this shared geometry is driven by low-level visual features or by more abstract, conceptual content, we compared the neural data to two distinct classes of computational models. Deep vision models, which learn representations directly from pixel data, provide a template for a visually-grounded processing hierarchy48, 49, 50. In contrast, large language models, which learn from text tokens, provide a template for modality-independent semantic and conceptual relationships. By systematically comparing cortical activity to both model types, we can therefore functionally characterize brain regions. This approach allows us to determine whether their organizing principles are primarily visual or more abstract in nature.
We analyzed a diverse collection of vision and language models based on Transformer and Vision Transformer (ViT) architectures, spanning supervised, self-supervised, and contrastive training objectives (see Table 1). For each model, we computed the layer-wise RSA between its layer activations and every HCP-MMP parcel (Fig. 1; Methods). To obtain a single summary measure of model-parcel correspondence, we then identified the maximum RSA value across all layers. This max-RSA value estimates the peak alignment achieved between a model and a brain region, irrespective of the specific hierarchical stage at which it occurs. This approach yields a cortex-wide alignment map directly comparable to the IS-RSA map (Fig. 2c–d). Statistical significance was assessed via a two-tailed permutation test (10,000 permutations; see Supplementary Fig. D1 for full atlas results).
Vision-trained models closely reproduced the inter-subject map. Across HCP parcels, a parcel’s vision-model alignment score scaled tightly with its IS-RSA value, following a power-law relationship ($R^{2}=0.94$;Fig. 2e; see Supl. analysis A.6 for details). Spatially, the cortical alignment map (Fig. 2c) mirrored the inter-subject pattern (Fig. 2b), with the highest alignment in the Early Visual, Ventral, and LOTC hubs, and significant, albeit weaker, alignment throughout the remaining visual areas and prefrontal cortex. This tight correspondence suggests that the cortical parcels where stimulus geometry is shared across individuals are precisely those whose representational structure is captured by image-trained networks, indicating that these models and the human brain converge on a common, stimulus-driven coding scheme for natural scenes.
Language models trained on image captions displayed a distinct and more localized alignment pattern. The cortical alignment map showed no strong alignment in early visual or ventral areas. Specifically, the classical visual streams exhibited negative RSA scores, indicating dissimilar representational geometries that contrast with visually grounded organizations (Fig.2d; Supplementary Fig.D2). In contrast, robust positive alignment emerged in the LOTC hub and neighboring parcels (Fig.2f), indicating that alignment in these regions is not visually grounded. A direct comparison of modalities revealed that language-model alignment exceeded vision-model alignment within this LOTC hub (Fig.2g).
These results serve as a modality control, indicating that while alignment in early and ventral areas is tied to visually grounded features, the information shared in the LOTC hub is organized around a non-visual structure captured by language embeddings of image descriptions. Beyond this modality dissociation, we next aimed to situate cortical regions along the visual-to-semantic transformation trajectory. To this end, we examined their layer-wise alignment profiles, using the model hierarchy as a proxy for the brain’s own computational stages.
Figure 3: Hierarchical convergence between models and cortex. (A–C) Layer-wise alignment (RSA) between vision models and brain activity for representative parcels within the three reference hubs. The alignment curves are averaged across all vision models used in the study. Panels show Early Visual Cortex (A), the Ventral Hub (B), and the LOTC Hub (C). Lines show the mean alignment across participants ($N=8$), with shaded areas representing the standard error of the mean (SEM). Alignment profiles differ by hub, showing early, distributed, and late-stage peaks, respectively. (D) Cortical surface map of the vision model layer with the highest alignment for each parcel. The color scale indicates the normalized depth of the peak layer (0% shallower layer, 100% deeper layer), revealing a posterior-to-anterior gradient from shallow (green) to deep (purple) layers. (E) Layer-wise alignment between language model and representative parcels from each of the three hubs. In contrast to vision models, only parcels within the LOTC hub show strong alignment, which follows a step-like function rather than a smooth progression. (F) Box plots showing the distribution of peak alignment depths for the 20 parcels with the highest overall vision-model alignment. Parcels are sorted by their median peak depth, illustrating a continuous hierarchical organization spanning from early visual areas to the LOTC hub.
2.3 Hierarchical correspondence Between Models and Cortex
To obtain a more precise mapping of how brain regions correspond to different levels of model abstraction, we moved beyond single summary score and examined the layer-wise alignment profiles for each cortical parcel. While numerous studies have established a coarse hierarchical correspondence showing that early model layers align with V1–V4 and deeper layers map onto downstream regions 15, 16, 17, 51, 52, it remains unclear whether this is a simple one-to-one mapping. Alternatively, certain regions might exhibit a mixed alignment by integrating features from multiple computational stages. To address this, we conducted a layer-wise RSA for each parcel, yielding an alignment profile across model depth (see Methods; Fig. 3).
This layer-wise analysis of vision models showed distinct alignment profiles for each of the three main cortical hubs (Fig.3a–c). Alignment with the early visual cortex peaked in the shallowest model layers before decreasing, consistent with its role in processing low-level features (Fig.3a). In contrast, the ventral hub exhibited a broad plateau of moderate alignment across many layers, suggesting it integrates both visual features from earlier stages and more abstract information from deeper stages (Fig.3b). The LOTC hub, in turn, showed a progressively increasing alignment that peaked only in the deepest, most semantic layers (Fig.3c; see Supplementary Figs. D3D4 for extended parcels and model-family details).
To make the functional hierarchy explicit, we mapped the peak alignment layer of vision models for each parcel. Despite the mixed correspondence observed in some ventral regions, this analysis revealed a smooth posterior-to-anterior gradient across the visual system (Fig. 3d). This gradient was most apparent within visual cortex, whereas frontal areas exhibited uniformly flattened alignment profiles, reflecting mixed correspondence with features spanning multiple model stages (see Supplementary Fig. D3). Nevertheless, the ordered progression from low-level to high-level representations was highly consistent across participants in visual areas (Fig.3f), providing evidence for a shared progression toward abstraction mirrored between cortex and models.
Language models, in contrast, showed no cortex-wide progression; strong alignment emerged only in the LOTC hub (Fig. 3e). Elsewhere, values hovered near zero across layers (Supplementary Fig. D3). Within LOTC, the alignment curve followed a step-like profile: it began with a weak score that was highly correlated with the token-level embeddings fed into the first layer (Pearson’s $r=0.78$; tokenizer control, Supplementary Fig. A4), consistent with work that suggest that brain–language alignment may reflect discrete word properties rather than compositional meaning 53. After this initial stage, alignment increased sharply before reaching a stable maximum, with the final value proportional to the first-layer alignment (Pearson’s $r=0.97$). This pattern indicates that LOTC representations are not late-stage vision, but are instead organized around semantic dimensions driven by token representations in language, with alignment increasing once language models have formed stable embedding representations.
Although these analyses showed that some parcels exhibited mixed correspondence with features from different model stages, they revealed a cortical processing hierarchy that was partially mirrored in the architecture of deep vision models. The distinct, non-hierarchical profile observed for language models further highlighted this structure, isolating the LOTC hub as a region organized around a different, non-visual representational principle. Having established the vision model hierarchy as a proxy for the brain’s own computational stages, we proceeded to examine how information is routed between cortical regions.
Figure 4: Inter-subject representational connectivity network. (A) Inter-subject connectivity matrix. Parcel-wise representational connectivity (RSA, Pearson’s $r$) is shown for 30 key cortical parcels ($N=8$). Each cell represents the mean RSA score between the RDMs of two parcels, computed across all pairs of different individuals. The matrix reveals a clear block structure corresponding to three principal hubs: Early Visual Cortex (V1–V4), the Ventral Hub (PHA1-3, VMV1–3), and the LOTC Hub (MT, MST, V4t, FST, TPOJ2–3). Warm colours denote shared representational geometry; cool colours denote dissimilarity. (B) Directed connectivity graph. The same 30 parcels are plotted as nodes. Node colour indicates each parcel’s peak alignment depth with vision models, node size reflects its inter-subject alignment strength, and the node border colour marks its macro-anatomical group. The edges shown are the union of the three minimum-spanning trees (see Methods). Edge colour encodes the RSA value from panel (a), while edge width indicates the spanning tree order (from first to third), highlighting the network’s most central pathways. Arrowheads denote the putative direction of information flow, inferred from significant differences in peak-depth between connected parcels (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$). The layout reveals two primary processing streams emerging from early visual cortex: a medial-ventral route and a lateral-dorsal route projected to the LOTC hub.
2.4 Representational connectivity reveals an ordered processing network
To characterize the large-scale organization of these functionally distinct hubs, we extended our analysis from isolated regional analysis to the connectivity level. We constructed a parcel-by-parcel representational connectivity matrix by correlating the representational geometry of each parcel in one individual with that of every other parcel in all other individuals. This inter-subject representational connectivity approach directly captures a network structure that is reproducibly shared across the population. To estimate the direction of information flow within this network, we used the peak alignment depth from our vision model analysis as a proxy for hierarchical position, under the assumption that information flows from regions preferring shallower layers to those preferring deeper ones–a principle consistent with feed-forward processing in both the visual system and deep networks 16, 17. The same network structure was obtained using a conventional within-subject approach, confirming the robustness of our findings (see Supplementary Fig. D5).
The parcel-by-parcel connectivity matrix revealed three internally connected clusters corresponding to the Early Visual, Ventral, and LOTC hubs (Fig. 4a). The Early Visual and Ventral hubs were moderately connected, forming an integrated system for visually grounded representation. In contrast, the LOTC hub exhibited a dissimilar representational geometry (negative RSA) relative to both the Ventral hub and classical dorsal regions (e.g., intraparietal sulcus). This pattern supports its functional dissociation as a distinct processing pathway, not only from ventral scene-perception areas like the PPA 54, 55, but also from other dorsal regions involved in spatial processing—a dissociation consistent with proposals for a distinct lateral or “third” visual pathway 30, 56.
The directed graph, which highlights the network’s main connections, made the two-stream topology of information flow explicit (Fig. 4b). Representational similarity originated in early visual cortex and bifurcated into two primary branches: a medial-ventral stream that passed through the Ventral hub, and a lateral-dorsal stream that traversed the LOTC hub—mirroring the dual-network organization previously described for scene processing 54. Our whole-cortex rendering (Supplementary Fig. D6) showed these streams extending anteriorly into posterior parietal, prefrontal, and default-mode regions, outlining a plausible route by which perceptual information can reach circuits for attention, memory, decision-making, and motor planning, consistent with known anatomical and functional pathways.
In sum, our representational connectivity analysis identifies a stimulus-driven network whose core structure is consistent across individuals and is consistent with the dual-network topology previously reported for scene perception 54. This network originates in early visual cortex and bifurcates into two primary processing streams. The medial-ventral stream, terminating in the Ventral hub, corresponds to the classical pathway for scene and object recognition 26, 27. The lateral-dorsal stream, culminating in the LOTC hub, is highly consistent with recent proposals for a third visual pathway specialized for social perception 30. This functional dissociation raises the question of what information defines each of these processing routes—a question we address in the next section by isolating the dominant representational dimensions within each hub.
2.5 Content Decomposition of Representational Hubs
Figure 5: Dominant representational dimensions in cortical hubs. (A–C) Stimuli projection onto the first two shared components for each hub, extracted via Kernel Multi-view CCA (KMCCA) using the voxel data from all eight NSD participants for the common image set. In Early Visual Cortex (A), stimuli form a category-free cloud reflecting low-level visual similarity. In the Ventral Hub (B), the dominant axis arranges stimuli along a scene-to-object gradient. In the LOTC Hub (C), the first component separates animate from inanimate stimuli. (D) Impact on inter-subject RSA of controlling for the first KMCCA component via partial RSA. Controlling for the first component significantly reduced alignment in all hubs: Early Visual (6.6%), Ventral (3.1%), and LOTC (21.8%). The pronounced RSA reduction in LOTC underscores the animacy axis as its dominant organizing principle. (E, F) Cortical maps of inter-subject RSA recomputed separately for those scenes without (E) versus with (F) biological agents (people or animals). Alignment in the LOTC hub collapses when biological content is absent. Partial RSA values in (D) were computed parcel-wise and averaged by hub. All paired $t$-tests are Bonferroni-corrected ($t(7)$, $p<0.001$).
The preceding analyses showed that the Early Visual, Ventral, and LOTC hubs form three distinct clusters embedded in two major processing streams. To uncover the content encoded in each hub, we decomposed the shared representational geometry using Kernel Multi-view Canonical Correlation Analysis (KMCCA)33. Standard RSA collapses two RDMs into a single Pearson’s $r$ that quantifies their alignment. By contrast, KMCCA treats each participant’s RDM as a separate “view” of the same stimulus-pair dissimilarities and identifies a shared subspace of orthogonal components whose projections are maximally correlated across all views. Our use of KMCCA parallels multi-view CCA techniques57–but instead of operating on raw voxel data, it works directly on the correlation kernels. These canonical axes therefore decompose the overall inter-subject alignment into ranked representational dimensions, isolating the latent components that drive shared geometry.
We focused our analysis on the first KMCCA component, which explains the largest fraction of inter-subject alignment. First, we identified the stimulus features associated with this component to interpret its semantic content (details in Supplementary Fig. A6). Second, we used partial RSA to quantify this component’s contribution to the overall alignment. In this analysis, a large reduction in RSA after controlling for the component would indicate a geometry dominated by that single axis, whereas a minimal change would imply a more complex, high-dimensional space with no single dominant factor.
The KMCCA projections revealed a distinct organization for each hub. In Early Visual Cortex (Figs. 5a and D8), stimuli formed a category-free cloud organized primarily by low-level visual features, consistent with its retinotopic properties rather than high-level semantics. In the Ventral hub (Figs. 5b and D9), the dominant axis traced a gradient from panoramic scenes to isolated objects, recapitulating the classic scene-to-object gradient in ventromedial cortex 27, 54. By contrast, the LOTC hub (Figs. 5c and D10) was organized along an animacy axis, with its first canonical component separating stimuli containing biological agents (people or animals) from inanimate scenes.
[…]
To determine whether this alignment drop in the LOTC hub was uniform across all images or driven specifically by biological content, we split the image set into stimuli containing animate agents (65.3%) and those without, then recomputed inter-subject alignment. The results confirmed the LOTC hub’s specialization: alignment vanished in the absence of biological elements in the scene, while it remained robust for scenes containing people or animals (Fig. 5e-f). An analysis of the connectivity network using this same data split showed that the entire lateral stream from the Early Visual Cortex to the LOTC hub was significantly reduced without biological content (Supplementary Fig. D7). Together, these findings provided support from a representational approach for a lateral “third visual pathway” specialized in social and animate information 30, 56.
Figure 6: Shared representational geometry is modulated by stimulus content and task. Alignment computed using the symmetric HCP atlas combining both hemispheres. (A–D) Replication in BOLD5000, using complex natural scenes. With a different task (valence rating) but with comparable complex scene stimuli that include social content, the main findings were replicated. Both the inter-subject alignment (A) and its correspondence with vision models (B) show a similar posterior-to-anterior progression. The layer-wise curves (C) reproduce the distinct hierarchical trend for each hub, and the connectivity matrix (D) reveals a similar network structure with a representationally distinct LOTC hub. (E–H) Replication in THINGS-fMRI, using object-centric images. With a stimulus set of isolated objects and a rapid oddball task, the extent of shared geometry is more restricted. Inter-subject alignment (E) and its correspondence with vision models (F) are largely confined to early visual cortex. While the layer-wise curves (G) retain their characteristic shapes, the connectivity matrix (H) is dominated by a single Early Visual hub. This suggests that rich scene and social content are necessary to engage the full processing hierarchy captured by shared representational geometry.
2.6 Generalization across tasks and stimulus sets
Our findings derived from NSD picture dataset showed that stimulus statistics can align brains with vision models during a long-format memory task, but representational geometry may depend on both image distribution and task context58, 59. To assess the degree to which our findings generalize to other encoding task environments, we replicated our analysis pipeline on two independent fMRI datasets: BOLD5000 (valence judgments on  5,000 natural scenes)24 and THINGS-fMRI (oddball detection on  9,000 object images)25. For each dataset, we derived single-trial voxel responses, computed inter-subject RSA maps, generated parcel-wise alignment scores for vision models, and computed a representational connectivity matrix using inter-subject, parcel-pair RSA (see Supplementary Section B for details).
To test the generalization of these findings, we first analyzed the BOLD5000 dataset 24. The stimulus set for BOLD5000 is highly comparable to NSD, containing complex natural scenes with multiple objects and social content, including a large subset of images drawn directly from the same database. Despite this stimulus similarity, participants performed a different task (valence judgment) with a shorter exposure time. Nevertheless, the results on images from BOLD500 dataset results largely replicated our main findings. Inter-subject and vision-model alignment spanned early visual cortex, ventral occipito-temporal parcels, and a LOTC hub (Fig. 6a,b), and the parcel-wise relationship between these two measures was well-described by the same power-law model ($R^{2}=0.70$). Furthermore, the layer-wise alignment curves recapitulated the same hierarchical trend observed in NSD: decreasing for the Early Visual Cortex, mixed for the Ventral Hub, and increasing for the LOTC Hub (Fig. 6c). The representational connectivity matrix also revealed a similar core network structure, with the LOTC hub forming a separate subsystem (Fig. 6d). These results indicate that the stimulus-driven brain-model convergence is robust to changes in task and scanning parameters.
Finally, we analyzed the THINGS-fMRI dataset 25, which presents a strong contrast in both stimulus content and task design. This dataset used briefly flashed (500ms), object-centric images that lack the complex scene and social content of NSD, and participants performed an orthogonal oddball task. We found that both inter-subject and vision-model alignment were largely confined to the early visual cortex (Fig. 6e,f), though the strong parcel-wise relationship between the two measures was preserved (power-law fit, $R^{2}=0.81$). Despite this restricted spatial extent, the underlying hierarchical profiles for each hub type were maintained: layer-wise curves still showed the characteristic decreasing, distributed, and increasing patterns (Fig. 6g). Consequently, the connectivity analysis revealed a single cluster in early visual cortex (Fig. 6h). This is consistent with a stimulus set lacking complex scene information and social interactions, and a passive task design that likely reduced the representational signal in higher-order cortical areas 55.
Across three distinct datasets, vision-model alignment was spatially co-located with inter-subject alignment, a finding consistent with models capturing the stimulus-driven geometry shared across brains. However, the extent of this processing cascade was context-dependent. Rich, complex scenes engaged the full posterior-to-LOTC hierarchy, whereas isolated objects under brief exposure primarily engaged only early visual cortex. These results therefore showed that shared representational geometry is a robust framework for revealing the specific cortical networks where stimulus information is processed across both individuals and models.
3 Discussion
Our study establishes a general, data-driven approach for mapping information processing routes in the human brain. By combining inter-subject representational alignment, model-based comparisons, and explainability techniques, we identified a large-scale, stimulus-driven cortical network that is conserved across individuals and robust to task and dataset variation. This network, recruited during natural scene perception, is organized into two primary streams: a medial-ventral pathway for visually grounded scene analysis, and a lateral-dorsal pathway specialized for social and biological content. This result provides a data-driven refinement of the canonical dual-stream model of vision, functionally supporting recent proposals for a “third visual pathway”30, 56 and demonstrating a complex functional architecture that is integrated into a whole-cortex network.
The convergence we observed between brain and model representations—though not mechanistically identical—suggests that deep networks provide a rich "bag-of-features" capable of capturing multiple stages of cortical processing. This nuanced view is necessary to explain our findings for the high-level hubs. The Ventral hub, with its cumulative integration of both visual and abstract information, showed alignment across a wide range of model layers. The LOTC hub, in contrast, showed a highly specialized alignment: it corresponded almost exclusively with the deepest semantic layers and only for those stimuli containing biological content. These results indicate that a simple one-to-one hierarchical model cannot account for the distinct computational strategies of the cortex, where recurrence and functional specialization contrast with the strictly sequential architecture of the models used for comparison60, 61, 62. However, the strong correspondence between the alignment found across brains and with models suggests that despite these mechanistic differences, both systems are sensitive to the same underlying stimulus structure, captured in regions where this stimulus-driven information is most functionally relevant.
The observation that brains and deep networks converge on similar representational geometries likely reflects shared constraints for efficiently encoding the structured statistics of the environment 5, 6, 63. The principle of efficient coding posits that the optimal way to compress sensory information is to accurately predict it; therefore, any system that learns to predict the world will necessarily develop compressed, efficient representations. The convergence we observe across brains and models likely arises because both systems, through different optimization processes (evolution and gradient descent), are finding similar solutions to this fundamental problem. This is increasingly recognized as a general principle in both neuroscience and artificial intelligence, with modern models converging on similar latent features across tasks and modalities 64, 65. Characterizing shared representational spaces thus provides a principled way to understand how both biological and artificial systems learn to extract meaningful information from their environments 4.
Methodologically, this work underscores the potential of population-scale, data-driven analyses to move beyond static anatomical maps, enabling systematic tracing of functional routes in the brain. No single method used here would have sufficed; it is the combination of identifying a shared space across subjects, using model hierarchies to probe its content, and deploying explainability techniques to interpret the results that provides a complete picture. With increasing access to high-quality, large-scale datasets, this integrated approach provides a path for hypothesis generation across sensory, cognitive, and cross-modal domains. The integration of explainability methods further allows for direct interpretation of what dimensions or features are routed through specific pathways, with implications for theory and for applications such as decoding and brain-computer interfaces.
Limitations remain, including the need for deeper mechanistic models linking representational geometry to neural computation and behavior, and for extending these methods to causal and cross-modal analyses. As recent work highlights, high RSA scores can arise from non-mechanistic similarities when using complex stimuli with inherent confounds 59. While our explainability analysis is a first step toward mitigating this, future work will require more precise methods to isolate the specific features driving alignment and to bridge the gap from representational similarity to true mechanistic understanding. Nonetheless, this study demonstrates that combining shared representational geometry across individuals with deep model features is a powerful framework for mapping and interpreting brain function at scale.
Our findings highlight the importance of tracing not just where, but how and what information is routed through the brain. This provides a principled, generalizable methodology for the field—moving from describing correspondence to generating more precise, mechanistically testable hypotheses about information processing in the brain.
4 Methods
4.1 Functional-MRI data
For the main analyses of this study, we used data from the Natural Scenes Dataset (NSD)23, a high-resolution 7T fMRI dataset comprising recordings from 8 participants performing a visual memory task. On each trial, participants viewed a naturalistic image for 3 seconds and indicated whether it was old or new. The stimulus set included a wide variety of natural scenes, spanning indoor and outdoor environments, static object views, and complex multi-element situations, including a broad range of social contexts. Each participant completed between 32 and 40 sessions, with 750 trials per session conducted over the course of a year. For participants who completed all 40 sessions, each image was presented three times, yielding a total of 30,000 trials per participant. Full experimental details are provided in Allen et al. (2022)23.
Analyses were based on the 1.0-mm volumetric preprocessing of the NSD dataset and the corresponding single-trial $\beta$-estimates (version 3) 23. We used GLM-denoised single-trial $\beta$-estimates 66, as their image-related information content has been validated in multiple decoding and image reconstruction studies 42, 67, 68. Voxels were grouped into cortical parcels using the HCP-MMP1.0 atlas 31, which defines 180 cortical regions per hemisphere. We employed the pre-aligned parcel masks provided by the NSD authors to extract voxel activity within each cortical region 23. Analyses were performed using both the symmetric version of the atlas (combining hemispheres) and the asymmetric version (preserving hemisphere separation); a supplementary analysis confirms that symmetric parcels display highly correlated representational profiles (Supplementary Fig. A1). In all figures, parcels are colored by macro-anatomical groups as defined in Huang et al. (2022) 47, to facilitate anatomical interpretation.
For each participant $p~{}\in~{}\mathcal{P}$, parcel $r~{}\in~{}\mathcal{R}$, and session $s~{}\in~{}\mathcal{T}p$, we define a response matrix $X_{p,r,s}~{}\in~{}\mathbb{R}^{n\times v_{p,r}}$ containing the voxel responses to each stimulus in that session. Here, $n$ denotes the number of trials (750 per session), and $v_{p,r}$ the number of voxels in parcel $r$ for participant $p$.
Comparative analyses were performed using fMRI data from the BOLD500024 and THINGS25 datasets. Preprocessing and analysis followed the same procedure described above, with minor adaptations to accommodate differences in stimulus sampling and repetition structure. Full details for these datasets are provided in Supplementary Section B.
All data analysed are fully de-identified and publicly available. The original NSD collection was approved by the University of Minnesota Institutional Review Board (Allen et al.23). The THINGS dataset was collected under NIH Institutional Review Board approval (Hebart et al.25), and BOLD5000 under the Institutional Review Board of Carnegie Mellon University (Chang et al.24). No new human-subjects data were acquired for this work.
4.2 Model Features Extraction
To compare how deep learning models encode visual and linguistic stimuli relative to fMRI responses, we selected vision and language models previously used by Huh et al. (2024) in studies of representational convergence across modalities 64. Model selection was conducted prior to analysis, balancing computational constraints and the goal of covering diverse training objectives and model scales. A full list of models is provided in Table 1.
We selected 17 vision models spanning families with distinct training regimes, including ViT AugReg69, CLIP70, 71, DINOv272, and MAE73. All architectures were based on sequential Vision Transformer (ViT) blocks74. Each NSD image was processed through the backbone of each model, and activations were extracted from the output of each ViT block (see Fig. 1A).
For language models, we selected 14 transformer-based encoders from the BLOOMZ75, Gemma 276, LLaMA77, 78, and LLaMA 379 families80. For each NSD image, we used text captions annotated by trained raters and originally sourced from the MS-COCO dataset81. Captions were processed through each model’s encoder, and activations were extracted at the output of each transformer block.
Feature extraction was performed using the Transformers82 and TIMM83 libraries, adapting procedures from Huh et al. (2024)64. All models were implemented in PyTorch84 and retrieved from the Hugging Face Model Hub.
For each model $m\in\mathcal{M}$, layer $l\in\mathcal{L}_{m}$, and stimulus set $s\in\mathcal{T}$ (images or their corresponding captions), we constructed a matrix of layer activations $Z_{m,l,s}\in\mathbb{R}^{n_{s}\times v_{m}}$, where $n_{s}$ is the number of stimuli in set $s$ and $v_{m}$ is the model’s embedding size, which is constant across all layers in the selected models. For each model and stimulus set, the activations across layers define a sequence of latent representations, reflecting the hierarchical transformation of stimulus content within each model.
Modality Family Training Regime Model # Params # Blocks Embedding size
Vision ViT (AugReg)69 Supervised (21 K classes) ViT (AugReg) Tiny 10M 12 192
ViT (AugReg) Small 30M 12 384
ViT (AugReg) Base 103M 12 768
ViT (AugReg) Large 326M 24 1024
CLIP (Vision)70, 71 Contrastive Image–Text CLIP (ViT) Base 86M 12 768
CLIP (ViT) Large 304M 24 1024
CLIP (ViT) Huge 632M 32 1280
Contrastive I-T + fine-tuning (12 K) CLIP (ViT) Base ft-12k 95M 12 768
CLIP (ViT) Large ft-12k 315M 24 1024
CLIP (ViT) Huge ft-12k 646 M 32 1280
DINOv272 Self-Supervised DinoV2 (ViT) Small 22M 12 384
DinoV2 (ViT) Base 87M 12 768
DinoV2 (ViT) Large 304M 24 1024
DinoV2 (ViT) Giant 1.1B 40 1536
MAE73 Self-Supervised (Masked AE) MAE (ViT) Base 86M 12 768
MAE (ViT) Large 303M 24 1024
MAE (ViT) Huge 631M 32 1280
Language BLOOMZ75 Causal LM + Instruction FT BloomZ (560M) 559M 25 1024
BloomZ (1b1) 1.1B 25 1536
BloomZ (1b7) 1.7B 25 2048
BloomZ (3B) 3B 31 2560
BloomZ (7b1) 7.1B 31 4096
Gemma 276 Causal LM Gemma 2 (2B) 2.6B 27 2304
Gemma 2 (9B) 9.2B 43 3584
LLaMA77, 78 Causal LM OpenLlama (3B) 3.4B 27 3200
OpenLlama (7B) 6.7B 33 4096
OpenLlama (13B) 13B 41 5120
HuggyLLaMa (7B) 6.7B 33 4096
HuggyLLaMa (13B) 13B 41 5120
LLaMA 379 Causal LM Llama 3 (8B) 8B 33 4096
Llama 3.1 (8B) 8B 33 4096
Table 1: Vision and language models used in the analyses.
4.3 Alignment Measure
We used Representational Similarity Analysis (RSA)32 to quantify the correspondence between brain parcels and model layers geometries. For each system, the representational dissimilarity matrix (RDM) $D(X)$ was defined as $D(X)_{ij}=1-\rho(x_{i},x_{j})$, where $x_{i}$ and $x_{j}$ are rows of $X$ and $\rho$ is the Pearson correlation. The RSA score between two systems $X$ and $Y$ was the correlation between the upper triangles of their RDMs:
$\operatorname{RSA}(X,Y)=\rho\left(\operatorname{vec}_{u}D(X),\;\operatorname{vec}_{u}D(Y)\right)$ (1)
Main analyses were replicated with Spearman correlation and Centered Kernel Alignment (CKA)85, yielding consistent results (see Supplementary Fig.A5). Details of the matrix reformulation that enabled efficient, large-scale RSA computations are provided in Supplementary Section C.
4.4 Inter-Subject alignment
Inter-subject RSA was computed using the subset of 1,000 images viewed by all NSD participants (the shared1000 set)23. Each image in this subset was presented up to three times per participant, with corresponding trial indices across sessions. Trials were uniquely identified by participant, image ID, and repetition index $k\in{1,2,3}$. For each pair of participants $p,q\in\mathcal{P}$ and cortical parcels $r,r^{\prime}\in\mathcal{R}$, we extracted the corresponding voxel responses $X_{p,r}$ and $X_{q,r^{\prime}}$ and computed RSA between matched trials.
Images shared across participants were always presented at the same trial positions for all participants, interleaved with unique images specific to each individual. To attenuate alignment driven by task or session structure rather than by true stimulus processing, we applied a cyclic repetition shift when matching trials across participants. Specifically, for each image $i$ and repetition index $k$, the trial $(i,k)$ in participant $p$ was matched with the trial $(i,(k\bmod 3)+1)$ in participant $q$, thereby preserving stimulus identity while disrupting repetition-locked confounds. This matching strategy and its impact are illustrated and compared in Supplementary Fig. A2. Let $\boldsymbol{i}_{pq}$ and $\boldsymbol{j}_{pq}$ denote the matched stimulus indices for participants $p$ and $q$, respectively. The inter-subject alignment between parcels $r$ and $r^{\prime}$ was then computed as:
$\mathcal{A}^{\mathrm{IS}}_{p,q,r,r^{\prime}}=\operatorname{RSA}\left(X_{p,r}[\boldsymbol{i}_{pq},:],\;X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]\right).$ (2)
To compute a subject-level alignment measure for participant $p$, we aggregated alignment values with respect to all other participants:
$\mathcal{A}^{\mathrm{IS}}_{p,r,r^{\prime}}=\frac{1}{|\mathcal{P}|-1}\sum_{q\in\mathcal{P}\setminus\{p\}}\mathcal{A}^{\mathrm{IS}}_{p,q,r,r^{\prime}}.$ (3)
Group-level alignment was then obtained by averaging subject-level measures across participants. Cortical surface maps (e.g., Fig.2b) display group-level alignment, specifically comparing the same parcels across participant pairs ($r=r^{\prime}$). To benchmark the degree of shared geometry relative to individual consistency, we also compared inter-subject and within-subject alignment (see Supplementary Fig. D5), which yielded consistent results.
A parcel-by-parcel matrix of group-level inter-subject alignment values ($n_{\mathrm{rois}}\times n_{\mathrm{rois}}$) defined the whole-cortex connectivity network. We visualized the main network backbone and multiple major information routes using an iterative minimum spanning tree (k-MST) approach on the group-averaged connectivity matrix. At each iteration, a standard MST was computed (with edge weights defined as $1-\mathrm{similarity}$), penalizing previously selected edges to ensure the inclusion of new, critical pathways. The final backbone comprised the union of edges from all $k$ MSTs (see Supplementary Fig. D6), preserving multiple principal paths, avoiding arbitrary edge thresholding, and ensuring that all parcels remained connected while allowing for richer topology. MST computations were performed using the NetworkX and SciPy libraries86, 87.
Statistical significance was evaluated by comparing the observed alignment values to a null distribution generated by randomly shuffling stimulus labels (10,000 permutations, using a shared permutation scheme across all comparisons) with a two-tailed test. All p-values were FDR-corrected 88.
4.5 Brain–Model alignment
Brain–model representational alignment was computed using all available trials from NSD participants (i.e., not restricted to the set of shared images), with analyses performed separately by session to control for session-specific variance. For each participant $p\in\mathcal{P}$, model $m\in\mathcal{M}$, and session $s\in\mathcal{T}_{p}$, voxel responses from each cortical parcel $r\in\mathcal{R}$ were compared to model activations at each layer $l\in\mathcal{L}_{m}$ using RSA:
$\mathcal{C}_{p,r,s,m}(l)=\operatorname{RSA}\bigl{(}X_{p,r,s},\;Z_{m,l,s}\bigr{)}.$ (4)
To compute a subject-level alignment measure for each model, we identified the layer with the largest absolute alignment for each session, preserving the sign of the value (denoted as $\operatorname{sign\,max}$ in Eq. 5). These peak values were then averaged across sessions:
$\mathcal{A}^{\mathrm{BM}}_{p,r,m}=\frac{1}{|\mathcal{T}_{p}|}\sum_{s\in\mathcal{T}_{p}}\operatorname*{sign\,max}_{l\in\mathcal{L}_{m}}\,\mathcal{C}_{p,r,s,m}(l).$ (5)
This procedure, standard in recent literature 51, 64, ensures that alignment is captured independently of model depth. In cases—such as for language models—where correlations may be systematically negative, this approach preserves the interpretability of dissimilar representational geometries: negative alignment values indicate that stimuli which are close in one space (e.g., visual cortex) are systematically distant in another (e.g., language model embedding)89.
To identify where alignment peaked within the model hierarchy (e.g., Fig. 3b–c), we computed the normalized depth of the maximally aligned layer for each session and averaged across sessions:
$d^{*}_{p,r,m}=\frac{1}{|\mathcal{T}_{p}|}\sum_{s\in\mathcal{T}_{p}}\frac{1}{|\mathcal{L}_{m}|-1}\arg\max_{l\in\mathcal{L}_{m}}\mathcal{C}_{p,r,s,m}(l),$ (6)
where layers are indexed from 0 (first layer) to $|\mathcal{L}_{m}|-1$ (last layer), so $d^{*}_{p,r,m}\in[0,1]$. This procedure is standard for mapping the correspondence between brain regions and network hierarchy 19, 51. To obtain a subject-level measure for a set of models (e.g., all vision or language models), both the peak alignment and normalized depth measures were averaged across models. Group-level measures were then obtained by averaging subject-level values across participants.
To study how alignment changes across the model hierarchy, we constructed continuous alignment curves for each subject parcel and model by averaging layer-wise RSA across sessions. To compare models with different numbers of layers, linear interpolation was applied, normalizing layer indices to a depth variable $d\in[0,1]$ ($d=0$ for the first layer, $d=1$ for the last layer):
$\tilde{\mathcal{C}}_{p,r,m}(d)=\operatorname*{interp}_{l\to d}\left(\frac{1}{|\mathcal{T}_{p}|}\sum_{s\in\mathcal{T}_{p}}\mathcal{C}_{p,r,s,m}(l)\right),\quad d\in[0,1].$ (7)
Alignment curves for each subject’s parcel across a set of models $\mathcal{M}$ (e.g., all vision models or all language models) were then obtained by averaging interpolated curves across models:
$\tilde{\mathcal{C}}_{p,r,\mathcal{M}}(d)=\frac{1}{|\mathcal{M}|}\sum_{m\in\mathcal{M}}\tilde{\mathcal{C}}_{p,r,m}(d).$ (8)
Statistical significance was assessed using a permutation test with 10,000 stimulus-label shuffles. The null distribution was generated by applying the same set of permutations across all model layers and parcels, allowing direct comparison and group averaging. Two-tailed p-values were calculated by comparing the observed group-mean alignment to this null, and corrected for multiple comparisons using the Benjamini–Hochberg FDR procedure.
4.6 Shared Component Extraction
To identify the dominant dimensions underlying shared representational geometry across participants, we applied kernel multi-view canonical correlation analysis (KMCCA) 33. KMCCA was performed on the set of image trials shared by all participants in NSD, extracting a low-dimensional subspace that maximized correlation across participants’ RDMs for each hub.
We used the KMCCA implementation from mvlearn90. For each hub, we extracted the first two components for exploratory purposes, and assessed the influence of the first (dominant) component in control analyses. While our focus was on the dominant axis, further work is needed to interpret all statistically significant components. Image-level semantic annotations and category labels were derived from MS-COCO 81. For automated labeling of the scene-to-object gradient, we used Pixtral-12B 91 (see Fig. 5a).
Further details of the KMCCA procedure, semantic projections, and partial RSA controls are provided in Supplementary Section A.5.
Data and code availability
Functional-MRI data and image stimuli were obtained from the Natural Scenes Dataset (accessed via https://registry.opendata.aws/nsd), BOLD5000 (10.18112/openneuro.ds001499.v1.3.0 and 10.1184/R1/c.5325683), and THINGS-fMRI (10.25452/figshare.plus.c.6161151.v1 and 10.17605/osf.io/jum2f). Image metadata were extracted from MS-COCO (https://cocodataset.org). Model checkpoints analyzed in this work are available on Hugging Face (https://huggingface.co/collections/pablomm/convergent-transformations-6808f7de248fa9674acac588). All resources were publicly accessible in July 2025.
Code to reproduce analyses is available at https://github.com/memory-formation/convergent-transformations.
Acknowledgments
We thank J. García-Arch, M. Domínguez-Orfila, D. Pacheco-Estefan, C. Baldassano, and E. Cámara for helpful comments and discussions, the creators of the Natural Scenes Dataset, THINGS, and BOLD5000 for making their data publicly available, and the teams who developed and released the models on Hugging Face. This work was supported by the Spanish Ministerio de Ciencia, Innovación y Universidades, which is part of Agencia Estatal de Investigación (AEI), through the project PID2022 - 140426NB (Co-funded by European Regional Development Fund. ERDF, a way to build Europe). We thank CERCA Programme/Generalitat de Catalunya for institutional support.
Author contributions
P.M. and L.F. conceived the project, formulated the methodology and drafted the manuscript. P.M. performed the analyses. L.F. secured funding and supervised the work. Both authors revised and approved the final manuscript.
Supplementary Material
A Supplementary analyses
A.1 Hemispheric Asymmetry and the Validity of Symmetric Analyses
Figure A1: Hemispheric comparison of representational alignment. (A–C) Comparison of representational alignment (RSA, Pearson’s $r$; $N=8$, NSD) computed independently for the left (x-axis) and right (y-axis) hemispheres: (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment. Each point is a cortical parcel, colored by macro-anatomical group. The diagonal line indicates equal alignment. Squares denote parcels with a statistically significant hemispheric difference (paired $t$-test, $p<0.05$, FDR-corrected). (D) Boxplots of alignment within principal hubs. Significant hemisphere effects are marked. (E) Correlation between cross-hemispheric and bilateral inter-subject alignment. Cross-hemispheric RSA (y-axis) was computed between a parcel in one subject’s left hemisphere and the corresponding parcel in another subject’s right hemisphere, compared to bilateral RSA (x-axis). The near-perfect correlation ($r=0.99$) shows that combining hemispheres preserves geometry.
We quantified hemispheric asymmetry in representational geometry by computing alignment measures separately for the left and right hemispheres using the HCP-MMP1.0 atlas31. For each cortical parcel, we assessed (i) inter-subject alignment, (ii) brain–vision model alignment, and (iii) brain–language model alignment, using the same Natural Scenes Dataset (NSD)23 responses as in the main analyses (see Fig. A1).
Across all modalities, a largely symmetric pattern was observed (Fig. A1a–c): the rank-order of parcels was highly similar between hemispheres (Spearman’s $r$: inter-subject $=0.93$, vision $=0.92$, language $=0.82$), although some significant asymmetries were present. Specifically, early visual cortex (V1–V4) showed stronger alignment in the left hemisphere, while alignment was higher in the right hemisphere for ventral and LOTC hub parcels (hMT+, LO, TPOJ). Vision- and language-model alignment followed similar trends.
Grouping parcels by functional hub (Fig. A1d) highlighted these asymmetries: left hemisphere alignment was higher in early visual cortex, while right hemisphere alignment was higher in ventral and LOTC hubs.
To determine whether these asymmetries reflected differences in representational content or only magnitude, we compared cross-hemispheric RSA (e.g., participant $p$’s left V1 vs subject $q$’s right V1) to bilateral RSA (subject $p$’s left+right V1 vs subject $q$’s left+right V1). These measures were almost perfectly correlated across parcels ($r=0.99$, Fig. A1e), with no parcel deviating from the regression line. This demonstrates that combining hemispheres increases overall alignment strength without distorting the underlying representational geometry—no region showed a unique lateralized axis absent in the contralateral hemisphere.
Together, these results justify our approach of using symmetric (hemisphere-combined) parcels for the main analyses: (i) combining ROIs across hemispheres maximizes statistical power for multivariate comparisons without introducing distortions in representational geometry; (ii) hemisphere-separated ROIs remain informative for quantifying lateralization effects. This flexible approach supports both detailed lateralization analysis and robust, large-scale cross-dataset RSA experiments with joined hemispheres.
Figure A2: Repetition-matching strategies and their impact on inter-subject RSA. To compute inter-subject RSA, we compared different strategies for matching the three trial repetitions for each shared image between any pair of participants. (A) Schematic of unshifted (repetition 1→1, 2→2, 3→3) versus shifted (cyclic 1→2, 2→3, 3→1) trial matching with example RDMs. (B, C)Group-average cortical RSA maps under unshifted (B) and shifted (C) matching; both reveal the early-visual, ventral, and LOTC hubs. (D) Parcel-wise comparison of shifted vs. unshifted RSA values shows a uniform attenuation under shifted matching, with two linear regimes reflecting stronger prefrontal attenuation.
A.2 Repetition-Matching Strategies for Inter-Subject RSA
The Natural Scenes Dataset (NSD) presents 10,000 unique images across up to 40 scanning sessions per participant (30,000 trials total). A subset of 1,000 images (3,000 trials) was shared by all eight participants and always appeared in the same trial positions—interleaved among participant-unique stimuli—so that each shared image was viewed under identical practice, fatigue, and session-context conditions. This locked structure raises the possibility that inter-subject correlations might reflect non-perceptual factors (e.g., trial timing or memory demands) rather than purely stimulus-driven geometry.
To isolate the stimulus-driven component, we adopted a “shifted-repetition” matching strategy when computing inter-subject representational similarity (IS-RSA). Rather than pairing each repetition index $k\leftrightarrow k$ across participants (“unshifted”), we cyclically permute so that repetition $k$ in participant $p$ matches repetition $(k\bmod 3)+1$ in participant $q$ (i.e. $1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$; see Fig. A2A). Because all shared trials occupy the same session slots, this shift preserves image identity while breaking any exact repetition-locked confounds and also allows direct comparison with within-subject analyses (which must compare different repetitions; see Supplementary Fig. D5). Due to symmetry across participant-pair comparisons, shifts of 1 and 2 yield identical group-average maps.
Recomputing the group-average cortical RSA under both unshifted (Fig. A2B) and shifted (Fig. A2C) schemes recovers the same three hubs—early visual cortex, ventral occipito-temporal parcels, and LOTC. As expected, shifting is more conservative: alignment magnitudes are uniformly lower (Supplementary Fig. A2D), yet the relative rank-order of all parcels remains unchanged (Spearman’s $\rho=0.96$). Frontal parcels (e.g. FEF, inferior frontal gyrus) suffer proportionally larger attenuation under shifting than occipito-temporal hubs, suggesting that these frontal regions encode task- or repetition-related variance (e.g. decision or attentional control) that the shifted criterion minimizes. By contrast, the robust persistence of occipital and temporal alignments underscores their stable, stimulus-driven representational geometry.
To illustrate how these differences manifest in network structure, Fig. A3 shows representational connectivity graphs for both matching strategies. We first identified parcels whose inter-subject RSA exceeded an absolute threshold of $r=0.05$ in either the shifted (shift = 1 or 2) or unshifted maps. We then constructed a reduced connectivity matrix among this common set of nodes and extracted a two-iteration minimum-spanning-tree backbone (using edge weights defined as $1-\mathrm{RSA}$; see main Methods). The core two-stream topology (Early Visual $\rightarrow$ Ventral hub and Early Visual $\rightarrow$ LOTC hub) is preserved under both schemes. However, only the unshifted graph highlights FEF as a high-centrality node, further suggesting that frontal regions carry shared task- or repetition-locked signals that are attenuated by the shifted criterion.
While these frontal effects are exploratory—given their lower RSA magnitudes and sensitivity to thresholding—they point to important future work on how task and attention signals coexist with stimulus-driven representations. At the same time, they underscore the necessity of shifted-repetition controls and of complementary interpretability techniques to isolate which stimulus-content features drive convergence, enabling mechanistic hypotheses about information transmitted across cortical routes.
[…]
A.3 Token-occurrence control for language–brain RSA
To quantify how much of the language–brain alignment reflects initial tokenization, we constructed RDMs from token-occurrence vectors. Each caption was encoded as a count vector—each entry recording the number of times a given token appears—and pairwise Pearson dissimilarities among these vectors yielded a tokenizer-based RDM. We applied this procedure to six vocabularies (BLOOMZ, Gemma 2, LLaMA 1/2/3; see Table 1), covering all tokenizers used in our language models, and then computed parcel-wise RSA against NSD fMRI responses using the symmetric HCP atlas. Parcel-wise alignments were nearly identical across tokenizers (pairwise Spearman’s $\rho$ $\leq 0.978$), so we report their average.
The tokenizer–brain alignment map closely matches that of full language models: it peaks in the LOTC hub and shows minimal or negative alignment in early visual parcels (Fig.A4A,D). Token-count geometry correlates strongly with each model’s first-layer alignment (Pearson’s $\rho=0.78$; Fig.A4B) and with maximal alignment across layers ($\rho=0.68$; Fig. A4C), indicating that token frequencies account for a substantial portion of the language–brain correspondence.
Repeating the analysis with a word tokenizer—splitting captions on whitespace rather than into subword units—again yielded pronounced LOTC alignment (Fig. A4E). These findings suggest that LOTC’s alignment to language models largely reflects sensitivity to discrete lexical tokens common across descriptions—functioning as a “bag-of-words” representation—rather than to compositional or syntactic structure53.
Figure A4: Tokenizer-level contributions to language–brain alignment. We derived RDMs from token-occurrence count vectors over each model’s vocabulary and computed parcel-wise RSA against NSD fMRI responses (averaged across the six tokenizers used in our language models). (A) Scatter of tokenizer-brain vs. inter-subject alignment, with peak values in the LOTC hub. (B) Token-count alignment against each model’s first-layer alignment. (C) Token-count alignment against each model’s maximal alignment across layers. (D) Cortical map of tokenizer-brain alignment, showing pronounced LOTC localization. (E) Alignment map for a whitespace-based word tokenizer, which likewise peaks in LOTC, consistent with a “bag-of-words” explanation rather than richer compositional structure.
A.4 Robustness to Similarity Metric (Spearman-RSA & CKA)
Recent work has cautioned that representational-similarity findings can depend not only on stimulus design but also on the choice of similarity metric—some measures probe local, pairwise geometry while others emphasize global, space-wide dependence 59. To test our results’ robustness to these different measures, we repeated key analyses using two alternatives to Pearson-RSA: (i) a rank-based Spearman-RSA, which preserves only the ordering of each stimulus pair’s dissimilarity, and (ii) an unbiased Centered Kernel Alignment (CKA) that quantifies global statistical dependence via an unbiased HSIC estimator 92.
Replacing Pearson with Spearman correlations yields a purely rank-based, nonparametric measure agnostic to absolute distance magnitudes. As shown in Supplementary Fig. A5A–C, the inter-subject map, the parcel-by-parcel scatter against model alignment, and the representational connectivity matrix all mirror our original Pearson-RSA results (parcel-wise $r=0.99$). This confirms that neither the linearity nor distributional assumptions of Pearson correlation drive our findings—our local pairwise geometry holds under a fully nonparametric test.
CKA measures dependence across entire representational spaces, emphasizing alignment along principal axes rather than isolated pairwise distances. We computed linear-kernel CKA (using the unbiased HSIC estimator 92) under the same shifted-repetition protocol. Supplementary Fig. A5D–F shows that CKA again recovers the three hubs but inflates LOTC alignment and attenuates early visual alignment in the parcel scatter, reflecting CKA’s sensitivity to low-dimensional, dominant components. Parcel-wise CKA and Pearson-RSA still correlate strongly (Pearson’s $r=0.91$, Spearman’s $\rho=0.93$), and the CKA connectivity matrix preserves the two-stream topology, despite CKA’s unsigned nature preventing a sign-based “dissimilarity” distinction.
Across both local (Spearman-RSA) and global (CKA) criteria, the identification of Early Visual, Ventral, and LOTC hubs and their two-stream connectivity remains unchanged. This convergence across similarity measures, spanning local versus global criteria, strengthens confidence that our observed hierarchies and pathways reflect genuine stimulus-driven geometry rather than artifacts of any single metric.
Figure A5: Comparison of alignment metrics. (A–C) Spearman-RSA. Replacing Pearson’s $r$ with Spearman’s rank correlation yields virtually identical maps: the inter-subject alignment (A), the parcel-wise scatter against vision- and language-model alignment (B), and the representational connectivity matrix (C) all match the Pearson-RSA findings, with parcel-wise strengths correlating at $r=0.99$. (D–F) Unbiased CKA. Centered Kernel Alignment, which captures global dependence via an unbiased HSIC estimator, again highlights Early Visual, Ventral, and LOTC hubs (D). In the parcel-wise comparison (E), CKA inflates LOTC alignment and attenuates early visual alignment—reflecting its sensitivity to low-dimensional, dominant components—yet CKA and Pearson-RSA remain strongly correlated ($r=0.91$). The CKA-based connectivity matrix (F) preserves the two-stream topology, though as a dependence measure it does not distinguish signed similarity.
A.5 Extended details of Shared Component Decomposition and Partial RSA Controls
Figure A6: Extended analysis of representational dimensions. (A) KMCCA projections onto the top two canonical axes for six representative parcels: V1, V4, VMV2, PHA2, MT, and TPOJ2. V1 shows no clear clustering; V4 begins to separate by content; Ventral and LOTC parcels replicate the hub-level semantic gradients. (B) Words ranked by loading on the first KMCCA component (positive loadings in blue, negative in red) for each hub. Early Visual words share low-level shape features; Ventral words span scene vs. object terms; LOTC words distinguish animate vs. inanimate labels. (C) Partial RSA boxplots for the three hubs, showing inter-subject alignment under no control (baseline), controlling for the scene-object gradient, or biological presence. LOTC alignment is markedly reduced by each control, whereas Early Visual and Ventral hubs are minimally affected.
This supplementary note details our use of Kernel Multi-view Canonical Correlation Analysis (KMCCA)33 to isolate the axes driving our RSA findings32. While RSA provides a single global similarity score, KMCCA decomposes this correspondence into a ranked set of orthogonal “representational channels,” revealing which latent axes contribute most strongly to the shared geometry.
We performed KMCCA using the Python package mvlearn (v0.5.0)90 with a correlation kernel and a regularization parameter of $\lambda=0.3$, using default parameters of implementation. For each cortical parcel (or hub), we formed a data matrix for each participant from the GLM-denoised single-trial $\beta$-estimates. We included only those trials for which each (image, repetition) pair was available for all eight participants. This procedure yielded a set of eight matrices ($X_{p,r}\in\mathbb{R}^{m\times v_{p,r}}$), one for each participant, all representing an identical sequence of stimulus trials. The KMCCA algorithm was then applied to this set of matrices, returning projections onto a common shared subspace. For visualization (Fig. A6A; Fig 5A-C), we projected each subject’s data onto the first two canonical axes and then averaged the projections for the repetitions of each unique image to produce a single, stable point per stimulus.
The parcel-level KMCCA projections (Supplementary Fig.A6A) confirmed that the organizational principles observed at the hub level were consistent within their constituent parcels. In the Early Visual Cortex, the V1 parcel showed no clear semantic organization in its first two dimensions, whereas the V4 parcel already exhibited emergent clustering by image content. Within the Ventral hub, representative parcels such as VMV2 and PHA2 reproduced the hub-level scene-to-object gradient and showed clear organization by semantic category. Similarly, within the LOTC hub, parcels like MT and TPOJ2 showed the same primary separation between biological and non-biological stimuli. Within these main clusters, stimuli were further organized by more fine-grained semantic categories, mirroring the main hub-level analysis (Fig.5C).
To interpret these KMCCA axes semantically, we projected the image captions onto the shared subspace learned for each hub. We first created a vocabulary from the union of all words present in the captions ($s=11,090$ unique words), encoding each caption as a binary word-occurrence vector (a “bag-of-words” representation). We then projected this caption matrix onto the first KMCCA component derived for each hub, which yielded a loading weight for each word in the vocabulary. By ranking these weights, we identified the words that contributed most positively and negatively to each dimension (Fig. A6B). This word-loading analysis confirmed our interpretation of each hub’s organizing principle: in Early Visual Cortex, top-weighted words were not semantically coherent but appeared to share coarse visual features; in the Ventral hub, positive tokens index scene contexts (e.g., ‘kitchen’, ‘bathroom’) while negative tokens name isolated objects (e.g., ‘plate’, ‘frisbee’); and in the LOTC hub, tokens cleanly separated animate from inanimate labels. This analysis provides a semantic interpretation for the organization of each cluster, which can be visually confirmed in the corresponding image atlases (See Supplementary Figs. D8-D10)
Finally, we quantified the impact of the identified dimensions on the overall inter-subject alignment using partial RSA. Denoting by $\rho(X,Y)$ the standard Pearson correlation use to compute the alignment between two RDMs, $Z$ a third RDM (e.g., one built from the first KMCCA component or from a binary “biological presence” vector), the partial RSA was computed as:
$\mathrm{pRSA}(X,Y\mid Z)=\frac{\rho(X,Y)\;-\;\rho(X,Z)\,\rho(Y,Z)}{\sqrt{(1-\rho(X,Z)^{2})\,(1-\rho(Y,Z)^{2})}}.$ (A1)
As shown in the main text (Fig. 5d-f), controlling for the continuous, data-driven KMCCA axis produced only minimal drops in the Early Visual and Ventral hubs, but a substantial reduction in the LOTC hub. A complementary analysis confirmed this result: controlling for the discrete categorical variables yielded a drop in alignment of a similar magnitude to that produced by the continuous KMCCCA components (Supplementary Fig. A6C).
A.6 Power-Law Attenuation Model of RSA Metrics
In the main text, we compared several RSA measures—including within-subject (WS), inter-subject (IS-1: shift=1 and IS-0: unshifted), and vision-model-brain (VM)–that probed a common cortical geometry but were subject to different levels of measurement noise. While all metrics revealed a similar cortical alignment pattern, their magnitudes varied considerably. Parcel-wise scatter plots (e.g., Fig. 2E; Fig. A2) showed a systematic relationship between the metrics, which were tightly related by power-law curves. Although these curves differed in scale and curvature, their consistent form suggested that each metric captured the same latent similarity while imposing a unique, noise-dependent attenuation.
We formalise this observation by modelling the correlation measured for parcel r with metric j as a power-law transform of a latent correlation:
$\rho^{\mathrm{obs}}_{rj}=a_{j}\bigl{(}\rho^{\mathrm{latent}}_{r}\bigr{)}^{\,b_{j}}+\epsilon_{rj},$ (A2)
where
  • $\rho^{\mathrm{latent}}_{r}\in[0,1]$ is the metric-independent similarity common to all measures,
  • $a_{j}\in[0,1]$ is a linear attenuation factor for metric j, and
  • $b_{j}\geq 1$ captures signal-dependent (heteroscedastic) noise. A value of $b_{j}=1$ corresponds to simple homoscedastic attenuation, whereas $b_{j}>1$ creates a convex curve that suppresses weak latent correlations more strongly than strong ones. This constraint is motivated by our observation that the noisiest metrics (e.g., IS-shift) systematically underestimate weakly aligned parcels relative to cleaner metrics (e.g., WS), which is consistent with a noise regime that disproportionately affects weaker signals.
  • $\epsilon_{rj}$ is a residual term that absorbs any remaining parcel–metric–specific variability not captured by the deterministic power-law component.
We jointly fitted the parameters of Eq. A2 to four core metrics (WS, IS-0, IS-1, VM) using data from the symmetric-hemisphere ROIs. Parameter estimation was performed via the L-BFGS-B algorithm by minimizing a least-squares loss, with parameters bounded as follows: $a_{j}\in[0,1]$, $b_{j}\in[1,\infty)$, and $\rho^{\mathrm{true}}_{r}\in[0,1]$.
Figure A7A shows the model’s high goodness-of-fit by plotting the observed correlations against the model-fitted values for the four core metrics. In all cases, the model explained over 94% of the variance ($R^{2}>0.94$), confirming that the power-law attenuation model accurately predicts the trends in the observed data. Complementing this, Figure A7B visualizes the distinct attenuation profile for each RSA strategy, allowing for a direct observation of how each measured correlation is attenuated in comparison with the estimated latent correlation.
Because all metrics are modeled as functions of the same latent variable, $\rho^{\mathrm{true}}_{r}$, the relationship between any two metrics, j and j’, can be derived by eliminating this common term:
$y=\frac{a_{j^{\prime}}}{a_{j}^{\,b_{j^{\prime}}/b_{j}}}\;x^{\,b_{j^{\prime}}/b_{j}},$ (A3)
where $x=\rho^{\mathrm{obs}}_{rj}$ and $y=\rho^{\mathrm{obs}}_{rj^{\prime}}$. Figure A7C validates this prediction by overlaying the derived curve (Eq. A3) on the empirical parcel scatters. This confirmed that a single latent similarity, subject to metric-specific attenuation, explains all pairwise relations. For completeness, Figure A8 extends this validation to a set of 12 metrics, demonstrating the model’s power to unify all RSA strategies used in our analyses.
The resulting cortical map of the latent variable, $\rho^{\mathrm{latent}}_{r}$, is displayed in Figure A7D. This map represents a noise-corrected estimate of similarity that unifies the four core RSA measures. It reveals a three-hub pattern of high alignment observed in the main text, with peak correlations ($\rho^{\mathrm{latent}}_{r}\sim 0.50$) in the EVC, Ventral stream, and LOTC. Furthermore, the model shows that prefrontal areas (FEF and pFC) are more prominent after correcting for the attenuation present in the individual raw measurements. The fact that these hubs align neatly with those highlighted by individual metrics confirms that our model (Eq. A2) successfully isolates a common signal while factoring out metric-specific noise.
Because Eq. (A2) is scale-free, absolute values of $\hat{\rho}^{\mathrm{latent}}$ and $a_{j}$ are defined only up to a common factor. We fixed no anchor during optimisation, yet repeated fits converged to the same scale, and all qualitative findings (relative parcel ranking, cross-metric relations) are invariant under rescaling.
The power-law attenuation model offers a compact, quantitative account of how measurement differences degrades RSA scores across diverse metrics. Its success supports the view that all metrics probe a shared latent representational geometry, differing only in the reliability with which each metric samples that geometry.
Figure A7: (A) Goodness-of-fit. Parcel-wise observed RSA values (y-axis) are plotted against the corresponding predictions from the joint power-law model (Eq. A2; x-axis). Colours denote the different metrics (WS: within-subject; IS-1: inter-subject shifted (shift=1); IS-0: inter-subject unshifted; VM: vision model). The legend provides the explained variance ($R^{2}$) for each metric. (B) Correlation attenuation curves. The estimated latent correlation ($\hat{\rho}^{\mathrm{latent}}$, x-axis) is plotted against the observed correlation for each metric (y-axis). Coloured lines show the fitted metric-specific attenuation curves ($y=a_{j}\,x^{b_{j}}$). (C) Pairwise metric comparisons. Scatter plots for six representative metric pairs (grey points) are shown. Each plot includes an independent power-law fit to the empirical data (solid red line, with 95% CI) and the prediction derived from the joint model (dashed black line, Eq. A3), demonstrating high correspondence. (D) Latent similarity map. The cortical surface map displays the estimated latent correlation values, $\hat{\rho}^{\mathrm{latent}}_{r}$ (HCP-MMP1 atlas, symmetric parcels). The model reveals major hubs of high latent similarity in early visual cortex, ventral temporal cortex, and LOTC, with pFC becoming more prominent after accounting for metric-specific adjustment. Together, these panels demonstrate that a single latent geometry, subject to metric-specific power-law attenuation, effectively explains the relationships among all measured RSA metrics.
Figure A8: Global power-law model captures every pair-wise relation among RSA metrics. Scatter plots (grey points) show parcel-wise observed RSA values for all metric combinations: inter-subject (shifted and unshifted), within-subject, subject-versus-group, and vision-model correlations, each additionally split by hemisphere sampling (left, right, both). Red curves depict an independent power-law fit for the specific scatter, whereas black dashed curves are the predictions from a single joint fit of Eq. (A3) applied to all metrics simultaneously (no panel–specific re–fitting). The accompanying table reports the fitted scale factors $a_{j}$, exponents $b_{j}$, and explained variance $R^{2}$ for every metric (model-predicted vs observed rsa). All metric trends are explained with $R^{2}\!>\!0.92$, confirming that one latent similarity map, together with metric-specific attenuation parameters, suffices to account for the entire measurement set.
B THINGS-fMRI and BOLD5000 Processing
B.1 THINGS-fMRI dataset
We processed the THINGS-fMRI dataset 25 using a pipeline analogous to that for NSD. This dataset comprises 3T fMRI recordings from three participants performing an oddball detection task on 8,740 unique object images. A subset of 100 images was repeated once per session (12 presentations total); all other images appeared only once. For our primary computations, we used only the first presentation of each of the 8,740 images.
Pre-processed, ICA-regressed single-trial estimates were obtained from the authors’ public repository. We extracted voxel responses into the HCP-MMP atlas parcels provided by the THINGS team.
To control for within-session variability, inter-subject alignment was computed on a session-wise basis. For any pair of participants $(p,q)$ and parcels $(r,r^{\prime})$, we computed IS-RSA by correlating the RDMs generated from the images within a given session $s$:
$\mathcal{A}_{p,q,r,r^{\prime},s}^{\mathrm{IS}}=\mathrm{RSA}\bigl{(}X_{p,r,s},X_{q,r^{\prime},s}\bigr{)}.$ (B4)
Group-level maps were then obtained by averaging these scores across sessions and all participant pairs. Model-brain RSA followed the same session-wise procedure. As a control, we confirmed that computing IS-RSA over the full 8,740-trial set (using $,740\times 8$ 740 × 8740 RDMs) yielded parcel-wise results that were almost perfectly correlated with our session-wise computation (Pearson’s $r=0.995$). While the full-set computation produced systematically lower alignment values (linear slope $\approx 0.53$; $R^{2}=0.99$), this correlation validates that our session-wise approach preserves the relative spatial pattern of the findings.
B.2 BOLD5000 dataset
We processed the BOLD5000 dataset 24 using the same session-wise pipeline as for THINGS-fMRI. This dataset comprised 3T recordings from four participants performing a valence rating task on 4,916 unique images drawn from the MS-COCO, ImageNet, and SUN databases. We used the pre-processed, single-trial $\beta$-estimates provided by the authors, selecting only the first presentation of each image. Inter-subject and model-brain RSA were computed following the same session-wise procedure described above.
For voxel extraction, we registered the HCP-MMP1.0 atlas from MNI152 standard space to each subject’s native T1-weighted scan and then to their mean functional volume, using FSL’s FLIRT and FNIRT. While qualitative inspection confirmed consistent parcel definitions across participants, this volume-to-volume registration has a lower effective resolution compared to the specialized registration available for the NSD atlas. This may account for some parcel-specific results, such as the lower-than-expected alignment values observed in certain regions (e.g., MST). All other preprocessing and computational steps paralleled the main NSD pipeline.
C Implementation details for large-scale RSA comparisons
To efficiently handle the computational demands associated with large-scale RSA analyses (due to dataset size and permutation complexity), we developed a GPU-accelerated implementation using PyTorch84. Specifically, we reformulated representational dissimilarity matrix (RDM) calculations as optimized matrix operations. For each participant pair $p,q\in\mathcal{P}$, and each region $r\in\mathcal{R}$, we constructed stacked matrices consisting of flattened RDMs across all cortical regions, significantly reducing computational overhead and runtime compared to conventional RSA implementations.
Given data matrices for matched trials $X_{p,r}[\boldsymbol{i}_{pq},:]$ and $X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]$, we computed upper-triangular vectors of dissimilarities:
$\mathbf{x}^{(p)}_{pq,r}=\operatorname{vec}_{u}D\left(X_{p,r}[\boldsymbol{i}_{pq},:]\right),\quad\mathbf{x}^{(q)}_{pq,r^{\prime}}=\operatorname{vec}_{u}D\left(X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]\right),$ (C5)
where $D(\cdot)$ denotes the dissimilarity matrix and $\operatorname{vec}_{u}$ denotes vectorization of the upper triangle.
Each RDM vector was centered and $l_{2}$-normalized before stacking into a matrix containing all ROIs:
$\mathbf{X}^{(p)}_{pq}=\begin{bmatrix}\tilde{\mathbf{x}}^{(p)}_{pq,1}\\ \tilde{\mathbf{x}}^{(p)}_{pq,2}\\ \vdots\\ \tilde{\mathbf{x}}^{(p)}_{pq,|\mathcal{R}|}\end{bmatrix}\quad\in\mathbb{R}^{|\mathcal{R}|\times\frac{n(n-1)}{2}},$ (C6)
and similarly for $\mathbf{X}^{(q)}_{pq}$.
The inter-subject connectivity matrix was then obtained as:
$\mathcal{A}^{\mathrm{IS}}_{p,q}[r,r^{\prime}]=\mathbf{X}^{(p)}_{pq}\left(\mathbf{X}^{(q)}_{pq}\right)^{\top}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{R}|}$ (C7)
Analogous formulations were applied for Spearman-based RSA, with rank-normalization applied to RDM vectors prior to computing dot products.
For brain–model alignment, similar stacked matrices were constructed separately for brain RDMs and model RDMs across sessions. The brain matrix was organized as:
$\mathbf{X}^{(\text{brain})}_{p}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{T}_{p}|\times\frac{n(n-1)}{2}},$ (C8)
and the model matrix as:
$\mathbf{X}^{(\text{model})}_{m}\in\mathbb{R}^{|\mathcal{L}_{m}|\times|\mathcal{T}_{p}|\times\frac{n(n-1)}{2}}.$ (C9)
Brain–model RSA alignments across regions, layers, and sessions were computed as the inner product between normalized RDM vectors for each session:
$\mathcal{C}_{p,r,m,s}(l)=\langle\mathbf{X}^{(\text{brain})}_{p,r,s},\;\mathbf{X}^{(\text{model})}_{m,l,s}\rangle.$ (C10)
This operation was implemented as a contraction corresponding to the einsum notation $r\,s\,d,l\,s\,d\to r\,l$.
For permutation testing, we permuted stimulus labels defined by the permutation $\sigma\in S_{n}$ prior to RSA computation. Rather than recomputing dissimilarity matrices after each permutation, we constructed an equivalent permutation $\sigma^{\prime}$ over the vectorized upper-triangular RDM entries, matching the induced pairwise index shifts. This allowed permutation of the flattened RDMs directly via:
$\mathbf{X}^{(p)}_{pq}[:,\sigma^{\prime}],$ (C11)
preserving the correct stimulus-pair structure with minimal computational overhead.
D Extended figures
Figure D1: Detailed parcel-level alignment for all modalities. Box plots show representational alignment scores computed for each of the 180 cortical parcels of the symmetric HCP atlas 31, organized by macro-anatomical groups 47. Each box represents the distribution of alignment scores across the eight participants ($N=8$). The red line and shaded area in each panel denote the mean and standard deviation of the null distribution, respectively, estimated via permutation testing. (A) Inter-subject alignment (IS-RSA). (B) Brain-to-vision-model alignment. (C) Brain-to-language-model alignment. This figure provides a detailed view of the results summarized in the main text, showing the parcel-by-parcel variability and confirming the concentration of high alignment within the Early Visual, Ventral, and LOTC hubs.
Figure D2: Cortical flat map projections of representational alignment. To provide a comprehensive view of their spatial distribution, group-level alignment scores are projected onto a flattened cortical surface. The figure shows (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment. Values represent group-averaged RSA scores (Pearson’s $\rho$, $N=8$ from the NSD dataset), and major sulci are labeled for anatomical orientation. This visualization makes the full extent of the alignment patterns clear, particularly highlighting the widespread negative alignment (cool colors) between language models and the ventral visual stream, in contrast to the positive alignment seen for inter-subject and vision-model comparisons.
Figure D3: Detailed layer-wise alignment profiles across cortical parcels. This figure provides a detailed, parcel-by-parcel view of the brain-model alignment data summarized in the main text (Fig. 3), showing results for the 15 macro-anatomical groups with the highest inter-subject alignment. Only parcels that were themselves statistically significant in the inter-subject analysis are displayed. For each modality, the plotted alignment curves represent the RSA score averaged across all models of that type. Each individual line shows the mean alignment for a single parcel across participants ($N=8$), with shaded areas indicating the SEM. (A) Brain-to-vision-model alignment. This detailed view confirms that the distinct alignment profiles—decreasing, distributed, and increasing—are characteristic of the main hubs and often extend to anatomically adjacent parcels (e.g., the increasing profile of the LOTC hub is also observed in nearby STS regions). In contrast, areas with lower overall correspondence, such as in the prefrontal cortex, tend to show weak alignment across all model layers without a clear hierarchical preference. (B) Brain-to-language-model alignment. This view confirms that positive alignment is restricted to the LOTC hub and a small number of anatomically proximal parcels, all of which consistently exhibit the same step-like alignment profile.
Figure D4: Brain-model alignment profiles for different model families. Layer-wise alignment (RSA, Pearson’s $r$) is shown for representative parcels from the three main hubs, computed separately for each family of models (see Table 1 for a full list). Each curve represents the mean alignment across participants ($N=8$), with shaded areas indicating the SEM. Vision Models: Most vision model families (AugReg, CLIP, DINOv2) reproduce the characteristic hierarchical profiles: a decreasing alignment with Early Visual Cortex, a distributed alignment with the Ventral Hub, and an increasing alignment with the LOTC Hub. The Masked Autoencoder (MAE) models are a notable exception; consistent with their reconstructive training objective, they maintain a high alignment with Early Visual Cortex across all layers. Language Models: Despite differences in architecture, training data, and fine-tuning (e.g., instruction tuning in BLOOMZ), all language model families exhibit a similar pattern. Alignment is negligible or negative for Early Visual and Ventral hub parcels, while all families show the same characteristic step-function profile for parcels in the LOTC hub.
Figure D5: Comparison of within-subject and inter-subject representational analyses. (A) Cortical map of within-subject alignment (RSA, Pearson’s $r$). This was computed analogously to the inter-subject analysis by comparing RDMs from different trial repetitions within each participant, and then averaging across participants ($N=8$). The map reveals a similar spatial distribution to the inter-subject version, with a posterior-to-anterior gradient and three distinct hubs. (B) Parcel-wise comparison of within-subject and inter-subject alignment. The two measures are tightly correlated across the cortex, following a power-law relationship (fit shown in red; $R^{2}=0.98$). This indicates that brain regions that are consistent within an individual are also the ones that are consistent across individuals. (C) Within-subject representational connectivity matrix. Computed by correlating RDMs between all pairs of parcels within each participant, this analysis reveals the same three-hub block structure discovered using the inter-subject measurements. (D) Directed connectivity graph of the within-subject data, pruned using a 2-minimum spanning tree to highlight the network backbone. Edge color and width encode the inter-region alignment (RSA value), while node color represents each parcel’s peak alignment depth with vision models. Directionality is inferred from this depth measure based on a low-to-high sequential hierarchy, with arrows pointing from shallower- to deeper-aligning regions. The graph confirms the same information flow as the inter-subject analysis, with two streams emerging from early visual cortex: a medial-ventral stream to the Ventral hub and a lateral-dorsal stream to the LOTC hub.
Figure D6: Whole-cortex representational connectivity network. Representational connectivity network among the 157 cortical parcels that exhibited statistically significant inter-subject alignment. The network is pruned using a 2-minimum spanning tree to highlight the strongest connections. To minimize confounds from anatomical variability, connectivity was computed on a within-subject basis ($N=8$); for each participant, we correlated the RDM of every parcel with every other parcel across different repetitions of the same stimuli, then averaged the resulting connectivity matrices. Edge color encodes the strength of this inter-region alignment (RSA), while edge width indicates the spanning tree order (1st or 2nd) to highlight the most central pathways. Node color represents each parcel’s peak alignment depth with vision models, providing a proxy for its hierarchical position. Arrows are added to the pruned edges to denote a statistically significant difference in this depth between connected parcels (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$), indicating the putative direction of information flow. The graph confirms the two primary visual processing streams identified in the main text. A medial-ventral stream connects early visual areas to core ventral stream regions involved in scene and object processing. In parallel, a lateral-dorsal stream connects early visual areas to the LOTC hub, including motion-sensitive areas (MT+ complex) and higher-order regions in the superior temporal sulcus (STS) and temporoparietal junction (TPOJ). The analysis also reveals key bridges between these streams, with nodes such as LO3 and PGp linking the ventral and LOTC hubs. Furthermore, despite weaker overall alignment, frontal and parietal regions like the Frontal Eye Fields (FEF) and the Medial Intraparietal Area (MIP) emerge as distinct hubs, suggesting their integration into this large-scale, stimulus-driven network.
Figure D7: Biological content selectively drives LOTC alignment and lateral-stream connectivity. Scenes were split into those that contained biological agents (people or animals; 65.3% of the shared image set) and those that did not, and all inter-subject analyses were recomputed on each subset. (A) Parcel-wise inter-subject RSA within the three hubs. Blue: scenes with biological agents; orange: scenes without them. Early Visual parcels (V1–V4) show no reliable difference, Ventral parcels (VMV1–PHA3) align more for non-biological scenes, whereas LOTC parcels (V4t–TPOJ3) align almost exclusively for biological scenes. Stars mark two-tailed paired $t$-tests across eight participants (${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$, FDR-corrected). (B,C) Group-level representational-connectivity matrices (left) and backbone graphs (right) for scenes without (B) and with (C) biological agents. Backbones were extracted with a three-iteration minimum-spanning-tree procedure. Node colour encodes each parcel’s peak vision-model layer (early $\rightarrow$ late: green $\rightarrow$ purple), node size scales with within-parcel IS-RSA, edge hue shows pairwise RSA, and edge width denotes the iteration (1–3) at which the edge entered the spanning tree. The lateral stream linking Early Visual Cortex to LOTC is largely absent for non-biological scenes (B) but re-emerges when biological agents are present (C), mirroring the parcel-level results in (A).
Figure D8: Image-based atlas of the dominant representational dimensions in Early Visual Cortex. The first two shared representational components for the Early Visual Cortex hub, derived using KMCCA on the responses of all participants to the shared image set. Each point is an individual stimulus image, and its colored border indicates its general semantic category (see legend). The layout reveals no clear clustering by semantic category. Instead, images that are close to each other in this shared space often have similar low-level visual properties (e.g., color palettes, textures, or global shapes), confirming that the geometry of this hub is primarily driven by visual similarity rather than abstract content.
Figure D9: Image-based atlas of the dominant representational dimensions in the Ventral Hub. The first two shared representational components for the Ventral hub, derived using KMCCA. Each point is an individual stimulus image. The layout reveals an organizing principle: a smooth gradient from scene-dominated images to object-dominated images. For example, the upper-left quadrant contains images of indoor spaces, the lower portion contains outdoor scenes, and the upper-right quadrant contains images where a single object is the primary focus. This evidences that the dominant representational axis of this hub recapitulates the classic scene-to-object gradient of ventromedial cortex.
Figure D10: Image-based atlas of the dominant representational dimensions in the LOTC Hub. The first two shared representational components for the LOTC hub, derived using KMCCA. Each point is an individual stimulus image, with its border colored by its general semantic category (see legend). The first KMCCA component (the x-axis) separates images containing biological agents (left cluster) from those without (right cluster). Within the biological cluster, there is a further separation, with images of animals generally located to the right of the images of people. This confirms that the LOTC hub’s geometry is primarily organized around the presence and type of animate entities in a scene.

### v2

@@
-Convergent transformations of visual representation in brains and models
+Shared representations in brains and models reveal a two-route cortical organization during scene perception
 Abstract
-A fundamental question in cognitive neuroscience is what shapes visual perception: the external world’s structure or the brain’s internal architecture.
-Although some perceptual variability can be traced to individual differences, brain responses to naturalistic stimuli evoke similar activity patterns across individuals, suggesting a convergent representational principle.
-Here, we test if this stimulus-driven convergence follows a common trajectory across people and deep neural networks (DNNs) during its transformation from sensory to high-level internal representations.
-We introduce a unified framework that traces representational flow by combining inter-subject similarity with alignment to model hierarchies.
-Applying this framework to three independent fMRI datasets of visual scene perception, we reveal a cortex-wide network, conserved across individuals, organized into two pathways: a medial-ventral stream for scene structure and a lateral-dorsal stream tuned for social and biological content.
-This functional organization is captured by the hierarchies of vision DNNs but not language models, reinforcing the specificity of the visual-to-semantic transformation.
-These findings show a convergent computational solution for visual encoding in both human and artificial vision, driven by the structure of the external world.
+The brain transforms visual inputs into high-dimensional cortical representations that support diverse cognitive and behavioral goals.
+Characterizing how this information is organized and routed across the human brain is essential for understanding how we process complex visual scenes.
+Here, we applied representational similarity analysis to 7T fMRI data collected during natural scene viewing.
+We quantified representational geometry shared across individuals and compared it to hierarchical features from vision and language neural networks.
+This analysis revealed two distinct processing routes: a ventromedial pathway specialized for scene layout and environmental context, and a lateral occipitotemporal pathway selective for animate content.
+Vision models aligned with shared structure in both routes, whereas language models corresponded primarily with the lateral pathway.
+These findings refine classical visual-stream models by characterizing scene perception as a distributed cortical network with separable representational routes for context and animate content.
 1 Introduction
-Visual perception arises from the brain’s ability to internalize statistical regularities in the environment, transforming raw sensory input into structured, meaningful experience.
-Yet whether perception primarily reflects these external patterns or is shaped by the intrinsic computational architecture of the brain remains a central issue in cognitive science.
-This longstanding debate traces from philosophical perspectives like Plato’s allegory of the cave and theories of direct realism 1, which posit that perception is a direct reflection of the external world, to modern constructivist theories, which frame perception not as a passive imprint but as an active, model-based inference process 2, 3.
-A stimulus-driven view predicts that individuals will form convergent neural representations aligned with environmental structure4.
-Conversely, if internal constraints dominate, representations may diverge, reflecting idiosyncratic predictive models 5, 6, 7.
-Understanding the balance between these external and internal constraints is therefore foundational to any theory of cortical representation.
-Empirical evidence supports both sides of the debate.
-Substantial individual variability has been observed in low-level perceptual processes such as motion detection, contrast sensitivity, and color perception, linked to differences in cortical anatomy and function 8, 9, 10.
-In contrast, studies using naturalistic paradigms show that when different individuals process identical stimuli, they display similar activity patterns, indicating a shared representational organization across brains 11, 12, 13, 14.
-This suggests that despite anatomical and experiential variability, common transformation pathways reliably lead neural computation toward convergent representational outcomes.
-The principle of stimulus-driven convergence extends beyond the human brain to deep neural networks (DNNs).
-A substantial body of research has established a hierarchical correspondence between DNN layers and the processing stages of the visual system—a relationship that holds across diverse sensory modalities15, 16, 17, 18, 19.
-This alignment makes DNNs particularly effective probes for how sensory input is transformed into abstract, conceptual representations20, 21, 22.
-By providing an explicit, layer-by-layer mapping from low-level features to high-level semantics, DNNs offer a powerful model system for isolating the transformations that occur along the cortical hierarchy.
-Thus, characterizing a brain region’s alignment across the DNN hierarchy provides a functional fingerprint of its abstraction level and enables a data-driven approach for tracing information flow through the cortex.
-Here, we investigated whether the transformation of visual information into abstract meaning unfolds along a shared, systematic trajectory across individuals and whether this trajectory is reflected in the representational dynamics of DNNs.
-To address this, we introduce a unified, data-driven framework that combines inter-subject alignment to identify cortical areas and networks with shared geometry, model-based alignment to map the hierarchy, and decomposition techniques to interpret the stimuli features driving the correspondence (see Fig.
+Sensory inputs arrive at the brain as physical signals, but perception depends on how those signals are converted into structured patterns of cortical activity 37, 20.
+In vision, this conversion spans multiple stages, from early visual cortex to downstream regions whose activity patterns encode objects, scenes, and higher-order interactions 20, 35, 27, 77.
+Characterizing how these representations are transformed across the cortex is central to understanding how the brain performs this computation: identifying which regions encode specific aspects of the sensory input and mapping the networks that route this information between them.
+Naturalistic neuroimaging provides a way to approach this problem.
+When different individuals view the same stimuli, activity patterns show reliable synchrony across widespread regions 42, 5, 12.
+This inter-subject synchrony is interpreted as evidence that, despite individual variability, a substantial stimulus-locked component is encoded similarly across brains.
+This shared component can be used to localize regions that reliably encode stimulus-driven information across individuals 62, 29.
+Representational geometry formalizes this idea by describing each region in terms of pairwise similarities among stimulus-evoked responses, enabling comparisons across individuals and regions, and relating representational variance to stimulus properties that organize it 55, 43, 85.
+Representational geometry also enables direct comparisons between brains and deep neural networks (DNNs).
+Across vision and language models, internal activations can resemble cortical representations when both systems process the same stimuli 79, 10, 82, reflecting shared organizing dimensions that group stimuli according to similar coarse principles 13.
+In the visual domain, vision models show graded correspondence with cortical organization, with deeper layers capturing increasingly complex shape- and object-based features 51, 99, 17, 52.
+Complementarily, language models capture semantic regularities that abstract away from visual details, with recent work identifying correspondence in higher-level ventral and lateral regions where perceptual information interfaces with conceptual knowledge 72, 94.
+Comparing model-derived representations to brain responses therefore provides a compact way to situate regions along an ordered computational hierarchy and to distinguish visually grounded features from more abstract semantic representations 16, 23.
+Here, we investigated how information about visual scenes is encoded and routed across cortex by analyzing representational geometry shared among individuals and DNNs.
+We introduced a unified framework that links shared representational structure across individuals, its correspondence to vision and language models, and the network topology implied by representational connectivity, while using dimensionality decomposition to relate these patterns to coarse stimulus properties (Fig.
 1).
-We applied this framework to the massive-scale Natural Scenes Dataset (NSD) 23 and further validated the generalization of our findings in two independent fMRI datasets–BOLD5000 24 and THINGS-fMRI 25–with different image-viewing tasks.
-Our analysis applied to scene perception revealed a cortex-wide representational network that is shared across individuals and organized into two principal streams with distinct functional roles.
-One stream follows a medial-ventral trajectory dedicated to encoding visually grounded scene structure, aligning with the classical ventral pathway 26, 27, 28.
-The other traces a lateral-dorsal route, selectively tuned to biological and social content, offering empirical support and a data-driven refinement for recent proposals of a ‘third visual pathway’ for social perception 29, 30.
-This convergence across brains and artificial systems provides a principled framework for mapping how stimulus-driven information propagates through cortical systems, revealing the functional architecture that supports the transformation of sensory information into abstract representations.
-Figure 1: A unified framework for tracing representational pathways.
+Integrating these complementary approaches, which are often used in isolation, allowed us to move beyond localizing regions with shared information to characterizing how these regions organize as a distributed network and identifying the specific stimulus features that drive this organization.
+We applied this framework to the Natural Scenes Dataset (NSD) 2, in which participants performed a memory recognition task while viewing thousands of natural scenes spanning diverse environments, objects, and social interactions.
+We recovered a cortex-wide organization with two dissociable processing routes, each specialized to represent distinct types of stimulus information.
+The medial-ventral stream was specialized for scene layout and environmental context, while the lateral stream was selectively tuned to animate and social content.
+DNNs captured the shared geometry of both routes, though with clear modality-specific differences: vision models aligned with both pathways, whereas language models corresponded primarily with the lateral stream.
+These findings are consistent with recent refinements of classical visual-system accounts 35, 54, including proposals of a ventromedial route for scene context and a lateral “third visual pathway” for social perception 78, 70.We then compared these results using two independent datasets (BOLD5000 11 and THINGS-fMRI 46) and found that, although the same cortical areas consistently emerged in the alignment analysis, replicating the full two-stream organization required rich scene and social content.
+Figure 1: Overview of the analysis pipeline.
 (A) Feature extraction.
-For each image stimulus, we extracted equivalent representations from brain activity and deep neural networks.
-Single-trial fMRI responses were aggregated within HCP-MMP cortical parcels 31 to create vector representations of the brain’s response.
+For each image stimulus, we extracted corresponding representations from brain activity and deep neural networks.
+Single-trial fMRI responses were aggregated within cortical parcels to create vector representations of the brain’s response.
 Concurrently, layer-wise activations were extracted from pre-trained vision and language models to obtain representations across the full model hierarchies.
-Both brain and model vectors were used to compute Representational Dissimilarity Matrices (RDMs) based on their pairwise correlations.
+Both brain and model vectors were used to compute representational dissimilarity matrices (RDMs).
+Example input image adapted from Wikimedia (Bengt Nyman, $C$ BY 2.0).
 (B) Representational alignment.
-Representational Similarity Analysis (RSA) 32 was used to compare RDMs, quantifying the alignment between brain parcels and model layers.
+Representational Similarity Analysis (RSA) was used in two ways: (i) inter-subject RSA (IS-RSA), correlating parcel-wise RDMs across participants to estimate shared representational geometry; and (ii) brain–model RSA, correlating parcel RDMs with model-layer RDMs to quantify brain–model and layer-wise alignment profiles.
 (C) Representational connectivity.
-Inter-parcel RSA was used to construct a cortical network based on shared representational geometry.
-Analyzing this network revealed the hierarchical flow of information and identified key representational hubs.
+IS-RSA between pairs of parcels was used to construct a cortical network based on how similarly regions encode the stimulus set.
+The directionality of information flow was inferred from the peak model-layer alignment established in (B).
 (D) Shared dimensions.
-Within these hubs, Kernel Multi-view Canonical Correlation Analysis (KMCCA) 33 was used to isolate a subspace of representational dimensions shared across all participants.
-This final step links the data-driven representational axes to specific properties of the scene content, allowing for a functional interpretation of the information driving the alignment across the network.
+Within the main hubs identified in (C), Kernel Multi-view Canonical Correlation Analysis (KMCCA) was used to decompose the shared geometry into latent dimensions common across participants and to relate these dimensions to coarse scene features that drive alignment in different parts of the network.
 2 Results
-2.1 Inter-subject convergence in stimulus representation
-A central prediction of stimulus-driven perception is that different individuals processing the same stimuli should converge on a common representational geometry 13, 32.
-For the natural image-viewing tasks studied here, this implies that stimulus structure should drive representational alignment across observers in relevant cortical areas–from early visual cortex through high-level visual regions, and potentially into frontal areas implicated in visually guided decisions 34, 35.
-While inter-subject measures in naturalistic fMRI are often interpreted as ceiling estimates of signal-to-noise 23, 36, 37, we instead treated high inter-subject RSA (IS-RSA) as direct evidence that a region encodes stimulus information in a format shared across individuals.
-Identifying this cortical map of shared representational geometry thus provides a foundation for characterizing the spatial organization of stimulus-driven representations and serves as a starting point for examining how visual information is transformed across distinct brain regions.
-To evaluate this prediction, we focused our analysis on the Natural Scenes Dataset (NSD) 23.
-This dataset comprises 7T fMRI recordings in which eight participants viewed thousands of naturalistic scenes images in a memory task, each presented up to three times across sessions spanning a full year.
-For each stimulus, we formed a beta vector for each HCP-MMP parcel by concatenating the single-trial responses of its voxels 31.
-We then computed a parcel-wise Representational Dissimilarity Matrix (RDM) for each subject and used Representational Similarity Analysis (RSA) to quantify the geometry alignment of these neural patterns 32.
-To isolate the stable, perception-driven component of the response from potential confounds like session timing, we computed inter-subject alignment (IS-RSA) using a ‘shifted-repetition’ matching.
-This method compares RDMs between subjects using responses to the same images but from different trial repetitions (e.g., the first repetition for one subject vs.
-the second for another), ensuring alignment is driven by stimulus content alone (see Supplementary Fig.
-A2 for a schematic; main pipeline in Fig.
-1).
-Statistical significance was then assessed against a stimulus-shuffled null distribution (10,000 repetitions) to identify parcels with a significant shared geometry (see Methods).
-IS-RSA captured significant inter-subject alignment throughout the visual hierarchy (Figs.
-2a–b; see Supplementary Fig.
-D1 for full parcel-level statistics).
-We observed an alignment gradient that peaked in early visual cortex and extended into high-level ventral occipito-temporal and dorsal occipito-parietal regions–the main areas forming the visual system hierarchy34, 38, 35.
-These regions are also linked to high-level functions like scene perception and place memory 27, 39.
-Weaker, yet significant, alignment was also found in prefrontal cortex (PFC) areas typically linked to visual attention and task-dependent semantic processing and anterior areas associated to 40, 41.
-This spatial pattern is consistent with prior analyses of the NSD dataset using signal-to-noise and decoding methods23, 42, 43.
-Although some hemispheric asymmetries in alignment strength were present (Fig.
-2b), the representational profiles of symmetric parcels in the left and right hemispheres were highly correlated, suggesting that both hemispheres encode a similar geometric structure (Supplementary Fig.
-A1).
-These results support the hypothesis that scene statistics impose a shared representational organization from low-level visual areas to higher-order association cortex.
-Three anatomically clustered groups of parcels showed a consistent alignment across participants (Fig.
+2.1 Inter-subject shared geometry during scene viewing
+To localize cortical regions whose response patterns reflect reliable stimulus information in a shared format across people, we used inter-subject representational similarity analysis (IS-RSA) 29, 55.
+We applied this approach to the Natural Scenes Dataset (NSD) 2, which provides high-resolution 7T fMRI responses from eight participants performing an inter-session long-term recognition memory task in which they detected repeated images while viewing thousands of natural scenes.
+Our goal was to measure the stimulus-driven representational geometry related to visual content.
+To isolate this geometry, we parcellated the cortical surface into 180 regions per hemisphere using the Human Connectome Project multimodal parcellation (HCP-MMP) atlas 34 and extracted the provided single-trial responses for each image.
+For each subject and parcel, we constructed representational dissimilarity matrices (RDMs) from pairwise distances between multivoxel patterns.
+We then quantified inter-subject alignment by correlating parcel-wise RDMs across participants, using a shifted-repetition scheme to reduce mnemonic and session-related confounds (Supplementary Note 1).
+This yielded a parcel-wise map of shared representational geometry.
+IS-RSA captured significant alignment across occipital cortex and extended into higher-order association cortex (Fig.
+2a; Supplementary Figs.
+1-2).
+Alignment was strongest in early visual cortex (V1-V4) and extended into ventral occipitotemporal and dorsal occipitoparietal regions, consistent with hierarchical accounts of visual organization 20, 54, 28.
+We also observed significant, albeit weaker, alignment in prefrontal regions, which may reflect shared contributions of visual attention and task-dependent semantic processing30, 95, 8.
+Within this map, three anatomically clustered sets of parcels showed especially strong and consistent alignment across participants (Fig.
 2b).
-The first is early visual cortex (V1–V4), region associated with processing low-level visual features 44.
-The second is a ventral hub spanning key regions for processing scene context and spatial layout (parcels: VMV1-3, PHA1-3) 27, 45, 28.
-The third is a dorsal hub in the lateral occipitotemporal cortex (LOTC), encompassing the MT+ complex and the temporoparietal junction (parcels: MT, MST, FST, V4t, TPOJ2-3).
-These regions, which constitute an extension of the dorsal stream beyond classic spatial processing, are typically linked to motion integration, social cognition, and multimodal processing 12, 30, 46, 39.
-In the following sections, we use these three hubs—early visual, ventral, and LOTC—as reference points for our subsequent analyses.
-Figure 2: Cortical distribution of representational alignment.
-(A) Inter-subject alignment (RSA, Pearson $r$) computed per parcel and grouped by macro-anatomical clusters 47.
-Box-plots show the ten clusters with the highest mean alignment in the NSD sample ($N=8$; symmetric HCP-MMP atlas 31).
-Red lines mark the parcel-wise null distribution (mean ± s.d.; 10 000 label permutations).
-All displayed parcels exceed chance (two-tailed, FDR-corrected) at $p<0.001$, except EC and Pfm ($p<0.01$), 31pd ($p<0.05$) and 23d/31pv (n.s.).
-Peak clusters include early visual cortex (V1–V4), the ventral hub (VMV1–3, PHA1–3) and the LOTC hub (V4t, MT, MST, FST, TPOJ2–3).
-Full-atlas results appear in Supplementary Fig.
-D1.
-(B–D) Surface maps of inter-subject (B), vision-model (C) and language-model (D) alignment.
-RSA values are averaged across subjects and—where applicable—across models.
-(E–F) Parcel-wise relation between inter-subject alignment and model–brain alignment for vision (E) and language (F).
-Vision parcels follow a power-law fit ($R^{2}=0.94$, 95% bootstrap CI); language parcels cluster near zero except for those in the LOTC hub.
-(G) Mean alignment by modality within the three hubs.
-Paired $t$-tests ($t(7)$, two-tailed, FDR-corrected) show higher vision-model alignment in early-visual and ventral hubs, and higher language-model alignment in the LOTC hub (all $p<0.0001$).
-2.2 Model–Brain Alignment Mirrors Inter-Subject Shared Geometry
-Our previous analysis revealed a shared representational geometry across individuals; however, the nature of the information underlying this alignment remains unspecified.
-To determine whether this shared geometry is driven by low-level visual features or by more abstract, conceptual content, we compared the neural data to two distinct classes of computational models.
-Deep vision models, which learn representations directly from pixel data, provide a template for a visually-grounded processing hierarchy48, 49, 50.
-In contrast, large language models, which learn from text tokens, provide a template for modality-independent semantic and conceptual relationships.
-By systematically comparing cortical activity to both model types, we can therefore functionally characterize brain regions.
-This approach allows us to determine whether their organizing principles are primarily visual or more abstract in nature.
-We analyzed a diverse collection of vision and language models based on Transformer and Vision Transformer (ViT) architectures, spanning supervised, self-supervised, and contrastive training objectives (see Table 1).
-For each model, we computed the layer-wise RSA between its layer activations and every HCP-MMP parcel (Fig.
-1; Methods).
-To obtain a single summary measure of model-parcel correspondence, we then identified the maximum RSA value across all layers.
-This max-RSA value estimates the peak alignment achieved between a model and a brain region, irrespective of the specific hierarchical stage at which it occurs.
-This approach yields a cortex-wide alignment map directly comparable to the IS-RSA map (Fig.
-2c–d).
-Statistical significance was assessed via a two-tailed permutation test (10,000 permutations; see Supplementary Fig.
-D1 for full atlas results).
-Vision-trained models closely reproduced the inter-subject map.
-Across HCP parcels, a parcel’s vision-model alignment score scaled tightly with its IS-RSA value, following a power-law relationship ($R^{2}=0.94$;Fig.
-2e; see Supl.
-analysis A.6 for details).
-Spatially, the cortical alignment map (Fig.
-2c) mirrored the inter-subject pattern (Fig.
-2b), with the highest alignment in the Early Visual, Ventral, and LOTC hubs, and significant, albeit weaker, alignment throughout the remaining visual areas and prefrontal cortex.
-This tight correspondence suggests that the cortical parcels where stimulus geometry is shared across individuals are precisely those whose representational structure is captured by image-trained networks, indicating that these models and the human brain converge on a common, stimulus-driven coding scheme for natural scenes.
-Language models trained on image captions displayed a distinct and more localized alignment pattern.
-The cortical alignment map showed no strong alignment in early visual or ventral areas.
-Specifically, the classical visual streams exhibited negative RSA scores, indicating dissimilar representational geometries that contrast with visually grounded organizations (Fig.2d; Supplementary Fig.D2).
-In contrast, robust positive alignment emerged in the LOTC hub and neighboring parcels (Fig.2f), indicating that alignment in these regions is not visually grounded.
-A direct comparison of modalities revealed that language-model alignment exceeded vision-model alignment within this LOTC hub (Fig.2g).
-These results serve as a modality control, indicating that while alignment in early and ventral areas is tied to visually grounded features, the information shared in the LOTC hub is organized around a non-visual structure captured by language embeddings of image descriptions.
-Beyond this modality dissociation, we next aimed to situate cortical regions along the visual-to-semantic transformation trajectory.
-To this end, we examined their layer-wise alignment profiles, using the model hierarchy as a proxy for the brain’s own computational stages.
-Figure 3: Hierarchical convergence between models and cortex.
-(A–C) Layer-wise alignment (RSA) between vision models and brain activity for representative parcels within the three reference hubs.
-The alignment curves are averaged across all vision models used in the study.
+The first cluster corresponds to the early visual cortex (V1-V4), consistent with its role in encoding low-level visual features 93.
+The second comprised a ventral hub in ventromedial temporal cortex (VMV1-3, PHA1-3), overlapping scene-selective regions such as the parahippocampal place area (PPA) and adjacent cortex implicated in encoding scene layout and place memory 27, 77, 66.
+The third comprised a lateral occipitotemporal cortex (LOTC) hub including the motion-selective MT+ complex and posterior temporoparietal junction (MT, MST, FST, V4t, TPOJ2-3), near regions responsive to bodies and biological motion that have been implicated in social perception 70, 95, 6.
+In the following sections, we use these three hubs (early visual, ventral, and LOTC) as reference points for subsequent analyses.
+This spatial distribution is consistent with prior NSD work reporting variability in signal reliability and decodability across cortex 2, 87, 18, 22.
+Across parcels, higher IS-RSA values were associated with higher signal reliability (Supplementary Note 2), and a within-subject RSA analysis comparing different repetitions within each participant produced a similar cortical pattern (Supplementary Fig.
+3).
+Although alignment strength showed some hemispheric asymmetries (Fig.
+2b), representational profiles of left-right homologous parcels were highly correlated, consistent with a largely bilateral organization (Supplementary Note 3).
+Together, these results indicate that IS-RSA isolates a distributed set of regions whose scene-evoked representational geometry is shared across individuals, providing a benchmark for model comparisons described in the next sections.
+Figure 2: Inter-subject and model–brain alignment across cortex.
+(A) Inter-subject representational alignment (IS-RSA; Pearson’s $r$ between parcel-wise RDMs across participants) for parcels drawn from the macro-anatomical clusters with the highest mean IS-RSA (color legend).
+Boxplots show the distribution across subjects in the NSD sample ($N=8$, symmetric HCP–MMP atlas).
+Red lines indicate the parcel-wise null distribution (mean $\pm$ s.d.; 10,000 permutations).
+Unless otherwise annotated in the plot, parcels are significant at *** ($p<0.001$; two-tailed, FDR-corrected); additional annotations indicate ** ($p<0.01$), * ($p<0.05$), or n.s.
+($p\geq 0.05$).
+Peak alignment occurs in early visual cortex (V1–V4), a ventral hub (VMV1–3, PHA1–3), and a lateral occipitotemporal (LOTC) hub (V4t, MT, MST, FST, TPOJ2–3).
+Full-atlas results are shown in Supplementary Fig.
+1.
+(B–D) Cortical surface maps of alignment, averaged across subjects (and across models within each modality): (B) inter-subject IS-RSA; (C) vision–model RSA (maximum across layers, averaged across vision models); (D) language–model RSA (maximum across layers, averaged across language models).
+(E–F) Parcel-wise relationship between IS-RSA and model–brain alignment for (E) vision models and (F) language models.
+Points are colored by macro-anatomical group (as in panel A); other parcels are shown in white.
+Vision parcels follow an approximate power-law fit ($R^{2}=0.94$, shaded band: 95% bootstrap CI), whereas language-model alignment clusters near zero or negative except for parcels in and around the LOTC hub.
+(G) Modality comparison within the three hubs.
+Boxplots show hub alignment for vision and language models in early visual cortex, the ventral hub, and the LOTC hub (averaged across models).
+Paired two-tailed $t$-tests ($t(7)$, FDR-corrected) indicate stronger vision-model alignment in early visual and ventral hubs, and stronger language-model alignment in the LOTC hub.
+2.2 Model–brain alignment reveals modality-specific shared geometry
+Although IS-RSA localized parcels whose representational geometry was shared across observers, it did not by itself specify which stimulus information underlay that shared structure.
+To distinguish visually grounded from more abstract features, we compared cortical RDMs with RDMs from pretrained vision and language models.
+Vision models provide hierarchies learned directly from pixels 101, 56, 98, whereas large language models (LLMs) provide a reference for text-derived structure.
+By contrasting brain alignment with these two model classes, we characterized whether a parcel’s shared geometry was better captured by visually grounded features or by higher-level language representations.
+We analyzed 37 pretrained models based on Vision Transformer (ViT) 24 and Transformer 89 architectures, spanning supervised, self-supervised, and contrastive objectives (Table 1).
+For each model, we extracted layer-wise activations for each NSD image (vision models) or its caption (language models), constructed layer-specific RDMs, and quantified model-brain alignment by correlating each layer RDM with each parcel RDM.
+For each model, parcel-wise alignment was summarized as the maximum RSA across layers.
+Fig.
+2c-d show per-parcel maps averaged across subjects and, within each modality, across models (significance was assessed with a permutation-based null distribution).
+Vision models produced an alignment map that closely matched the IS-RSA spatial pattern (Fig.
+2c vs.
+Fig.
+2b).
+Parcels with high IS-RSA also showed strong vision-model alignment, with peaks in the Early Visual, Ventral, and LOTC hubs and weaker but significant effects across additional visual and prefrontal parcels.
+This overlap indicates that image-trained feature spaces capture a substantial component of the stimulus-structured geometry that is shared across individuals during scene viewing, consistent with prior NSD results and ventral-stream model-brain alignment work 51, 99, 17, 94, 2, 87.
+Across parcels, IS-RSA and vision-model alignment were strongly related by approximate power-law scaling ($R^{2}=0.94$; Fig.
+2e), consistent with both measures being largely driven by a common stimulus-locked component of scene organization (see Supplementary Note 4 for analysis of scaling and reliability).
+In contrast, autoregressive LLMs trained on next-token prediction showed alignment concentrated in LOTC and adjacent temporal cortex, and more weakly in anterior areas of visual cortex (Fig.2d,f), regions implicated in biological motion, social perception, and multimodal processing 72, 6, 70, 95.
+Alignment in these areas is consistent with prior reports of language-model correspondence in lateral/anterior temporal cortex and medial frontal regions in naturalistic paradigms associated with semantic processing 72, 94, 22.
+Within the LOTC hub, language-model alignment exceeded vision-model alignment (Fig.2g), indicating that the shared geometry expressed in this hub is better captured by caption-derived structure than by purely visual model features.
+Conversely, early and ventral visual regions showed near-zero or negative RSA with these LLM embeddings (Fig.
+2d; Supplementary Fig.
+2).
+Negative RSA indicates a systematic mismatch in geometry: stimulus features that separate responses in these parcels, mainly dominated by visual characteristics, are not emphasized by the language-model representations.
+This lack of alignment contrasts with studies finding ventral correspondence using language models optimized for sentence similarity or language-vision contrastive objectives 94, 22.
+Unlike the next-token prediction models used here, similarity-based models organize embeddings in their output layer by semantic content–grouping descriptions of similar scenes–which results in geometries that overlap more substantially with high-level visual representations.
+Figure 3: Hierarchical correspondence between model layers and cortical regions.
+(A–C) Layer-wise alignment (RSA) between vision models and representative parcels in the three hubs, averaged across all vision models.
 Panels show early visual cortex (A), the Ventral hub (B), and the LOTC hub (C).
-Lines show the mean alignment across participants ($N=8$), with shaded areas representing the standard error of the mean (SEM).
-Alignment profiles differ by hub, showing early, distributed, and late-stage peaks, respectively.
-(D) Cortical surface map of the vision model layer with the highest alignment for each parcel.
-The color scale indicates the normalized depth of the peak layer (0% shallower layer, 100% deeper layer), revealing a posterior-to-anterior gradient from shallow (green) to deep (purple) layers.
-(E) Layer-wise alignment between language model and representative parcels from each of the three hubs.
-In contrast to vision models, only parcels within the LOTC hub show strong alignment, which follows a step-like function rather than a smooth progression.
-(F) Box plots showing the distribution of peak alignment depths for the 20 parcels with the highest overall vision-model alignment.
-Parcels are sorted by their median peak depth, illustrating a continuous hierarchical organization spanning from early visual areas to the LOTC hub.
+The $x$-axis expresses depth as a percentage of the total number of layers in each vision model.
+Curves show the mean alignment across participants ($N=8$); shaded bands indicate the standard error of the mean (SEM).
+Early visual parcels peak in shallow layers, Ventral parcels show distributed alignment across the hierarchy, and LOTC parcels peak in the deepest layers.
+(D) Cortical surface map of the vision-model layer that yields the highest alignment for each parcel (averaged across vision models).
+Colors encode normalized peak-layer depth (0% = shallowest layer, 100% = deepest layer), revealing a posterior–anterior gradient from shallow (green) to deep (purple) alignment.
+(E) Layer-wise alignment between language models and representative parcels from the three hubs (mean across language models).
+Alignment values are negligible in early visual and Ventral parcels and are restricted to LOTC, where curves rise quickly and then form a plateau rather than a smooth hierarchical progression.
+(F) Distribution across participants ($N=8$) of peak alignment depths for the 20 parcels with the highest overall vision–model alignment.
+For each subject and parcel, peak depth is defined as the mean depth (across vision models) of the layer at which RSA is maximal.
+Boxplots are ordered by median peak depth and colored by macro-anatomical cluster.
 2.3 Hierarchical correspondence between models and cortex
-To obtain a more precise mapping of how brain regions correspond to different levels of model abstraction, we moved beyond single summary score and examined the layer-wise alignment profiles for each cortical parcel.
-While numerous studies have established a coarse hierarchical correspondence showing that early model layers align with V1–V4 and deeper layers map onto downstream regions 15, 16, 17, 51, 52, it remains unclear whether this is a simple one-to-one mapping.
-Alternatively, certain regions might exhibit a mixed alignment by integrating features from multiple computational stages.
-To address this, we conducted a layer-wise RSA for each parcel, yielding an alignment profile across model depth (see Methods; Fig.
-3).
-This layer-wise analysis of vision models showed distinct alignment profiles for each of the three main cortical hubs (Fig.3a–c).
-Alignment with the early visual cortex peaked in the shallowest model layers before decreasing, consistent with its role in processing low-level features (Fig.3a).
-In contrast, the ventral hub exhibited a broad plateau of moderate alignment across many layers, suggesting it integrates both visual features from earlier stages and more abstract information from deeper stages (Fig.3b).
-The LOTC hub, in turn, showed a progressively increasing alignment that peaked only in the deepest, most semantic layers (Fig.3c; see Supplementary Figs.
-D3D4 for extended parcels and model-family details).
-To make the functional hierarchy explicit, we mapped the peak alignment layer of vision models for each parcel.
-Despite the mixed correspondence observed in some ventral regions, this analysis revealed a smooth posterior-to-anterior gradient across the visual system (Fig.
+To relate cortical parcels to the ordered representational stages learned by DNNs, we examined how model-brain alignment varied across model depth.
+Prior work has shown that, for vision models, shallow layers tend to align with early visual cortex and deeper layers tend to align with higher-level ventral and downstream regions 79, 51, 99, 17, 38.
+Here, we assessed this depth dependence during natural scene viewing in NSD by analyzing layer-wise RSA curves, focusing on the three hubs where inter-subject shared geometry was strongest (Fig.
+3; full parcel-wise curves and model-family profiles in Supplementary Figs.
+4-5).
+The three hubs showed distinct depth profiles for vision models (Fig.
+3a-c).
+Early visual parcels peaked in the shallowest layers and then decreased (Fig.
+3a), consistent with correspondence to lower-level visual features.
+Ventral hub parcels showed broad alignment spanning earlier and later layers (Fig.
+3b), indicating that their shared geometry relates to features distributed across multiple stages of the vision-model hierarchy rather than a single best-matching depth.
+LOTC parcels increased more steadily with depth and reached their maxima in deeper layers (Fig.
+3c), indicating that the stimulus organization expressed in LOTC is best matched by later-stage vision-model features.
+To summarize this organization across cortex, we identified for each parcel the vision-model layer with maximal alignment and projected this peak depth onto the cortical surface.
+The resulting map formed a posterior-to-anterior gradient in which early visual cortex was best explained by shallower layers, while more anterior occipitotemporal, STS-adjacent, and prefrontal regions were best explained by deeper layers (Fig.
 3d).
-This gradient was most apparent within visual cortex, whereas frontal areas exhibited uniformly flattened alignment profiles, reflecting mixed correspondence with features spanning multiple model stages (see Supplementary Fig.
-D3).
-Nevertheless, the ordered progression from low-level to high-level representations was highly consistent across participants in visual areas (Fig.3f), providing evidence for a shared progression toward abstraction mirrored between cortex and models.
-Language models, in contrast, showed no cortex-wide progression; strong alignment emerged only in the LOTC hub (Fig.
-3e).
-Elsewhere, values hovered near zero across layers (Supplementary Fig.
-D3).
-Within LOTC, the alignment curve followed a step-like profile: it began with a weak score that was highly correlated with the token-level embeddings fed into the first layer (Pearson’s $r=0.78$; tokenizer control, Supplementary Fig.
-A4), consistent with work that suggest that brain–language alignment may reflect discrete word properties rather than compositional meaning 53.
-After this initial stage, alignment increased sharply before reaching a stable maximum, with the final value proportional to the first-layer alignment (Pearson’s $r=0.97$).
-This pattern indicates that LOTC representations are not late-stage vision, but are instead organized around semantic dimensions driven by token representations in language, with alignment increasing once language models have formed stable embedding representations.
-Although these analyses showed that some parcels exhibited mixed correspondence with features from different model stages, they revealed a cortical processing hierarchy that was partially mirrored in the architecture of deep vision models.
-The distinct, non-hierarchical profile observed for language models further highlighted this structure, isolating the LOTC hub as a region organized around a different, non-visual representational principle.
-Having established the vision model hierarchy as a proxy for the brain’s own computational stages, we proceeded to examine how information is routed between cortical regions.
+This gradient is consistent with hierarchical accounts of visual processing 28, 20, 54, 77 and with prior work showing that model depth tracks representational complexity across the human ventral pathway 38, 99, 98, 51, 17, 79.
+We use peak depth as a descriptive index that orders parcels along the model hierarchy (Fig.
+3f).
+Language models showed a different pattern.
+Alignment was concentrated in the LOTC hub, and layer-wise profiles did not show the early-to-late progression observed for vision models.
+Instead, alignment was present at the first layer, increased, and then approached a plateau.
+Across the language models we evaluated, first-layer alignment strongly predicted the plateau magnitude (Pearson’s $r=0.97$) and was related to the tokenizer representation (Pearson’s $r=0.78$; Supplementary Note 5).
+This pattern is consistent with evidence that, in visual contexts, brain-language correspondence can be driven by relatively coarse semantic content, closer to a bag-of-concepts organization than to progressively composed syntax 100.
+Interpreting this contrast depends on a linking assumption about inputs: vision models and cortex are compared while transforming pixel-level images, whereas language models operate on tokenized captions that already provide an abstracted description.
+Under this input mismatch, language-model alignment can emerge in regions that express higher-level scene organization, such as LOTC, without producing a depth-dependent gradient across the visual hierarchy.
 Figure 4: Inter-subject representational connectivity network.
 (A) Inter-subject connectivity matrix.
-Parcel-wise representational connectivity (RSA, Pearson’s $r$) is shown for 30 key cortical parcels ($N=8$).
-Each cell represents the mean RSA score between the RDMs of two parcels, computed across all pairs of different individuals.
-The matrix reveals a clear block structure corresponding to three principal hubs: Early Visual Cortex (V1–V4), the Ventral Hub (PHA1-3, VMV1–3), and the LOTC Hub (MT, MST, V4t, FST, TPOJ2–3).
-Warm colours denote shared representational geometry; cool colours denote dissimilarity.
+Parcel-wise representational connectivity (RSA; Pearson’s $r$) for the 30 parcels with the highest mean IS-RSA ($N=8$).
+Each cell shows the mean RSA between the RDMs of two parcels, computed across all cross-subject pairs.
+The matrix shows three clusters corresponding to early visual cortex (V1–V4), the Ventral hub (VMV1–3, PHA1–3), and the LOTC hub (V4t, MT, MST, FST, TPOJ2–3).
+Warm colors indicate higher representational similarity; cool colors indicate dissimilarity.
 (B) Directed connectivity graph.
-The same 30 parcels are plotted as nodes.
-Node colour indicates each parcel’s peak alignment depth with vision models, node size reflects its inter-subject alignment strength, and the node border colour marks its macro-anatomical group.
-The edges shown are the union of the three minimum-spanning trees (see Methods).
-Edge colour encodes the RSA value from panel (a), while edge width indicates the spanning tree order (from first to third), highlighting the network’s most central pathways.
-Arrowheads denote the putative direction of information flow, inferred from significant differences in peak-depth between connected parcels (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$).
-The layout reveals two primary processing streams emerging from early visual cortex: a medial-ventral route and a lateral-dorsal route projected to the LOTC hub.
-2.4 Representational connectivity reveals an ordered processing network
-To characterize the large-scale organization of these functionally distinct hubs, we extended our analysis from isolated regional analysis to the connectivity level.
-We constructed a parcel-by-parcel representational connectivity matrix by correlating the representational geometry of each parcel in one individual with that of every other parcel in all other individuals.
-This inter-subject representational connectivity approach directly captures a network structure that is reproducibly shared across the population.
-To estimate the direction of information flow within this network, we used the peak alignment depth from our vision model analysis as a proxy for hierarchical position, under the assumption that information flows from regions preferring shallower layers to those preferring deeper ones–a principle consistent with feed-forward processing in both the visual system and deep networks 16, 17.
-The same network structure was obtained using a conventional within-subject approach, confirming the robustness of our findings (see Supplementary Fig.
-D5).
-The parcel-by-parcel connectivity matrix revealed three internally connected clusters corresponding to the Early Visual, Ventral, and LOTC hubs (Fig.
+The same parcels are shown as nodes.
+Node color encodes each parcel’s peak alignment depth with vision models (heuristic proxy for hierarchical position), node size indicates inter-subject alignment strength, and node border color indicates macro-anatomical group membership.
+Edges correspond to the union of three minimum spanning trees; edge color reflects the RSA value in panel A, and edge width indicates the spanning-tree iteration (first to third).
+Arrow direction is assigned from shallower to deeper peak-depth parcels when the depth difference is significant across subjects (paired $t$-test, $t(7)$, FDR-corrected $p<0.05$).
+The graph highlights two dominant routes from early visual cortex: a medial-ventral route toward the Ventral hub and a lateral route toward the LOTC hub.
+2.4 A representational network linking ventromedial and lateral hubs
+To characterize how regions with shared scene-evoked information relate to each other, we repurposed IS-RSA as a measure of representational connectivity.
+This defines a task- and stimulus-dependent network based on similarity in how stimuli are organized across regions, rather than on anatomical proximity or raw BOLD covariance.
+For each pair of parcels, we correlated their RDMs across different individuals, yielding a parcel-by-parcel connectivity matrix that quantifies how similarly distinct regions organize the stimulus set across observers (Fig.
 4a).
-The Early Visual and Ventral hubs were moderately connected, forming an integrated system for visually grounded representation.
-In contrast, the LOTC hub exhibited a dissimilar representational geometry (negative RSA) relative to both the Ventral hub and classical dorsal regions (e.g., intraparietal sulcus).
-This pattern supports its functional dissociation as a distinct processing pathway, not only from ventral scene-perception areas like the PPA 54, 55, but also from other dorsal regions involved in spatial processing—a dissociation consistent with proposals for a distinct lateral or “third” visual pathway 30, 56.
-The directed graph, which highlights the network’s main connections, made the two-stream topology of information flow explicit (Fig.
-4b).
-Representational similarity originated in early visual cortex and bifurcated into two primary branches: a medial-ventral stream that passed through the Ventral hub, and a lateral-dorsal stream that traversed the LOTC hub—mirroring the dual-network organization previously described for scene processing 54.
-Our whole-cortex rendering (Supplementary Fig.
-D6) showed these streams extending anteriorly into posterior parietal, prefrontal, and default-mode regions, outlining a plausible route by which perceptual information can reach circuits for attention, memory, decision-making, and motor planning, consistent with known anatomical and functional pathways.
-In sum, our representational connectivity analysis identifies a stimulus-driven network whose core structure is consistent across individuals and is consistent with the dual-network topology previously reported for scene perception 54.
-This network originates in early visual cortex and bifurcates into two primary processing streams.
-The medial-ventral stream, terminating in the Ventral hub, corresponds to the classical pathway for scene and object recognition 26, 27.
-The lateral-dorsal stream, culminating in the LOTC hub, is highly consistent with recent proposals for a third visual pathway specialized for social perception 30.
-This functional dissociation raises the question of what information defines each of these processing routes—a question we address in the next section by isolating the dominant representational dimensions within each hub.
-2.5 Content Decomposition of Representational Hubs
-Figure 5: Dominant representational dimensions in cortical hubs.
-(A–C) Stimuli projection onto the first two shared components for each hub, extracted via Kernel Multi-view CCA (KMCCA) using the voxel data from all eight NSD participants for the common image set.
-In Early Visual Cortex (A), stimuli form a category-free cloud reflecting low-level visual similarity.
-In the Ventral Hub (B), the dominant axis arranges stimuli along a scene-to-object gradient.
-In the LOTC Hub (C), the first component separates animate from inanimate stimuli.
-(D) Impact on inter-subject RSA of controlling for the first KMCCA component via partial RSA.
-Controlling for the first component significantly reduced alignment in all hubs: Early Visual (6.6%), Ventral (3.1%), and LOTC (21.8%).
-The pronounced RSA reduction in LOTC underscores the animacy axis as its dominant organizing principle.
-(E, F) Cortical maps of inter-subject RSA recomputed separately for those scenes without (E) versus with (F) biological agents (people or animals).
-Alignment in the LOTC hub collapses when biological content is absent.
-Partial RSA values in (D) were computed parcel-wise and averaged by hub.
-All paired $t$-tests are Bonferroni-corrected ($t(7)$, $p<0.001$).
-The preceding analyses showed that the Early Visual, Ventral, and LOTC hubs form three distinct clusters embedded in two major processing streams.
-To uncover the content encoded in each hub, we decomposed the shared representational geometry using Kernel Multi-view Canonical Correlation Analysis (KMCCA)33.
-Standard RSA collapses two RDMs into a single Pearson’s $r$ that quantifies their alignment.
-By contrast, KMCCA treats each participant’s RDM as a separate “view” of the same stimulus-pair dissimilarities and identifies a shared subspace of orthogonal components whose projections are maximally correlated across all views.
-Our use of KMCCA parallels multi-view CCA techniques57–but instead of operating on raw voxel data, it works directly on the correlation kernels.
-These canonical axes therefore decompose the overall inter-subject alignment into ranked representational dimensions, isolating the latent components that drive shared geometry.
-We focused our analysis on the first KMCCA component, which explains the largest fraction of inter-subject alignment.
-First, we identified the stimulus features associated with this component to interpret its semantic content (details in Supplementary Fig.
-A6).
-Second, we used partial RSA to quantify this component’s contribution to the overall alignment.
-In this analysis, a large reduction in RSA after controlling for the component would indicate a geometry dominated by that single axis, whereas a minimal change would imply a more complex, high-dimensional space with no single dominant factor.
-The KMCCA projections revealed a distinct organization for each hub.
-In Early Visual Cortex (Figs.
-5a and D8), stimuli formed a category-free cloud organized primarily by low-level visual features, consistent with its retinotopic properties rather than high-level semantics.
-In the Ventral hub (Figs.
-5b and D9), the dominant axis traced a gradient from panoramic scenes to isolated objects, recapitulating the classic scene-to-object gradient in ventromedial cortex 27, 54.
-By contrast, the LOTC hub (Figs.
-5c and D10) was organized along an animacy axis, with its first canonical component separating stimuli containing biological agents (people or animals) from inanimate scenes.
-[…]
-To determine whether this alignment drop in the LOTC hub was uniform across all images or driven specifically by biological content, we split the image set into stimuli containing animate agents (65.3%) and those without, then recomputed inter-subject alignment.
-The results confirmed the LOTC hub’s specialization: alignment vanished in the absence of biological elements in the scene, while it remained robust for scenes containing people or animals (Fig.
-5e-f).
-An analysis of the connectivity network using this same data split showed that the entire lateral stream from the Early Visual Cortex to the LOTC hub was significantly reduced without biological content (Supplementary Fig.
-D7).
-Together, these findings provided support from a representational approach for a lateral “third visual pathway” specialized in social and animate information 30, 56.
-Figure 6: Shared representational geometry is modulated by stimulus content and task.
-Alignment computed using the symmetric HCP atlas combining both hemispheres.
-(A–D) Replication in BOLD5000, using complex natural scenes.
-With a different task (valence rating) but with comparable complex scene stimuli that include social content, the main findings were replicated.
-Both the inter-subject alignment (A) and its correspondence with vision models (B) show a similar posterior-to-anterior progression.
-The layer-wise curves (C) reproduce the distinct hierarchical trend for each hub, and the connectivity matrix (D) reveals a similar network structure with a representationally distinct LOTC hub.
-(E–H) Replication in THINGS-fMRI, using object-centric images.
-With a stimulus set of isolated objects and a rapid oddball task, the extent of shared geometry is more restricted.
-Inter-subject alignment (E) and its correspondence with vision models (F) are largely confined to early visual cortex.
-While the layer-wise curves (G) retain their characteristic shapes, the connectivity matrix (H) is dominated by a single Early Visual hub.
-This suggests that rich scene and social content are necessary to engage the full processing hierarchy captured by shared representational geometry.
-2.6 Generalization across tasks and stimulus sets
-Our findings derived from NSD picture dataset showed that stimulus statistics can align brains with vision models during a long-format memory task, but representational geometry may depend on both image distribution and task context58, 59.
-To assess the degree to which our findings generalize to other encoding task environments, we replicated our analysis pipeline on two independent fMRI datasets: BOLD5000 (valence judgments on  5,000 natural scenes)24 and THINGS-fMRI (oddball detection on  9,000 object images)25.
-For each dataset, we derived single-trial voxel responses, computed inter-subject RSA maps, generated parcel-wise alignment scores for vision models, and computed a representational connectivity matrix using inter-subject, parcel-pair RSA (see Supplementary Section B for details).
-To test the generalization of these findings, we first analyzed the BOLD5000 dataset 24.
-The stimulus set for BOLD5000 is highly comparable to NSD, containing complex natural scenes with multiple objects and social content, including a large subset of images drawn directly from the same database.
-Despite this stimulus similarity, participants performed a different task (valence judgment) with a shorter exposure time.
-Nevertheless, the results on images from BOLD500 dataset results largely replicated our main findings.
-Inter-subject and vision-model alignment spanned early visual cortex, ventral occipito-temporal parcels, and a LOTC hub (Fig.
-6a,b), and the parcel-wise relationship between these two measures was well-described by the same power-law model ($R^{2}=0.70$).
-Furthermore, the layer-wise alignment curves recapitulated the same hierarchical trend observed in NSD: decreasing for the Early Visual Cortex, mixed for the Ventral Hub, and increasing for the LOTC Hub (Fig.
+This approach emphasizes the informational content of the connectivity by grouping regions that share a common representational format, complementing univariate and conventional functional-connectivity analyses 69, 48.
+The connectivity matrix showed three clusters aligned with the Early Visual, Ventral, and LOTC hubs (Fig.
+4a).
+Early visual and ventral parcels formed a broad posterior cluster that also included the occipital and parietal cortex, indicating similar stimulus organization across these regions.
+LOTC parcels formed a second cluster with low or negative similarity to the posterior cluster.
+The strongest links between these systems passed through intermediate parcels in lateral occipital and ventral temporal cortex (LO1-3, FFC, PH).
+Because IS-RSA reflects both shared structure and measurement noise, local reliability differences can affect edge strength.
+However, the same topology was recovered with noise-ceiling normalization (Supplementary Fig.
+6) and in a within-subject analysis (Supplementary Fig.
+3).
+To visualize the strongest routes, we extracted a sparse backbone graph by iteratively applying a minimum spanning tree algorithm to the parcel-wise connectivity matrix.
+We then oriented edges using each parcel’s peak alignment depth from the vision-model analysis (Fig.
+3d,f) as a descriptive proxy for representational level, from shallower-layer peaks toward deeper-layer peaks.
+This ordering is not anatomical and does not imply causal direction.
+In the resulting graph, early visual cortex bifurcated into two routes (Fig.
+4b): a medial-ventral path progressing through ventral occipitotemporal cortex into the Ventral hub, and a lateral path progressing through lateral occipital cortex into the LOTC hub.
+Extending the analysis to the whole brain (Supplementary Fig.
+7) showed both routes continuing into association cortex, consistent with these representational formats interfacing with downstream systems implicated in memory, action, and control60, 75.
+This two-route organization does not map cleanly onto the classical “ventral what / dorsal where” framing of visual processing 35, 28, 54, 90.
+The medial-ventral branch targeted parahippocampal and ventromedial visual parcels (including VMV and parahippocampal scene regions) that are consistently linked to scene layout and environmental context 27, 4, 26, and it matches proposals of a ventromedial, scene-based pathway that provides scene input to hippocampal memory systems 77, 78.
+The lateral route, by contrast, encompassed MT+ and LOTC and continued toward posterior STS and temporo-parietal cortex in the whole-brain graph, regions implicated in dynamic cues, biological motion, body perception, and social processing 6, 70, 78, 71, 15.
+In this representational network, regions typically associated with the classical dorsal stream in intraparietal and superior parietal cortex were less prominent in shared geometry, but they remained part of the broader posterior cluster.
+This pattern suggests that the LOTC-centered trajectory reflects a distinct lateral pathway whose shared stimulus organization is driven most strongly by biological and socially relevant content, rather than the spatial computations emphasized in traditional “where/how” accounts 70.
+2.5 Representational dimensions driving shared geometry
+Figure 5: KMCCA components and biological-content dependence of LOTC shared geometry.
+(A-C) Stimulus projections onto the first two KMCCA components for each hub, estimated from all eight NSD participants using the shared image set.
+Early Visual Cortex (A) shows no clear clustering by scene category.
+The Ventral hub (B) varies along a scene-object continuum.
+The LOTC hub (C) shows a categorical separation along an animacy-related axis (animate agents vs.
+non-animate scenes/objects).
+(D) Effect of removing the first KMCCA component (KMCCA1) using partial RSA.
+Orange boxplots show partial IS-RSA after regressing out KMCCA1 from the hub RDMs; blue shows the original IS-RSA.
+Removing this single dimension reduces inter-subject alignment in all hubs, with the largest reduction in LOTC.
+IS-RSA was computed per parcel and subject, averaged within hub, and then averaged across subjects; paired $t$-tests were Bonferroni-corrected ($t(7)$, $p<0.001$).
+Annotated $\Delta\rho$ reports the mean absolute change, and percentages denote the relative change normalized by baseline.
+(E-F) Inter-subject RSA maps recomputed after splitting the stimulus set into scenes with biological agents (E) versus without biological agents (F).
+LOTC alignment is strongly reduced when biological content is absent.
+The two-route organization of the scene network raises a concrete question: which stimulus information most strongly organizes the shared representational geometry within each hub and along each route.
+To address this directly from the data, under the present task context, we decomposed shared across-subject structure in the Early Visual, Ventral, and LOTC hubs using Kernel Multi-view Canonical Correlation Analysis (KMCCA) 40.
+Unlike RSA, which summarizes similarity between two RDMs with a single correlation, KMCCA treats each participant’s hub RDM as a separate view of the same stimulus set and learns components that maximize correlation across views.
+We applied multi-view CCA ideas 86 to RDM-derived kernels, yielding a low-dimensional set of components that captures the geometry shared across participants within each hub.
+To visualize this shared geometry, we projected stimuli onto the first two KMCCA components for each hub (Fig.
+5a-c).
+In early visual cortex, the projection showed a diffuse distribution with no clear clustering by scene category, consistent with representations dominated by low-level image structure under this stimulus set and task 22 (Fig.
+5a; Supplementary Fig.
+9).
+In the Ventral hub, the leading axis formed a continuous gradient from panoramic scenes to more object-like images, consistent with ventromedial scene-object organization and scene layout/context representations 27, 77, 78, 4, 26 (Fig.
+5b; Supplementary Fig.
+10).
+In the LOTC hub, the first component separated images containing biological agents (people/animals) from a broader set of non-biological images, including objects and landscapes (Fig.
+5c; Supplementary Fig.
+11), consistent with LOTC sensitivity to biological and social content 70, 95, 6.
+We then quantified how much this leading axis contributed to each hub’s shared geometry.
+We regressed out the first KMCCA component and recomputed inter-subject alignment using partial RSA (Fig.5d).
+Removing this single axis reduced IS-RSA in all three hubs (all $p<.001$, Bonferroni-corrected), but the reduction was substantially larger in LOTC ($30.4\%$) than in Early Visual Cortex ($8.0\%$; t(7)=10.78t(7)=-10.78, $p<.001$) or the Ventral hub ($3.7\%$; t(7)=12.21t(7)=-12.21, $p<.001$).
+This pattern indicates that an animacy-related dimension accounts for a large fraction of LOTC’s shared geometry, whereas Early Visual and Ventral hubs distribute shared variance across higher-dimensional structure (Supplementary Fig.20).
+Given the dominant agent-related axis in LOTC, we tested whether lateral-route alignment depended on biological content by splitting the stimulus set into scenes with biological agents (people or animals; 65.3% of images) versus scenes without them, and recomputing IS-RSA within each subset.
+Alignment in LOTC was strongly reduced for non-biological scenes but remained robust when biological agents were present (Fig.5e–f).
+The same split also weakened the lateral branch of the representational network linking early visual cortex to LOTC, reducing the detectability of the two-route backbone when biological content was absent (Supplementary Fig.
+8).
+This pattern indicates that biological-agent content is a major contributor to LOTC shared geometry in this stimulus set and task regime, consistent with proposals that lateral pathways are preferentially organized by socially and biologically relevant scene information.
+Figure 6: Comparison across independent datasets.
+Analyses were performed on the BOLD5000 and THINGS-fMRI datasets using the same pipeline as NSD.
+(A–E) BOLD5000 (complex scenes, valence task; N=4).
+Despite lower signal reliability and differences in task, the core network topology persists in this dataset, which shares similar scene statistics with NSD.
+(A, B) Inter-subject alignment (A) and vision-model alignment (B) recover the characteristic occipitotemporal footprint, including Ventral and LOTC hubs, though with reduced magnitude and spatial continuity compared to NSD.
+(C) Layer-wise alignment profiles (mean $\pm$ SEM across subjects) preserve the hierarchical ordering (Early Visual << Ventral << LOTC) but exhibit flatter curves.
+(D, E) Representational connectivity matrix (D) and backbone graph (E) successfully recover the dual-stream topology (based on a $k=3$ minimum spanning tree), segregating the medial–ventral route from the lateral route.
+In the graph (E), node fill indicates peak model depth and edge color indicates representational connectivity strength.
+(F–J) THINGS-fMRI (object-centric images, oddball task; N=3).
+With minimal scene layout and social content, the network contracts.
+(F, G) Inter-subject (F) and vision-model (G) alignment are restricted primarily to early and ventral visual cortex, with no prominent LOTC peak.
+(H) Layer-wise profiles (mean $\pm$ SEM) are reduced in magnitude and show flatter profiles, especially in ventral and LOTC hubs.
+(I, J) The lateral stream does not emerge as a distinct subsystem in the connectivity matrix (I), and the backbone graph (J) lacks the clear two-route segregation evident in NSD and BOLD5000, consistent with the content-dependence of the lateral pathway.
+2.6 Comparison across tasks and stimulus sets
+Our analyses in NSD identified a distributed scene network by quantifying representational geometry shared across individuals and relating it to model feature spaces.
+Because this geometry is stimulus- and task-dependent, its spatial expression can vary with the regularities present in the stimulus set 19, 41, task demands that reweight the stimulus features that are emphasized 8, 45, and measurement reliability, which bounds how well shared structure can be estimated from trial-level data 63, 21, 92.
+To assess how the inferred organization generalizes across experimental regimes, we applied the same pipeline to two independent datasets: BOLD5000 11 and THINGS-fMRI 46 (Fig.6).
+BOLD5000 11 uses natural images broadly similar to NSD but differs in task (valence judgment vs.
+memory), stimulus duration (1 s vs.
+3 s), and trial-level reliability under comparable estimation procedures 2.
+In this regime, inter-subject and vision-model alignment remained concentrated in occipital and occipitotemporal cortex, with clear involvement of the ventral and LOTC hubs identified in NSD (Fig.
+6a-b), consistent with both datasets sampling rich scene layout, context, and social content.
+However, the overall spatial profile differed from NSD in two main respects.
+Early visual cortex was not detected as a dominant peak, and the LOTC pattern was less spatially contiguous.
+These differences are likely attributable to reduced signal reliability resulting from the lower field strength (3T) and shorter presentation times.
+Despite these differences, vision-model alignment preserved a monotonic parcel-wise relationship with IS-RSA, as in NSD (power-law scaling $R^{2}=0.71$).
+Layer-wise profiles preserved the hub-specific ordering but were flatter than in NSD (Fig.
 6c).
-The representational connectivity matrix also revealed a similar core network structure, with the LOTC hub forming a separate subsystem (Fig.
+At the network level, representational connectivity in BOLD5000 still separated a broad posterior cluster spanning early visual, ventral, and adjacent cortex from a lateral occipitotemporal cluster that included LOTC together with nearby lateral occipital and ventral temporal parcels (Fig.
 6d).
-These results indicate that the stimulus-driven brain-model convergence is robust to changes in task and scanning parameters.
-Finally, we analyzed the THINGS-fMRI dataset 25, which presents a strong contrast in both stimulus content and task design.
-This dataset used briefly flashed (500ms), object-centric images that lack the complex scene and social content of NSD, and participants performed an orthogonal oddball task.
-We found that both inter-subject and vision-model alignment were largely confined to the early visual cortex (Fig.
-6e,f), though the strong parcel-wise relationship between the two measures was preserved (power-law fit, $R^{2}=0.81$).
-Despite this restricted spatial extent, the underlying hierarchical profiles for each hub type were maintained: layer-wise curves still showed the characteristic decreasing, distributed, and increasing patterns (Fig.
-6g).
-Consequently, the connectivity analysis revealed a single cluster in early visual cortex (Fig.
+LOTC was less segregated from neighboring lateral parcels than in NSD, but the backbone visualization still recovered a two-route organization from early visual cortex toward ventral versus lateral targets (Fig.
+6e), consistent with a ventromedial route associated with scene context and a lateral route associated with biological and social features.
+THINGS-fMRI differs from NSD in both stimulus content and task.
+It uses briefly presented (500 ms), object-centric images with minimal scene layout and reduced social content, and participants performed an orthogonal oddball task 46.
+In this regime, inter-subject alignment and vision-model alignment again remained tightly coupled (power-law scaling $R^{2}=0.80$; Fig.
+6f-g), but the spatial pattern was concentrated in occipital and ventral visual cortex.
+The NSD ventromedial/parahippocampal and LOTC hubs were not detected as prominent hubs, and LOTC resembled the NSD pattern obtained when restricting analyses to images without biological agents (Fig.
+5f).
+Layer-wise profiles retained the same ordering but were reduced in magnitude and less distinctive for the ventral and LOTC hubs (Fig.
 6h).
-This is consistent with a stimulus set lacking complex scene information and social interactions, and a passive task design that likely reduced the representational signal in higher-order cortical areas 55.
-Across three distinct datasets, vision-model alignment was spatially co-located with inter-subject alignment, a finding consistent with models capturing the stimulus-driven geometry shared across brains.
-However, the extent of this processing cascade was context-dependent.
-Rich, complex scenes engaged the full posterior-to-LOTC hierarchy, whereas isolated objects under brief exposure primarily engaged only early visual cortex.
-These results therefore showed that shared representational geometry is a robust framework for revealing the specific cortical networks where stimulus information is processed across both individuals and models.
+Consistent with this, representational connectivity did not separate an LOTC-centered subsystem (Fig.
+6i), and the backbone graph did not recover the two-route topology observed for NSD and BOLD5000 (Fig.
+6j).
+Across datasets, vision-model alignment continued to track inter-subject alignment, indicating that the same stimulus-locked geometry that is detectable across observers is also present in image-trained feature spaces.
+However, which hubs and routes emerged as prominent depended on the structure available in the stimulus set and on task/reliability constraints: scene-rich datasets expressed ventromedial/parahippocampal and LOTC involvement more clearly, whereas the object-centric THINGS regime emphasized early and ventral visual cortex and did not express a separable LOTC-centered route.
 3 Discussion
-Our study establishes a general, data-driven approach for mapping information processing routes in the human brain.
-By combining inter-subject representational alignment, model-based comparisons, and explainability techniques, we identified a large-scale, stimulus-driven cortical network that is conserved across individuals and robust to task and dataset variation.
-This network, recruited during natural scene perception, is organized into two primary streams: a medial-ventral pathway for visually grounded scene analysis, and a lateral-dorsal pathway specialized for social and biological content.
-This result provides a data-driven refinement of the canonical dual-stream model of vision, functionally supporting recent proposals for a “third visual pathway”30, 56 and demonstrating a complex functional architecture that is integrated into a whole-cortex network.
-The convergence we observed between brain and model representations—though not mechanistically identical—suggests that deep networks provide a rich "bag-of-features" capable of capturing multiple stages of cortical processing.
-This nuanced view is necessary to explain our findings for the high-level hubs.
-The Ventral hub, with its cumulative integration of both visual and abstract information, showed alignment across a wide range of model layers.
-The LOTC hub, in contrast, showed a highly specialized alignment: it corresponded almost exclusively with the deepest semantic layers and only for those stimuli containing biological content.
-These results indicate that a simple one-to-one hierarchical model cannot account for the distinct computational strategies of the cortex, where recurrence and functional specialization contrast with the strictly sequential architecture of the models used for comparison60, 61, 62.
-However, the strong correspondence between the alignment found across brains and with models suggests that despite these mechanistic differences, both systems are sensitive to the same underlying stimulus structure, captured in regions where this stimulus-driven information is most functionally relevant.
-The observation that brains and deep networks converge on similar representational geometries likely reflects shared constraints for efficiently encoding the structured statistics of the environment 5, 6, 63.
-The principle of efficient coding posits that the optimal way to compress sensory information is to accurately predict it; therefore, any system that learns to predict the world will necessarily develop compressed, efficient representations.
-The convergence we observe across brains and models likely arises because both systems, through different optimization processes (evolution and gradient descent), are finding similar solutions to this fundamental problem.
-This is increasingly recognized as a general principle in both neuroscience and artificial intelligence, with modern models converging on similar latent features across tasks and modalities 64, 65.
-Characterizing shared representational spaces thus provides a principled way to understand how both biological and artificial systems learn to extract meaningful information from their environments 4.
-Methodologically, this work underscores the potential of population-scale, data-driven analyses to move beyond static anatomical maps, enabling systematic tracing of functional routes in the brain.
-No single method used here would have sufficed; it is the combination of identifying a shared space across subjects, using model hierarchies to probe its content, and deploying explainability techniques to interpret the results that provides a complete picture.
-With increasing access to high-quality, large-scale datasets, this integrated approach provides a path for hypothesis generation across sensory, cognitive, and cross-modal domains.
-The integration of explainability methods further allows for direct interpretation of what dimensions or features are routed through specific pathways, with implications for theory and for applications such as decoding and brain-computer interfaces.
-Limitations remain, including the need for deeper mechanistic models linking representational geometry to neural computation and behavior, and for extending these methods to causal and cross-modal analyses.
-As recent work highlights, high RSA scores can arise from non-mechanistic similarities when using complex stimuli with inherent confounds 59.
-While our explainability analysis is a first step toward mitigating this, future work will require more precise methods to isolate the specific features driving alignment and to bridge the gap from representational similarity to true mechanistic understanding.
-Nonetheless, this study demonstrates that combining shared representational geometry across individuals with deep model features is a powerful framework for mapping and interpreting brain function at scale.
-Our findings highlight the importance of tracing not just where, but how and what information is routed through the brain.
-This provides a principled, generalizable methodology for the field—moving from describing correspondence to generating more precise, mechanistically testable hypotheses about information processing in the brain.
+In this study, we aimed to characterize how the human cortex encodes natural scenes by quantifying representational structure shared across individuals and comparing it to hierarchical features in DNNs.
+We identified a distributed network of cortical regions engaged by scene perception and characterized its major hubs and the stimulus features that organize the processing within it.
+Two representational routes during scene viewing
+We identified this distributed network engaged during scene perception using inter-subject representational connectivity, which captures similarity in how different regions organize the stimulus set.
+The resulting topology separated three prominent hubs connected by two dominant routes: an early visual hub, a ventral hub spanning ventromedial and parahippocampal scene regions, and an MT+/LOTC hub.
+The early visual hub bifurcated into branches toward the ventral and lateral hubs, and the ventral and lateral systems were additionally coupled via intermediate lateral occipital and ventral temporal parcels (Fig.
+4).
+This organization does not match the classical ventral “what” versus dorsal “where/how” partition 35, 54, but is instead consistent with proposals that fractionate visual processing into additional routes, including a ventromedial pathway that emphasizes scene context and a dorsolateral pathway centered on MT+ and STS-adjacent regions that is associated with social/biological information 78, 70, 58.
+The medial-ventral route aligns with accounts of a ventromedial “where/scene” pathway that supports environmental layout and spatial context, providing inputs to hippocampal circuitry77, 78, 76.
+In our data, parcels in this ventral hub showed broad correspondence to vision models across multiple depths (Fig.
+3b), indicating that their shared geometry relates to features distributed across multiple stages of the vision-model hierarchy.
+This pattern held for scenes with and without biological agents (Fig.
+5e–f).
+It was reduced in the object-centric THINGS dataset, which contains limited scene layout and context (Fig.
+6e–f), confirming that this route is most detectable when the stimulus set contains rich environmental information.
+Language-model alignment in these regions showed no correspondence (Fig.
+2d), diverging from findings using similarity-based models 94, 22, indicating that caption-derived similarity structure did not capture the visually grounded scene-layout geometry expressed along the medial-ventral route.
+The lateral route aligns with accounts of an MT+/STS-centered pathway specialized for socially relevant information, including “third visual pathway” and dorsolateral subdivision proposals 78, 70, 58.
+Here, shared geometry was dominated by biological-agent content.
+The leading LOTC dimension separated scenes with animate agents from those without, and restricting analyses to images without biological agents reduced LOTC inter-subject alignment and weakened the lateral branch of the network (Fig.
+5e-f; Supplementary Fig.
+8).
+LOTC parcels aligned most strongly with later stages of vision models (Fig.
+3c) and also aligned with language-model representations (Fig.
+2d,f,g), consistent with this route emphasizing socially structured distinctions that are expressed in deep visual features and are also present in text-derived semantic structure.
+Model-brain alignment across the visual hierarchy
+Model-brain RSA related the representational geometry measured in cortex to the ordered progression of features across model depth.
+For vision models, model-brain correspondence tracked the inter-subject shared geometry across parcels and scaled with IS-RSA with an approximate power-law relationship (Fig.
+2e).
+This suggests that the geometry shared across observers during scene viewing is also present in image-trained feature spaces, consistent with the view that a reduced set of coarse dimensions is shared between brains and models and accounts for a large fraction of brain-model correspondence in visual tasks 13.
+Layer-wise profiles revealed a parallel with the posterior-to-anterior ordering expected from the visual hierarchy (Fig.
+3) 20, 28, 54, but the mapping was not one-to-one.
+For example, while the ventral hub showed correspondence across multiple depths, the MT+/LOTC hub showed selective alignment restricted to specific stimulus categories.
+This is consistent with cortical representations drawing on features distributed across multiple stages and being modulated by stimulus content, rather than following a strictly feedforward depth-to-stage mapping of models.
+Language models exhibited a distinct profile, reflecting the nature of their input.
+Because these models processed tokenized captions–abstracted from pixel data–they did not follow the cortical depth gradient observed in vision models.
+Instead, alignment was concentrated in LOTC and adjacent temporal regions, driven primarily by biological-agent content.
+Layer-wise alignment rose early and plateaued, scaling with tokenizer alignment (Supplementary Note 5).
+This pattern suggests a correspondence driven by coarse semantic content–closer to a “bag-of-concepts” organization than to progressively composed syntax 100–indicating convergence on LOTC dimensions that are independent of low-level visual features.
+More generally, these findings align with observations that modern models tend to converge on similar latent representations across objectives and architectures as they scale 13, 49, 50.
+One interpretation is that shared stimulus statistics, combined with efficiency constraints 31, favor the emergence of common representational axes that organize natural inputs by capturing structure essential for prediction and compression.
+Reliability and interpreting RSA magnitudes
+Across analyses, vision-model alignment closely tracked inter-subject alignment (Fig.
+2e).
+RSA variants that differ in their noise sensitivity (e.g., within-subject RSA or alternative repetition-matching schemes) were also related by approximate power-law scaling (Supplementary Note 4).
+This pattern is consistent with these measures indexing a common stimulus-locked representational geometry, differing primarily in how strongly measurement noise attenuates the observed correlations.
+This explains cases where model-brain RSA numerically exceeded IS-RSA in parcels with low-to-moderate reliability (e.g., posterior cingulate cortex; Supplementary Fig.
+1).
+IS-RSA correlates two noisy brain estimates, whereas model-brain RSA pairs one noisy estimate with a effectively noiseless model RDM.
+When reliability is limited, the noise from both sources in IS-RSA compounds the attenuation, lowering the observed correlation more than in model-brain RSA, even if both track the same underlying geometry.
+These results indicate that raw IS-RSA values may not always serve as an appropriate numeric ceiling for model-brain correspondence 63, 21, 92.
+Instead, RSA magnitudes are most informative when interpreted relative to reliability.
+Crucially, while adequate signal quality is a prerequisite for observing shared geometry–meaning the spatial extent of these maps inherently reflects the distribution of reliable fMRI signal–reliability alone cannot explain the topological organization we observed.
+The recovery of dissociable streams with distinct feature selectivities (e.g., ventral vs.
+lateral) confirms that this framework identifies specific information-processing networks, rather than merely tracking cortex-wide variations in signal quality.
+The empirical scaling relations provided here offer a practical calibration for separating these effects.
+Conclusions and scope
+In this work, we characterized a cortical network engaged during scene viewing at the level of representational geometry.
+By integrating complementary analyses of alignment, representational connectivity, and dimensionality decomposition, we linked three questions often treated separately: where shared stimulus information is expressed, how that information is routed across regions, and which stimulus properties organize representations within different network hubs.
+Together, these analyses provide a comprehensive perspective on how scene information is structured across the human cortex.
+This representational approach offers a general and adaptable methodology for identifying task-engaged networks and comparing information transformation across diverse systems.
+However, this flexibility entails a tradeoff: inferences concern representational organization rather than the precise mechanisms that implement it 3, 9.
+Because these comparisons rely on global similarity structure, computational dimensions that are subtle but behaviorally relevant may contribute weakly to the measured geometry, or be missed entirely if not captured by the chosen metric or feature sampling 57.
+Consequently, these results should be interpreted as context-dependent–varying with stimulus sampling, task demands, and measurement reliability 25–motivating future work to causally isolate the specific stimulus factors driving these alignments.
+Despite these limitations, this framework provides a practical route for leveraging large-scale data to move beyond simply asking whether two systems encode information similarly.
+By decomposing shared geometry to isolate driving features, the approach helps identify candidate factors and interactions that explain what information is shared across cortex and how it is distributed.
+As dataset scale and reliability increase, these tools can bridge the gap between network-level descriptions and experiments probing how distributed cortical systems support perception and flexible behavior in dynamic environments.
 4 Methods
 4.1 Functional-MRI data
-For the main analyses of this study, we used data from the Natural Scenes Dataset (NSD)23, a high-resolution 7T fMRI dataset comprising recordings from 8 participants performing a visual memory task.
-On each trial, participants viewed a naturalistic image for 3 seconds and indicated whether it was old or new.
-The stimulus set included a wide variety of natural scenes, spanning indoor and outdoor environments, static object views, and complex multi-element situations, including a broad range of social contexts.
-Each participant completed between 32 and 40 sessions, with 750 trials per session conducted over the course of a year.
-For participants who completed all 40 sessions, each image was presented three times, yielding a total of 30,000 trials per participant.
-Full experimental details are provided in Allen et al.
-(2022)23.
-Analyses were based on the 1.0-mm volumetric preprocessing of the NSD dataset and the corresponding single-trial $\beta$-estimates (version 3) 23.
-We used GLM-denoised single-trial $\beta$-estimates 66, as their image-related information content has been validated in multiple decoding and image reconstruction studies 42, 67, 68.
-Voxels were grouped into cortical parcels using the HCP-MMP1.0 atlas 31, which defines 180 cortical regions per hemisphere.
-We employed the pre-aligned parcel masks provided by the NSD authors to extract voxel activity within each cortical region 23.
-Analyses were performed using both the symmetric version of the atlas (combining hemispheres) and the asymmetric version (preserving hemisphere separation); a supplementary analysis confirms that symmetric parcels display highly correlated representational profiles (Supplementary Fig.
-A1).
+For the main analyses, we used the Natural Scenes Dataset (NSD) 2, a high-resolution 7T fMRI dataset with 8 participants performing an image recognition-memory task.
+On each trial, participants viewed a natural scene for 3s and reported whether it was old or new.
+Each participant completed 32–40 scanning sessions (750 trials per session), yielding up to $\sim 30,000$ trials per participant with repeated presentations of images across sessions.
+Full acquisition and experimental details are reported in Allen et al.
+2.
+Analyses were based on the 1.0-mm volumetric preprocessing of the NSD dataset and the corresponding single-trial $\beta$-estimates (version 3) 2.
+We used GLM-denoised single-trial $\beta$-estimates 73, as their image-related information content has been validated in multiple decoding and image reconstruction studies 87, 65, 81.
+Voxels were grouped into cortical parcels using the HCP-MMP1.0 atlas 34, which defines 180 cortical regions per hemisphere.
+We employed the pre-aligned parcel masks provided by the NSD authors to extract voxel activity within each cortical region 2.
+Analyses were performed using both the symmetric version of the atlas (combining hemispheres) and the asymmetric version (preserving hemisphere separation); a supplementary analysis confirms that symmetric parcels display highly correlated representational profiles (Supplementary Note 3).
 In all figures, parcels are colored by macro-anatomical groups as defined in Huang et al.
 (2022) 47, to facilitate anatomical interpretation.
-For each participant $p~{}\in~{}\mathcal{P}$, parcel $r~{}\in~{}\mathcal{R}$, and session $s~{}\in~{}\mathcal{T}p$, we define a response matrix $X_{p,r,s}~{}\in~{}\mathbb{R}^{n\times v_{p,r}}$ containing the voxel responses to each stimulus in that session.
-Here, $n$ denotes the number of trials (750 per session), and $v_{p,r}$ the number of voxels in parcel $r$ for participant $p$.
-Comparative analyses were performed using fMRI data from the BOLD500024 and THINGS25 datasets.
+[…]
+Comparative analyses were performed using fMRI data from the BOLD5000 11, and THINGS 46 datasets.
+We selected these datasets because IS-RSA and KMCCA require a large number of matched stimuli across participants, which is often limited in larger-cohort naturalistic datasets.
 Preprocessing and analysis followed the same procedure described above, with minor adaptations to accommodate differences in stimulus sampling and repetition structure.
-Full details for these datasets are provided in Supplementary Section B.
+Details are provided in Supplementary Note 8.
 All data analysed are fully de-identified and publicly available.
-The original NSD collection was approved by the University of Minnesota Institutional Review Board (Allen et al.23).
-The THINGS dataset was collected under NIH Institutional Review Board approval (Hebart et al.25), and BOLD5000 under the Institutional Review Board of Carnegie Mellon University (Chang et al.24).
+The original NSD collection was approved by the University of Minnesota Institutional Review Board (Allen et al.2).
+The THINGS dataset was collected under NIH Institutional Review Board approval (Hebart et al.46), and BOLD5000 under the Institutional Review Board of Carnegie Mellon University (Chang et al.11).
 No new human-subjects data were acquired for this work.
@@
 To compare how deep learning models encode visual and linguistic stimuli relative to fMRI responses, we selected vision and language models previously used by Huh et al.
-(2024) in studies of representational convergence across modalities 64.
+(2024) in studies of representational convergence across modalities 49.
 Model selection was conducted prior to analysis, balancing computational constraints and the goal of covering diverse training objectives and model scales.
 A full list of models is provided in Table 1.
-We selected 17 vision models spanning families with distinct training regimes, including ViT AugReg69, CLIP70, 71, DINOv272, and MAE73.
-All architectures were based on sequential Vision Transformer (ViT) blocks74.
+We selected 17 vision models spanning families with distinct training regimes, including ViT AugReg84, CLIP74, 14, DINOv264, and MAE44.
+All architectures were based on sequential Vision Transformer (ViT) blocks24.
 Each NSD image was processed through the backbone of each model, and activations were extracted from the output of each ViT block (see Fig.
 1A).
-For language models, we selected 14 transformer-based encoders from the BLOOMZ75, Gemma 276, LLaMA77, 78, and LLaMA 379 families80.
-For each NSD image, we used text captions annotated by trained raters and originally sourced from the MS-COCO dataset81.
+For language models, we selected 14 transformer-based encoders from the BLOOMZ61, Gemma 232, LLaMA88, 33, and LLaMA 336 families89.
+For each NSD image, we used text captions annotated by trained raters and originally sourced from the MS-COCO dataset59.
 Captions were processed through each model’s encoder, and activations were extracted at the output of each transformer block.
-Feature extraction was performed using the Transformers82 and TIMM83 libraries, adapting procedures from Huh et al.
-(2024)64.
-All models were implemented in PyTorch84 and retrieved from the Hugging Face Model Hub.
+Feature extraction was performed using the Transformers97 and TIMM96 libraries, adapting procedures from Huh et al.49.
+All models were implemented in PyTorch67 and retrieved from the Hugging Face Model Hub.
 For each model $m\in\mathcal{M}$, layer $l\in\mathcal{L}_{m}$, and stimulus set $s\in\mathcal{T}$ (images or their corresponding captions), we constructed a matrix of layer activations $Z_{m,l,s}\in\mathbb{R}^{n_{s}\times v_{m}}$, where $n_{s}$ is the number of stimuli in set $s$ and $v_{m}$ is the model’s embedding size, which is constant across all layers in the selected models.
@@
 Modality Family Training Regime Model # Params # Blocks Embedding size
-Vision ViT (AugReg)69 Supervised (21 K classes) ViT (AugReg) Tiny 10M 12 192
+Vision ViT (AugReg)84 Supervised (21 K classes) ViT (AugReg) Tiny 10M 12 192
 ViT (AugReg) Small 30M 12 384
@@
 ViT (AugReg) Large 326M 24 1024
-CLIP (Vision)70, 71 Contrastive Image–Text CLIP (ViT) Base 86M 12 768
+CLIP (Vision)74, 14 Contrastive Image–Text CLIP (ViT) Base 86M 12 768
 CLIP (ViT) Large 304M 24 1024
 CLIP (ViT) Huge 632M 32 1280
-Contrastive I-T + fine-tuning (12 K) CLIP (ViT) Base ft-12k 95M 12 768
-CLIP (ViT) Large ft-12k 315M 24 1024
-CLIP (ViT) Huge ft-12k 646 M 32 1280
-DINOv272 Self-Supervised DinoV2 (ViT) Small 22M 12 384
+Contrastive I-T + fine‑tuning (12 K) CLIP (ViT) Base ft‑12k 95M 12 768
+CLIP (ViT) Large ft‑12k 315M 24 1024
+CLIP (ViT) Huge ft‑12k 646 M 32 1280
+DINOv264 Self‑Supervised DinoV2 (ViT) Small 22M 12 384
 DinoV2 (ViT) Base 87M 12 768
@@
 DinoV2 (ViT) Giant 1.1B 40 1536
-MAE73 Self-Supervised (Masked AE) MAE (ViT) Base 86M 12 768
+MAE44 Self‑Supervised (Masked AE) MAE (ViT) Base 86M 12 768
 MAE (ViT) Large 303M 24 1024
 MAE (ViT) Huge 631M 32 1280
-Language BLOOMZ75 Causal LM + Instruction FT BloomZ (560M) 559M 25 1024
+Language BLOOMZ61 Causal LM + Instruction FT BloomZ (560M) 559M 25 1024
 BloomZ (1b1) 1.1B 25 1536
@@
 BloomZ (7b1) 7.1B 31 4096
-Gemma 276 Causal LM Gemma 2 (2B) 2.6B 27 2304
+Gemma 232 Causal LM Gemma 2 (2B) 2.6B 27 2304
 Gemma 2 (9B) 9.2B 43 3584
-LLaMA77, 78 Causal LM OpenLlama (3B) 3.4B 27 3200
+LLaMA88, 33 Causal LM OpenLlama (3B) 3.4B 27 3200
 OpenLlama (7B) 6.7B 33 4096
@@
 HuggyLLaMa (13B) 13B 41 5120
-LLaMA 379 Causal LM Llama 3 (8B) 8B 33 4096
+LLaMA 336 Causal LM Llama 3 (8B) 8B 33 4096
 Llama 3.1 (8B) 8B 33 4096
@@
 4.3 Alignment Measure
-We used Representational Similarity Analysis (RSA)32 to quantify the correspondence between brain parcels and model layers geometries.
+We used Representational Similarity Analysis (RSA)55 to quantify the correspondence between brain parcels and model layers geometries.
 For each system, the representational dissimilarity matrix (RDM) $D(X)$ was defined as $D(X)_{ij}=1-\rho(x_{i},x_{j})$, where $x_{i}$ and $x_{j}$ are rows of $X$ and $\rho$ is the Pearson correlation.
@@
 $\operatorname{RSA}(X,Y)=\rho\left(\operatorname{vec}_{u}D(X),\;\operatorname{vec}_{u}D(Y)\right)$ (1)
-Main analyses were replicated with Spearman correlation and Centered Kernel Alignment (CKA)85, yielding consistent results (see Supplementary Fig.A5).
-Details of the matrix reformulation that enabled efficient, large-scale RSA computations are provided in Supplementary Section C.
+Main analyses were replicated with Spearman correlation and Centered Kernel Alignment (CKA)53, yielding consistent results (see Supplementary Note 6).
+Details of the matrix reformulation that enabled efficient, large-scale RSA computations are provided in Supplementary Note 9.
 4.4 Inter-Subject alignment
-Inter-subject RSA was computed using the subset of 1,000 images viewed by all NSD participants (the shared1000 set)23.
+Inter-subject RSA was computed using the subset of 1,000 images viewed by all NSD participants (the shared1000 set)2.
 Each image in this subset was presented up to three times per participant, with corresponding trial indices across sessions.
@@
 Images shared across participants were always presented at the same trial positions for all participants, interleaved with unique images specific to each individual.
-To attenuate alignment driven by task or session structure rather than by true stimulus processing, we applied a cyclic repetition shift when matching trials across participants.
+To attenuate alignment driven by task or session structure rather than by stimulus processing, we applied a cyclic repetition shift when matching trials across participants.
 Specifically, for each image $i$ and repetition index $k$, the trial $(i,k)$ in participant $p$ was matched with the trial $(i,(k\bmod 3)+1)$ in participant $q$, thereby preserving stimulus identity while disrupting repetition-locked confounds.
-This matching strategy and its impact are illustrated and compared in Supplementary Fig.
-A2.
+This matching strategy and its impact are illustrated and compared in Supplementary Note 1.
 Let $\boldsymbol{i}_{pq}$ and $\boldsymbol{j}_{pq}$ denote the matched stimulus indices for participants $p$ and $q$, respectively.
-The inter-subject alignment between parcels $r$ and $r^{\prime}$ was then computed as:
+The inter-subject alignment between parcels $r$ and $r^{\prime}$ was computed as:
 $\mathcal{A}^{\mathrm{IS}}_{p,q,r,r^{\prime}}=\operatorname{RSA}\left(X_{p,r}[\boldsymbol{i}_{pq},:],\;X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]\right).$ (2)
@@
 To benchmark the degree of shared geometry relative to individual consistency, we also compared inter-subject and within-subject alignment (see Supplementary Fig.
-D5), which yielded consistent results.
+3), which yielded consistent results.
 A parcel-by-parcel matrix of group-level inter-subject alignment values ($n_{\mathrm{rois}}\times n_{\mathrm{rois}}$) defined the whole-cortex connectivity network.
-We visualized the main network backbone and multiple major information routes using an iterative minimum spanning tree (k-MST) approach on the group-averaged connectivity matrix.
-At each iteration, a standard MST was computed (with edge weights defined as $1-\mathrm{similarity}$), penalizing previously selected edges to ensure the inclusion of new, critical pathways.
+We visualized the main network backbone and multiple major representation routes using an iterative minimum spanning tree (k-MST) approach on the group-averaged connectivity matrix.
+At each iteration, a standard MST was computed (with edge weights defined as $1-\mathrm{similarity}$), removing previously selected edges to ensure the inclusion of new pathways.
 The final backbone comprised the union of edges from all $k$ MSTs (see Supplementary Fig.
-D6), preserving multiple principal paths, avoiding arbitrary edge thresholding, and ensuring that all parcels remained connected while allowing for richer topology.
-MST computations were performed using the NetworkX and SciPy libraries86, 87.
+7), preserving multiple principal paths, avoiding arbitrary edge thresholding, and ensuring that all parcels remained connected while allowing for richer topology.
+MST computations were performed using the NetworkX and SciPy libraries91, 39.
 Statistical significance was evaluated by comparing the observed alignment values to a null distribution generated by randomly shuffling stimulus labels (10,000 permutations, using a shared permutation scheme across all comparisons) with a two-tailed test.
-All p-values were FDR-corrected 88.
+All p-values were FDR-corrected 7.
 4.5 Brain–Model alignment
@@
 $\mathcal{A}^{\mathrm{BM}}_{p,r,m}=\frac{1}{|\mathcal{T}_{p}|}\sum_{s\in\mathcal{T}_{p}}\operatorname*{sign\,max}_{l\in\mathcal{L}_{m}}\,\mathcal{C}_{p,r,s,m}(l).$ (5)
-This procedure, standard in recent literature 51, 64, ensures that alignment is captured independently of model depth.
-In cases—such as for language models—where correlations may be systematically negative, this approach preserves the interpretability of dissimilar representational geometries: negative alignment values indicate that stimuli which are close in one space (e.g., visual cortex) are systematically distant in another (e.g., language model embedding)89.
+This procedure, standard in recent literature 79, 49, ensures that alignment is captured independently of model depth.
+In cases-such as for language models-where correlations may be systematically negative, this approach preserves the interpretability of dissimilar representational geometries: negative alignment values indicate that stimuli which are close in one space (e.g., visual cortex) are systematically distant in another (e.g., language model embedding)80.
 To identify where alignment peaked within the model hierarchy (e.g., Fig. 3b–c), we computed the normalized depth of the maximally aligned layer for each session and averaged across sessions:
@@
 where layers are indexed from 0 (first layer) to $|\mathcal{L}_{m}|-1$ (last layer), so $d^{*}_{p,r,m}\in[0,1]$.
-This procedure is standard for mapping the correspondence between brain regions and network hierarchy 19, 51.
+This procedure is standard for mapping the correspondence between brain regions and network hierarchy 10, 79.
 To obtain a subject-level measure for a set of models (e.g., all vision or language models), both the peak alignment and normalized depth measures were averaged across models.
@@
 4.6 Shared Component Extraction
-To identify the dominant dimensions underlying shared representational geometry across participants, we applied kernel multi-view canonical correlation analysis (KMCCA) 33.
+To identify the dominant dimensions underlying shared representational geometry across participants, we applied kernel multi-view canonical correlation analysis (KMCCA) 40.
 KMCCA was performed on the set of image trials shared by all participants in NSD, extracting a low-dimensional subspace that maximized correlation across participants’ RDMs for each hub.
-We used the KMCCA implementation from mvlearn90.
+We used the KMCCA implementation from mvlearn68.
 For each hub, we extracted the first two components for exploratory purposes, and assessed the influence of the first (dominant) component in control analyses.
 While our focus was on the dominant axis, further work is needed to interpret all statistically significant components.
-Image-level semantic annotations and category labels were derived from MS-COCO 81.
-For automated labeling of the scene-to-object gradient, we used Pixtral-12B 91 (see Fig.
+Image-level semantic annotations and category labels were derived from MS-COCO 59.
+For automated labeling of the scene-to-object gradient, we used Pixtral-12B 1 (see Fig.
 5a).
-Further details of the KMCCA procedure, semantic projections, and partial RSA controls are provided in Supplementary Section A.5.
+Further details of the KMCCA procedure, semantic projections, and partial RSA controls are provided in Supplementary Note 7.
+4.7 Statistics and Reproducibility
+Statistical testing and analysis were carried out in Python 3.10 using standard libraries (SciPy, Statsmodels).
+No statistical methods were used to pre-determine sample size; sample sizes were determined by availability in the public datasets: Natural Scenes Dataset ($N=8$), BOLD5000 ($N=4$), and THINGS-fMRI ($N=3$).
+These sizes are standard for deep-sampling fMRI studies where individual-subject reliability is high.
+All subjects and sessions from these datasets were included in the analysis.
+To assess the statistical significance of RSA maps (e.g., Fig.
+2a-c), a two-tailed permutation test ($n=10,000$) was used to determine if RSA values significantly differed from a null distribution.
+For all other NSD group-level results (Figs.
+1–5), significance was assessed using two-tailed paired t-tests ($N=8$) unless otherwise specified.
+In all cases, $p$-values were corrected for multiple comparisons using the False Discovery Rate (FDR) or Bonferroni method, with significance defined as $p<0.05$.
 Data and code availability
@@
 Model checkpoints analyzed in this work are available on Hugging Face (https://huggingface.co/collections/pablomm/convergent-transformations-6808f7de248fa9674acac588).
-All resources were publicly accessible in July 2025.
+Precomputed derivative data generated from these raw sources are available in Figshare 10.6084/m9.figshare.30753239.
+All resources were publicly accessible in January 2026.
 Code to reproduce analyses is available at https://github.com/memory-formation/convergent-transformations.
@@
 Baldassano, and E.
-Cámara for helpful comments and discussions, the creators of the Natural Scenes Dataset, THINGS, and BOLD5000 for making their data publicly available, and the teams who developed and released the models on Hugging Face.
+Cámara for helpful comments and discussions.
 This work was supported by the Spanish Ministerio de Ciencia, Innovación y Universidades, which is part of Agencia Estatal de Investigación (AEI), through the project PID2022 - 140426NB (Co-funded by European Regional Development Fund.
@@
 Both authors revised and approved the final manuscript.
+Competing interests
+The authors declare no competing interests.
 Supplementary Material
-A Supplementary analyses
-A.1 Hemispheric Asymmetry and the Validity of Symmetric Analyses
-Figure A1: Hemispheric comparison of representational alignment.
-(A–C) Comparison of representational alignment (RSA, Pearson’s $r$; $N=8$, NSD) computed independently for the left (x-axis) and right (y-axis) hemispheres: (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment.
-Each point is a cortical parcel, colored by macro-anatomical group.
-The diagonal line indicates equal alignment.
-Squares denote parcels with a statistically significant hemispheric difference (paired $t$-test, $p<0.05$, FDR-corrected).
-(D) Boxplots of alignment within principal hubs.
-Significant hemisphere effects are marked.
-(E) Correlation between cross-hemispheric and bilateral inter-subject alignment.
-Cross-hemispheric RSA (y-axis) was computed between a parcel in one subject’s left hemisphere and the corresponding parcel in another subject’s right hemisphere, compared to bilateral RSA (x-axis).
-The near-perfect correlation ($r=0.99$) shows that combining hemispheres preserves geometry.
-We quantified hemispheric asymmetry in representational geometry by computing alignment measures separately for the left and right hemispheres using the HCP-MMP1.0 atlas31.
-For each cortical parcel, we assessed (i) inter-subject alignment, (ii) brain–vision model alignment, and (iii) brain–language model alignment, using the same Natural Scenes Dataset (NSD)23 responses as in the main analyses (see Fig.
-A1).
-Across all modalities, a largely symmetric pattern was observed (Fig.
-A1a–c): the rank-order of parcels was highly similar between hemispheres (Spearman’s $r$: inter-subject $=0.93$, vision $=0.92$, language $=0.82$), although some significant asymmetries were present.
-Specifically, early visual cortex (V1–V4) showed stronger alignment in the left hemisphere, while alignment was higher in the right hemisphere for ventral and LOTC hub parcels (hMT+, LO, TPOJ).
-Vision- and language-model alignment followed similar trends.
-Grouping parcels by functional hub (Fig.
-A1d) highlighted these asymmetries: left hemisphere alignment was higher in early visual cortex, while right hemisphere alignment was higher in ventral and LOTC hubs.
-To determine whether these asymmetries reflected differences in representational content or only magnitude, we compared cross-hemispheric RSA (e.g., participant $p$’s left V1 vs subject $q$’s right V1) to bilateral RSA (subject $p$’s left+right V1 vs subject $q$’s left+right V1).
-These measures were almost perfectly correlated across parcels ($r=0.99$, Fig.
-A1e), with no parcel deviating from the regression line.
-This demonstrates that combining hemispheres increases overall alignment strength without distorting the underlying representational geometry—no region showed a unique lateralized axis absent in the contralateral hemisphere.
-Together, these results justify our approach of using symmetric (hemisphere-combined) parcels for the main analyses: (i) combining ROIs across hemispheres maximizes statistical power for multivariate comparisons without introducing distortions in representational geometry; (ii) hemisphere-separated ROIs remain informative for quantifying lateralization effects.
-This flexible approach supports both detailed lateralization analysis and robust, large-scale cross-dataset RSA experiments with joined hemispheres.
-Figure A2: Repetition-matching strategies and their impact on inter-subject RSA.
-To compute inter-subject RSA, we compared different strategies for matching the three trial repetitions for each shared image between any pair of participants.
-(A) Schematic of unshifted (repetition 1→1, 2→2, 3→3) versus shifted (cyclic 1→2, 2→3, 3→1) trial matching with example RDMs.
-(B, C)Group-average cortical RSA maps under unshifted (B) and shifted (C) matching; both reveal the early-visual, ventral, and LOTC hubs.
-(D) Parcel-wise comparison of shifted vs.
-unshifted RSA values shows a uniform attenuation under shifted matching, with two linear regimes reflecting stronger prefrontal attenuation.
-A.2 Repetition-Matching Strategies for Inter-Subject RSA
-The Natural Scenes Dataset (NSD) presents 10,000 unique images across up to 40 scanning sessions per participant (30,000 trials total).
-A subset of 1,000 images (3,000 trials) was shared by all eight participants and always appeared in the same trial positions—interleaved among participant-unique stimuli—so that each shared image was viewed under identical practice, fatigue, and session-context conditions.
-This locked structure raises the possibility that inter-subject correlations might reflect non-perceptual factors (e.g., trial timing or memory demands) rather than purely stimulus-driven geometry.
-To isolate the stimulus-driven component, we adopted a “shifted-repetition” matching strategy when computing inter-subject representational similarity (IS-RSA).
-Rather than pairing each repetition index $k\leftrightarrow k$ across participants (“unshifted”), we cyclically permute so that repetition $k$ in participant $p$ matches repetition $(k\bmod 3)+1$ in participant $q$ (i.e.
-$1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$; see Fig.
-A2A).
-Because all shared trials occupy the same session slots, this shift preserves image identity while breaking any exact repetition-locked confounds and also allows direct comparison with within-subject analyses (which must compare different repetitions; see Supplementary Fig.
-D5).
-Due to symmetry across participant-pair comparisons, shifts of 1 and 2 yield identical group-average maps.
-Recomputing the group-average cortical RSA under both unshifted (Fig.
-A2B) and shifted (Fig.
-A2C) schemes recovers the same three hubs—early visual cortex, ventral occipito-temporal parcels, and LOTC.
-As expected, shifting is more conservative: alignment magnitudes are uniformly lower (Supplementary Fig.
-A2D), yet the relative rank-order of all parcels remains unchanged (Spearman’s $\rho=0.96$).
-Frontal parcels (e.g.
-FEF, inferior frontal gyrus) suffer proportionally larger attenuation under shifting than occipito-temporal hubs, suggesting that these frontal regions encode task- or repetition-related variance (e.g.
-decision or attentional control) that the shifted criterion minimizes.
-By contrast, the robust persistence of occipital and temporal alignments underscores their stable, stimulus-driven representational geometry.
-To illustrate how these differences manifest in network structure, Fig.
-A3 shows representational connectivity graphs for both matching strategies.
-We first identified parcels whose inter-subject RSA exceeded an absolute threshold of $r=0.05$ in either the shifted (shift = 1 or 2) or unshifted maps.
-We then constructed a reduced connectivity matrix among this common set of nodes and extracted a two-iteration minimum-spanning-tree backbone (using edge weights defined as $1-\mathrm{RSA}$; see main Methods).
-The core two-stream topology (Early Visual $\rightarrow$ Ventral hub and Early Visual $\rightarrow$ LOTC hub) is preserved under both schemes.
-However, only the unshifted graph highlights FEF as a high-centrality node, further suggesting that frontal regions carry shared task- or repetition-locked signals that are attenuated by the shifted criterion.
-While these frontal effects are exploratory—given their lower RSA magnitudes and sensitivity to thresholding—they point to important future work on how task and attention signals coexist with stimulus-driven representations.
-At the same time, they underscore the necessity of shifted-repetition controls and of complementary interpretability techniques to isolate which stimulus-content features drive convergence, enabling mechanistic hypotheses about information transmitted across cortical routes.
-[…]
-A.3 Token-occurrence control for language–brain RSA
-To quantify how much of the language–brain alignment reflects initial tokenization, we constructed RDMs from token-occurrence vectors.
-Each caption was encoded as a count vector—each entry recording the number of times a given token appears—and pairwise Pearson dissimilarities among these vectors yielded a tokenizer-based RDM.
-We applied this procedure to six vocabularies (BLOOMZ, Gemma 2, LLaMA 1/2/3; see Table 1), covering all tokenizers used in our language models, and then computed parcel-wise RSA against NSD fMRI responses using the symmetric HCP atlas.
-Parcel-wise alignments were nearly identical across tokenizers (pairwise Spearman’s $\rho$ $\leq 0.978$), so we report their average.
-The tokenizer–brain alignment map closely matches that of full language models: it peaks in the LOTC hub and shows minimal or negative alignment in early visual parcels (Fig.A4A,D).
-Token-count geometry correlates strongly with each model’s first-layer alignment (Pearson’s $\rho=0.78$; Fig.A4B) and with maximal alignment across layers ($\rho=0.68$; Fig.
-A4C), indicating that token frequencies account for a substantial portion of the language–brain correspondence.
-Repeating the analysis with a word tokenizer—splitting captions on whitespace rather than into subword units—again yielded pronounced LOTC alignment (Fig.
-A4E).
-These findings suggest that LOTC’s alignment to language models largely reflects sensitivity to discrete lexical tokens common across descriptions—functioning as a “bag-of-words” representation—rather than to compositional or syntactic structure53.
-Figure A4: Tokenizer-level contributions to language–brain alignment.
-We derived RDMs from token-occurrence count vectors over each model’s vocabulary and computed parcel-wise RSA against NSD fMRI responses (averaged across the six tokenizers used in our language models).
-(A) Scatter of tokenizer-brain vs.
-inter-subject alignment, with peak values in the LOTC hub.
-(B) Token-count alignment against each model’s first-layer alignment.
-(C) Token-count alignment against each model’s maximal alignment across layers.
-(D) Cortical map of tokenizer-brain alignment, showing pronounced LOTC localization.
-(E) Alignment map for a whitespace-based word tokenizer, which likewise peaks in LOTC, consistent with a “bag-of-words” explanation rather than richer compositional structure.
-A.4 Robustness to Similarity Metric (Spearman-RSA & CKA)
-Recent work has cautioned that representational-similarity findings can depend not only on stimulus design but also on the choice of similarity metric—some measures probe local, pairwise geometry while others emphasize global, space-wide dependence 59.
-To test our results’ robustness to these different measures, we repeated key analyses using two alternatives to Pearson-RSA: (i) a rank-based Spearman-RSA, which preserves only the ordering of each stimulus pair’s dissimilarity, and (ii) an unbiased Centered Kernel Alignment (CKA) that quantifies global statistical dependence via an unbiased HSIC estimator 92.
-Replacing Pearson with Spearman correlations yields a purely rank-based, nonparametric measure agnostic to absolute distance magnitudes.
-As shown in Supplementary Fig.
-A5A–C, the inter-subject map, the parcel-by-parcel scatter against model alignment, and the representational connectivity matrix all mirror our original Pearson-RSA results (parcel-wise $r=0.99$).
-This confirms that neither the linearity nor distributional assumptions of Pearson correlation drive our findings—our local pairwise geometry holds under a fully nonparametric test.
-CKA measures dependence across entire representational spaces, emphasizing alignment along principal axes rather than isolated pairwise distances.
-We computed linear-kernel CKA (using the unbiased HSIC estimator 92) under the same shifted-repetition protocol.
-Supplementary Fig.
-A5D–F shows that CKA again recovers the three hubs but inflates LOTC alignment and attenuates early visual alignment in the parcel scatter, reflecting CKA’s sensitivity to low-dimensional, dominant components.
-Parcel-wise CKA and Pearson-RSA still correlate strongly (Pearson’s $r=0.91$, Spearman’s $\rho=0.93$), and the CKA connectivity matrix preserves the two-stream topology, despite CKA’s unsigned nature preventing a sign-based “dissimilarity” distinction.
-Across both local (Spearman-RSA) and global (CKA) criteria, the identification of Early Visual, Ventral, and LOTC hubs and their two-stream connectivity remains unchanged.
-This convergence across similarity measures, spanning local versus global criteria, strengthens confidence that our observed hierarchies and pathways reflect genuine stimulus-driven geometry rather than artifacts of any single metric.
-Figure A5: Comparison of alignment metrics.
-(A–C) Spearman-RSA.
-Replacing Pearson’s $r$ with Spearman’s rank correlation yields virtually identical maps: the inter-subject alignment (A), the parcel-wise scatter against vision- and language-model alignment (B), and the representational connectivity matrix (C) all match the Pearson-RSA findings, with parcel-wise strengths correlating at $r=0.99$.
-(D–F) Unbiased CKA.
-Centered Kernel Alignment, which captures global dependence via an unbiased HSIC estimator, again highlights Early Visual, Ventral, and LOTC hubs (D).
-In the parcel-wise comparison (E), CKA inflates LOTC alignment and attenuates early visual alignment—reflecting its sensitivity to low-dimensional, dominant components—yet CKA and Pearson-RSA remain strongly correlated ($r=0.91$).
-The CKA-based connectivity matrix (F) preserves the two-stream topology, though as a dependence measure it does not distinguish signed similarity.
-A.5 Extended details of Shared Component Decomposition and Partial RSA Controls
-Figure A6: Extended analysis of representational dimensions.
-(A) KMCCA projections onto the top two canonical axes for six representative parcels: V1, V4, VMV2, PHA2, MT, and TPOJ2.
-V1 shows no clear clustering; V4 begins to separate by content; Ventral and LOTC parcels replicate the hub-level semantic gradients.
-(B) Words ranked by loading on the first KMCCA component (positive loadings in blue, negative in red) for each hub.
-Early Visual words share low-level shape features; Ventral words span scene vs.
-object terms; LOTC words distinguish animate vs.
-inanimate labels.
-(C) Partial RSA boxplots for the three hubs, showing inter-subject alignment under no control (baseline), controlling for the scene-object gradient, or biological presence.
-LOTC alignment is markedly reduced by each control, whereas Early Visual and Ventral hubs are minimally affected.
-This supplementary note details our use of Kernel Multi-view Canonical Correlation Analysis (KMCCA)33 to isolate the axes driving our RSA findings32.
-While RSA provides a single global similarity score, KMCCA decomposes this correspondence into a ranked set of orthogonal “representational channels,” revealing which latent axes contribute most strongly to the shared geometry.
-We performed KMCCA using the Python package mvlearn (v0.5.0)90 with a correlation kernel and a regularization parameter of $\lambda=0.3$, using default parameters of implementation.
-For each cortical parcel (or hub), we formed a data matrix for each participant from the GLM-denoised single-trial $\beta$-estimates.
-We included only those trials for which each (image, repetition) pair was available for all eight participants.
-This procedure yielded a set of eight matrices ($X_{p,r}\in\mathbb{R}^{m\times v_{p,r}}$), one for each participant, all representing an identical sequence of stimulus trials.
-The KMCCA algorithm was then applied to this set of matrices, returning projections onto a common shared subspace.
-For visualization (Fig.
-A6A; Fig 5A-C), we projected each subject’s data onto the first two canonical axes and then averaged the projections for the repetitions of each unique image to produce a single, stable point per stimulus.
-The parcel-level KMCCA projections (Supplementary Fig.A6A) confirmed that the organizational principles observed at the hub level were consistent within their constituent parcels.
-In the Early Visual Cortex, the V1 parcel showed no clear semantic organization in its first two dimensions, whereas the V4 parcel already exhibited emergent clustering by image content.
-Within the Ventral hub, representative parcels such as VMV2 and PHA2 reproduced the hub-level scene-to-object gradient and showed clear organization by semantic category.
-Similarly, within the LOTC hub, parcels like MT and TPOJ2 showed the same primary separation between biological and non-biological stimuli.
-Within these main clusters, stimuli were further organized by more fine-grained semantic categories, mirroring the main hub-level analysis (Fig.5C).
-To interpret these KMCCA axes semantically, we projected the image captions onto the shared subspace learned for each hub.
-We first created a vocabulary from the union of all words present in the captions ($s=11,090$ unique words), encoding each caption as a binary word-occurrence vector (a “bag-of-words” representation).
-We then projected this caption matrix onto the first KMCCA component derived for each hub, which yielded a loading weight for each word in the vocabulary.
-By ranking these weights, we identified the words that contributed most positively and negatively to each dimension (Fig.
-A6B).
-This word-loading analysis confirmed our interpretation of each hub’s organizing principle: in Early Visual Cortex, top-weighted words were not semantically coherent but appeared to share coarse visual features; in the Ventral hub, positive tokens index scene contexts (e.g., ‘kitchen’, ‘bathroom’) while negative tokens name isolated objects (e.g., ‘plate’, ‘frisbee’); and in the LOTC hub, tokens cleanly separated animate from inanimate labels.
-This analysis provides a semantic interpretation for the organization of each cluster, which can be visually confirmed in the corresponding image atlases (See Supplementary Figs.
-D8-D10)
-Finally, we quantified the impact of the identified dimensions on the overall inter-subject alignment using partial RSA.
-Denoting by $\rho(X,Y)$ the standard Pearson correlation use to compute the alignment between two RDMs, $Z$ a third RDM (e.g., one built from the first KMCCA component or from a binary “biological presence” vector), the partial RSA was computed as:
-$\mathrm{pRSA}(X,Y\mid Z)=\frac{\rho(X,Y)\;-\;\rho(X,Z)\,\rho(Y,Z)}{\sqrt{(1-\rho(X,Z)^{2})\,(1-\rho(Y,Z)^{2})}}.$ (A1)
-As shown in the main text (Fig.
-5d-f), controlling for the continuous, data-driven KMCCA axis produced only minimal drops in the Early Visual and Ventral hubs, but a substantial reduction in the LOTC hub.
-A complementary analysis confirmed this result: controlling for the discrete categorical variables yielded a drop in alignment of a similar magnitude to that produced by the continuous KMCCCA components (Supplementary Fig.
-A6C).
-A.6 Power-Law Attenuation Model of RSA Metrics
-In the main text, we compared several RSA measures—including within-subject (WS), inter-subject (IS-1: shift=1 and IS-0: unshifted), and vision-model-brain (VM)–that probed a common cortical geometry but were subject to different levels of measurement noise.
-While all metrics revealed a similar cortical alignment pattern, their magnitudes varied considerably.
-Parcel-wise scatter plots (e.g., Fig.
-2E; Fig.
-A2) showed a systematic relationship between the metrics, which were tightly related by power-law curves.
-Although these curves differed in scale and curvature, their consistent form suggested that each metric captured the same latent similarity while imposing a unique, noise-dependent attenuation.
-We formalise this observation by modelling the correlation measured for parcel r with metric j as a power-law transform of a latent correlation:
-$\rho^{\mathrm{obs}}_{rj}=a_{j}\bigl{(}\rho^{\mathrm{latent}}_{r}\bigr{)}^{\,b_{j}}+\epsilon_{rj},$ (A2)
-where
-  • $\rho^{\mathrm{latent}}_{r}\in[0,1]$ is the metric-independent similarity common to all measures,
-  • $a_{j}\in[0,1]$ is a linear attenuation factor for metric j, and
-  • $b_{j}\geq 1$ captures signal-dependent (heteroscedastic) noise.
-A value of $b_{j}=1$ corresponds to simple homoscedastic attenuation, whereas $b_{j}>1$ creates a convex curve that suppresses weak latent correlations more strongly than strong ones.
-This constraint is motivated by our observation that the noisiest metrics (e.g., IS-shift) systematically underestimate weakly aligned parcels relative to cleaner metrics (e.g., WS), which is consistent with a noise regime that disproportionately affects weaker signals.
-  • $\epsilon_{rj}$ is a residual term that absorbs any remaining parcel–metric–specific variability not captured by the deterministic power-law component.
-We jointly fitted the parameters of Eq.
-A2 to four core metrics (WS, IS-0, IS-1, VM) using data from the symmetric-hemisphere ROIs.
-Parameter estimation was performed via the L-BFGS-B algorithm by minimizing a least-squares loss, with parameters bounded as follows: $a_{j}\in[0,1]$, $b_{j}\in[1,\infty)$, and $\rho^{\mathrm{true}}_{r}\in[0,1]$.
-Figure A7A shows the model’s high goodness-of-fit by plotting the observed correlations against the model-fitted values for the four core metrics.
-In all cases, the model explained over 94% of the variance ($R^{2}>0.94$), confirming that the power-law attenuation model accurately predicts the trends in the observed data.
-Complementing this, Figure A7B visualizes the distinct attenuation profile for each RSA strategy, allowing for a direct observation of how each measured correlation is attenuated in comparison with the estimated latent correlation.
-Because all metrics are modeled as functions of the same latent variable, $\rho^{\mathrm{true}}_{r}$, the relationship between any two metrics, j and j’, can be derived by eliminating this common term:
-$y=\frac{a_{j^{\prime}}}{a_{j}^{\,b_{j^{\prime}}/b_{j}}}\;x^{\,b_{j^{\prime}}/b_{j}},$ (A3)
-where $x=\rho^{\mathrm{obs}}_{rj}$ and $y=\rho^{\mathrm{obs}}_{rj^{\prime}}$.
-Figure A7C validates this prediction by overlaying the derived curve (Eq.
-A3) on the empirical parcel scatters.
-This confirmed that a single latent similarity, subject to metric-specific attenuation, explains all pairwise relations.
-For completeness, Figure A8 extends this validation to a set of 12 metrics, demonstrating the model’s power to unify all RSA strategies used in our analyses.
-The resulting cortical map of the latent variable, $\rho^{\mathrm{latent}}_{r}$, is displayed in Figure A7D.
-This map represents a noise-corrected estimate of similarity that unifies the four core RSA measures.
-It reveals a three-hub pattern of high alignment observed in the main text, with peak correlations ($\rho^{\mathrm{latent}}_{r}\sim 0.50$) in the EVC, Ventral stream, and LOTC.
-Furthermore, the model shows that prefrontal areas (FEF and pFC) are more prominent after correcting for the attenuation present in the individual raw measurements.
-The fact that these hubs align neatly with those highlighted by individual metrics confirms that our model (Eq.
-A2) successfully isolates a common signal while factoring out metric-specific noise.
-Because Eq.
-(A2) is scale-free, absolute values of $\hat{\rho}^{\mathrm{latent}}$ and $a_{j}$ are defined only up to a common factor.
-We fixed no anchor during optimisation, yet repeated fits converged to the same scale, and all qualitative findings (relative parcel ranking, cross-metric relations) are invariant under rescaling.
-The power-law attenuation model offers a compact, quantitative account of how measurement differences degrades RSA scores across diverse metrics.
-Its success supports the view that all metrics probe a shared latent representational geometry, differing only in the reliability with which each metric samples that geometry.
-Figure A7: (A) Goodness-of-fit.
-Parcel-wise observed RSA values (y-axis) are plotted against the corresponding predictions from the joint power-law model (Eq.
-A2; x-axis).
-Colours denote the different metrics (WS: within-subject; IS-1: inter-subject shifted (shift=1); IS-0: inter-subject unshifted; VM: vision model).
-The legend provides the explained variance ($R^{2}$) for each metric.
-(B) Correlation attenuation curves.
-The estimated latent correlation ($\hat{\rho}^{\mathrm{latent}}$, x-axis) is plotted against the observed correlation for each metric (y-axis).
-Coloured lines show the fitted metric-specific attenuation curves ($y=a_{j}\,x^{b_{j}}$).
-(C) Pairwise metric comparisons.
-Scatter plots for six representative metric pairs (grey points) are shown.
-Each plot includes an independent power-law fit to the empirical data (solid red line, with 95% CI) and the prediction derived from the joint model (dashed black line, Eq.
-A3), demonstrating high correspondence.
-(D) Latent similarity map.
-The cortical surface map displays the estimated latent correlation values, $\hat{\rho}^{\mathrm{latent}}_{r}$ (HCP-MMP1 atlas, symmetric parcels).
-The model reveals major hubs of high latent similarity in early visual cortex, ventral temporal cortex, and LOTC, with pFC becoming more prominent after accounting for metric-specific adjustment.
-Together, these panels demonstrate that a single latent geometry, subject to metric-specific power-law attenuation, effectively explains the relationships among all measured RSA metrics.
-Figure A8: Global power-law model captures every pair-wise relation among RSA metrics.
-Scatter plots (grey points) show parcel-wise observed RSA values for all metric combinations: inter-subject (shifted and unshifted), within-subject, subject-versus-group, and vision-model correlations, each additionally split by hemisphere sampling (left, right, both).
-Red curves depict an independent power-law fit for the specific scatter, whereas black dashed curves are the predictions from a single joint fit of Eq.
-(A3) applied to all metrics simultaneously (no panel–specific re–fitting).
-The accompanying table reports the fitted scale factors $a_{j}$, exponents $b_{j}$, and explained variance $R^{2}$ for every metric (model-predicted vs observed rsa).
-All metric trends are explained with $R^{2}\!>\!0.92$, confirming that one latent similarity map, together with metric-specific attenuation parameters, suffices to account for the entire measurement set.
-B THINGS-fMRI and BOLD5000 Processing
-B.1 THINGS-fMRI dataset
-We processed the THINGS-fMRI dataset 25 using a pipeline analogous to that for NSD.
-This dataset comprises 3T fMRI recordings from three participants performing an oddball detection task on 8,740 unique object images.
-A subset of 100 images was repeated once per session (12 presentations total); all other images appeared only once.
-For our primary computations, we used only the first presentation of each of the 8,740 images.
-Pre-processed, ICA-regressed single-trial estimates were obtained from the authors’ public repository. We extracted voxel responses into the HCP-MMP atlas parcels provided by the THINGS team.
-To control for within-session variability, inter-subject alignment was computed on a session-wise basis.
-For any pair of participants $(p,q)$ and parcels $(r,r^{\prime})$, we computed IS-RSA by correlating the RDMs generated from the images within a given session $s$:
-$\mathcal{A}_{p,q,r,r^{\prime},s}^{\mathrm{IS}}=\mathrm{RSA}\bigl{(}X_{p,r,s},X_{q,r^{\prime},s}\bigr{)}.$ (B4)
-Group-level maps were then obtained by averaging these scores across sessions and all participant pairs.
-Model-brain RSA followed the same session-wise procedure.
-As a control, we confirmed that computing IS-RSA over the full 8,740-trial set (using $,740\times 8$ 740 × 8740 RDMs) yielded parcel-wise results that were almost perfectly correlated with our session-wise computation (Pearson’s $r=0.995$).
-While the full-set computation produced systematically lower alignment values (linear slope $\approx 0.53$; $R^{2}=0.99$), this correlation validates that our session-wise approach preserves the relative spatial pattern of the findings.
-B.2 BOLD5000 dataset
-We processed the BOLD5000 dataset 24 using the same session-wise pipeline as for THINGS-fMRI.
-This dataset comprised 3T recordings from four participants performing a valence rating task on 4,916 unique images drawn from the MS-COCO, ImageNet, and SUN databases.
-We used the pre-processed, single-trial $\beta$-estimates provided by the authors, selecting only the first presentation of each image.
-Inter-subject and model-brain RSA were computed following the same session-wise procedure described above.
-For voxel extraction, we registered the HCP-MMP1.0 atlas from MNI152 standard space to each subject’s native T1-weighted scan and then to their mean functional volume, using FSL’s FLIRT and FNIRT.
-While qualitative inspection confirmed consistent parcel definitions across participants, this volume-to-volume registration has a lower effective resolution compared to the specialized registration available for the NSD atlas.
-This may account for some parcel-specific results, such as the lower-than-expected alignment values observed in certain regions (e.g., MST).
-All other preprocessing and computational steps paralleled the main NSD pipeline.
-C Implementation details for large-scale RSA comparisons
-To efficiently handle the computational demands associated with large-scale RSA analyses (due to dataset size and permutation complexity), we developed a GPU-accelerated implementation using PyTorch84.
-Specifically, we reformulated representational dissimilarity matrix (RDM) calculations as optimized matrix operations.
-For each participant pair $p,q\in\mathcal{P}$, and each region $r\in\mathcal{R}$, we constructed stacked matrices consisting of flattened RDMs across all cortical regions, significantly reducing computational overhead and runtime compared to conventional RSA implementations.
-Given data matrices for matched trials $X_{p,r}[\boldsymbol{i}_{pq},:]$ and $X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]$, we computed upper-triangular vectors of dissimilarities:
-$\mathbf{x}^{(p)}_{pq,r}=\operatorname{vec}_{u}D\left(X_{p,r}[\boldsymbol{i}_{pq},:]\right),\quad\mathbf{x}^{(q)}_{pq,r^{\prime}}=\operatorname{vec}_{u}D\left(X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]\right),$ (C5)
-where $D(\cdot)$ denotes the dissimilarity matrix and $\operatorname{vec}_{u}$ denotes vectorization of the upper triangle.
-Each RDM vector was centered and $l_{2}$-normalized before stacking into a matrix containing all ROIs:
-$\mathbf{X}^{(p)}_{pq}=\begin{bmatrix}\tilde{\mathbf{x}}^{(p)}_{pq,1}\\ \tilde{\mathbf{x}}^{(p)}_{pq,2}\\ \vdots\\ \tilde{\mathbf{x}}^{(p)}_{pq,|\mathcal{R}|}\end{bmatrix}\quad\in\mathbb{R}^{|\mathcal{R}|\times\frac{n(n-1)}{2}},$ (C6)
-and similarly for $\mathbf{X}^{(q)}_{pq}$.
-The inter-subject connectivity matrix was then obtained as:
-$\mathcal{A}^{\mathrm{IS}}_{p,q}[r,r^{\prime}]=\mathbf{X}^{(p)}_{pq}\left(\mathbf{X}^{(q)}_{pq}\right)^{\top}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{R}|}$ (C7)
-Analogous formulations were applied for Spearman-based RSA, with rank-normalization applied to RDM vectors prior to computing dot products.
-For brain–model alignment, similar stacked matrices were constructed separately for brain RDMs and model RDMs across sessions. The brain matrix was organized as:
-$\mathbf{X}^{(\text{brain})}_{p}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{T}_{p}|\times\frac{n(n-1)}{2}},$ (C8)
-and the model matrix as:
-$\mathbf{X}^{(\text{model})}_{m}\in\mathbb{R}^{|\mathcal{L}_{m}|\times|\mathcal{T}_{p}|\times\frac{n(n-1)}{2}}.$ (C9)
-Brain–model RSA alignments across regions, layers, and sessions were computed as the inner product between normalized RDM vectors for each session:
-$\mathcal{C}_{p,r,m,s}(l)=\langle\mathbf{X}^{(\text{brain})}_{p,r,s},\;\mathbf{X}^{(\text{model})}_{m,l,s}\rangle.$ (C10)
-This operation was implemented as a contraction corresponding to the einsum notation $r\,s\,d,l\,s\,d\to r\,l$.
-For permutation testing, we permuted stimulus labels defined by the permutation $\sigma\in S_{n}$ prior to RSA computation.
-Rather than recomputing dissimilarity matrices after each permutation, we constructed an equivalent permutation $\sigma^{\prime}$ over the vectorized upper-triangular RDM entries, matching the induced pairwise index shifts.
-This allowed permutation of the flattened RDMs directly via:
-$\mathbf{X}^{(p)}_{pq}[:,\sigma^{\prime}],$ (C11)
-preserving the correct stimulus-pair structure with minimal computational overhead.
-D Extended figures
-Figure D1: Detailed parcel-level alignment for all modalities.
-Box plots show representational alignment scores computed for each of the 180 cortical parcels of the symmetric HCP atlas 31, organized by macro-anatomical groups 47.
+Extended figures
+Figure 1: Detailed parcel-level alignment for all modalities.
+Box plots show representational alignment scores computed for each of the 180 cortical parcels of the symmetric HCP atlas 34, organized by macro-anatomical groups 47.
 Each box represents the distribution of alignment scores across the eight participants ($N=8$).
@@
 This figure provides a detailed view of the results summarized in the main text, showing the parcel-by-parcel variability and confirming the concentration of high alignment within the Early Visual, Ventral, and LOTC hubs.
-Figure D2: Cortical flat map projections of representational alignment.
+Figure 2: Cortical flat map projections of representational alignment.
 To provide a comprehensive view of their spatial distribution, group-level alignment scores are projected onto a flattened cortical surface.
-The figure shows (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment.
+The figure shows (A) IS-RSA alignment, (B) vision-model alignment, and (C) language-model alignment.
 Values represent group-averaged RSA scores (Pearson’s $\rho$, $N=8$ from the NSD dataset), and major sulci are labeled for anatomical orientation.
 This visualization makes the full extent of the alignment patterns clear, particularly highlighting the widespread negative alignment (cool colors) between language models and the ventral visual stream, in contrast to the positive alignment seen for inter-subject and vision-model comparisons.
-Figure D3: Detailed layer-wise alignment profiles across cortical parcels.
-This figure provides a detailed, parcel-by-parcel view of the brain-model alignment data summarized in the main text (Fig.
-3), showing results for the 15 macro-anatomical groups with the highest inter-subject alignment.
-Only parcels that were themselves statistically significant in the inter-subject analysis are displayed.
-For each modality, the plotted alignment curves represent the RSA score averaged across all models of that type.
-Each individual line shows the mean alignment for a single parcel across participants ($N=8$), with shaded areas indicating the SEM.
-(A) Brain-to-vision-model alignment.
-This detailed view confirms that the distinct alignment profiles—decreasing, distributed, and increasing—are characteristic of the main hubs and often extend to anatomically adjacent parcels (e.g., the increasing profile of the LOTC hub is also observed in nearby STS regions).
-In contrast, areas with lower overall correspondence, such as in the prefrontal cortex, tend to show weak alignment across all model layers without a clear hierarchical preference.
-(B) Brain-to-language-model alignment.
-This view confirms that positive alignment is restricted to the LOTC hub and a small number of anatomically proximal parcels, all of which consistently exhibit the same step-like alignment profile.
-Figure D4: Brain-model alignment profiles for different model families.
-Layer-wise alignment (RSA, Pearson’s $r$) is shown for representative parcels from the three main hubs, computed separately for each family of models (see Table 1 for a full list).
-Each curve represents the mean alignment across participants ($N=8$), with shaded areas indicating the SEM.
-Vision Models: Most vision model families (AugReg, CLIP, DINOv2) reproduce the characteristic hierarchical profiles: a decreasing alignment with Early Visual Cortex, a distributed alignment with the Ventral Hub, and an increasing alignment with the LOTC Hub.
-The Masked Autoencoder (MAE) models are a notable exception; consistent with their reconstructive training objective, they maintain a high alignment with Early Visual Cortex across all layers.
-Language Models: Despite differences in architecture, training data, and fine-tuning (e.g., instruction tuning in BLOOMZ), all language model families exhibit a similar pattern.
-Alignment is negligible or negative for Early Visual and Ventral hub parcels, while all families show the same characteristic step-function profile for parcels in the LOTC hub.
-Figure D5: Comparison of within-subject and inter-subject representational analyses.
+Figure 3: Comparison of within-subject and inter-subject representational analyses.
 (A) Cortical map of within-subject alignment (RSA, Pearson’s $r$).
@@
 Computed by correlating RDMs between all pairs of parcels within each participant, this analysis reveals the same three-hub block structure discovered using the inter-subject measurements.
-(D) Directed connectivity graph of the within-subject data, pruned using a 2-minimum spanning tree to highlight the network backbone.
+(D) Directed connectivity graph of the within-subject data, pruned using a 3-minimum spanning tree to highlight the network backbone.
 Edge color and width encode the inter-region alignment (RSA value), while node color represents each parcel’s peak alignment depth with vision models.
@@
 The graph confirms the same information flow as the inter-subject analysis, with two streams emerging from early visual cortex: a medial-ventral stream to the Ventral hub and a lateral-dorsal stream to the LOTC hub.
-Figure D6: Whole-cortex representational connectivity network.
+Figure 4: Detailed layer-wise alignment profiles across cortical parcels.
+Detailed parcel-by-parcel view of the brain-model alignment curves summarized in the main text (Fig.
+3), showing results for the 15 macro-anatomical groups with the highest inter-subject alignment.
+Only parcels that were themselves statistically significant in the inter-subject analysis are displayed.
+For each modality, the plotted alignment curves represent the RSA score averaged across all models of that type.
+Each individual line shows the mean alignment for a single parcel across participants ($N=8$), with shaded areas indicating the SEM.
+(A) Brain-to-vision-model alignment.
+This detailed view confirms that the distinct alignment profiles–decreasing, distributed, and increasing–are characteristic of the main hubs and often extend to anatomically adjacent parcels (e.g., the increasing profile of the LOTC hub is also observed in nearby STS regions).
+In contrast, areas with lower overall correspondence, such as in the prefrontal cortex, tend to show weak alignment across all model layers without a clear hierarchical preference.
+(B) Brain-to-language-model alignment.
+This view confirms that positive alignment is restricted to the LOTC hub and a small number of anatomically proximal parcels, all of which consistently exhibit the same step-like alignment profile.
+Figure 5: Brain-model alignment profiles for different model families.
+Layer-wise alignment (RSA, Pearson’s $r$) is shown for representative parcels from the three main hubs, computed separately for each family of models (see Table 1 for a full list).
+Each curve represents the mean alignment across participants ($N=8$), with shaded areas indicating the SEM.
+Vision Models: Most vision model families (AugReg, CLIP, DINOv2) reproduce the characteristic hierarchical profiles: a decreasing alignment with Early Visual Cortex, a distributed alignment with the Ventral Hub, and an increasing alignment with the LOTC Hub.
+The Masked Autoencoder (MAE) models are a notable exception; consistent with their reconstructive training objective, they maintain a high alignment with Early Visual Cortex across all layers.
+Language Models: Despite differences in architecture, training data, and fine-tuning (e.g., instruction tuning in BLOOMZ), all language model families exhibit a similar pattern.
+Alignment is negligible or negative for Early Visual and Ventral hub parcels, while all families show the same characteristic step-function profile for parcels in the LOTC hub.
+Figure 6: Noise-ceiling normalization of IS-RSA representational connectivity.
+(A) IS-RSA representational connectivity matrix (Pearson’s $\rho$) for the 30 parcels with highest inter-subject alignment.
+Black squares indicate, for each parcel (column), the parcel with which it shows maximal between-parcel IS-RSA connectivity.
+(B) Backbone graph derived from (A) using three iterations of a minimum-spanning-tree procedure.
+(C) Noise-ceiling–normalized connectivity matrix, obtained by dividing each edge by the geometric mean of the corresponding diagonal self-alignment estimates, setting the diagonal to 1.
+(D) Backbone graph extracted from the normalized matrix in (C).
+The preserved topology indicates that the inferred dual-route architecture is robust to parcel-wise differences in measurement reliability.
+Figure 7: Whole-cortex representational connectivity network.
 Representational connectivity network among the 157 cortical parcels that exhibited statistically significant inter-subject alignment.
@@
 Furthermore, despite weaker overall alignment, frontal and parietal regions like the Frontal Eye Fields (FEF) and the Medial Intraparietal Area (MIP) emerge as distinct hubs, suggesting their integration into this large-scale, stimulus-driven network.
-Figure D7: Biological content selectively drives LOTC alignment and lateral-stream connectivity.
+Figure 8: Biological content selectively drives LOTC alignment and lateral‐stream connectivity.
 Scenes were split into those that contained biological agents (people or animals; 65.3% of the shared image set) and those that did not, and all inter-subject analyses were recomputed on each subset.
@@
 The lateral stream linking Early Visual Cortex to LOTC is largely absent for non-biological scenes (B) but re-emerges when biological agents are present (C), mirroring the parcel-level results in (A).
-Figure D8: Image-based atlas of the dominant representational dimensions in Early Visual Cortex.
+Figure 9: Image-based atlas of the dominant representational dimensions in Early Visual Cortex.
 The first two shared representational components for the Early Visual Cortex hub, derived using KMCCA on the responses of all participants to the shared image set.
@@
 Instead, images that are close to each other in this shared space often have similar low-level visual properties (e.g., color palettes, textures, or global shapes), confirming that the geometry of this hub is primarily driven by visual similarity rather than abstract content.
-Figure D9: Image-based atlas of the dominant representational dimensions in the Ventral Hub.
+Figure 10: Image-based atlas of the dominant representational dimensions in the Ventral Hub.
 The first two shared representational components for the Ventral hub, derived using KMCCA.
@@
 This evidences that the dominant representational axis of this hub recapitulates the classic scene-to-object gradient of ventromedial cortex.
-Figure D10: Image-based atlas of the dominant representational dimensions in the LOTC Hub.
+Figure 11: Image-based atlas of the dominant representational dimensions in the LOTC Hub.
 The first two shared representational components for the LOTC hub, derived using KMCCA.
@@
 This confirms that the LOTC hub’s geometry is primarily organized around the presence and type of animate entities in a scene.
+Supplementary Notes
+Figure 12: Repetition-matching strategies and their impact on inter-subject RSA.
+To compute inter-subject RSA, we compared different strategies for matching the three trial repetitions for each shared image between any pair of participants.
+(A) Schematic of unshifted (repetition 1→1, 2→2, 3→3) versus shifted (cyclic 1→2, 2→3, 3→1) trial matching with example RDMs.
+(B, C)Group-average cortical RSA maps under unshifted (B) and shifted (C) matching; both reveal the early-visual, ventral, and LOTC hubs.
+(D) Parcel-wise comparison of shifted vs.
+unshifted RSA values shows a uniform attenuation under shifted matching, with two linear regimes reflecting stronger prefrontal attenuation.
+Supplementary Note 1: Repetition-Matching Strategies for Inter-Subject RSA
+The Natural Scenes Dataset (NSD) presents 10,000 unique images across up to 40 scanning sessions per participant (30,000 trials total).
+A subset of 1,000 images (3,000 trials) was shared by all eight participants and always appeared in the same trial positions–interleaved among participant‐unique stimuli–so that each shared image was viewed under identical practice, fatigue, and session‐context conditions.
+This locked structure raises the possibility that inter‐subject correlations might reflect non‐perceptual factors (e.g., trial timing or memory demands) rather than purely stimulus‐driven geometry.
+To isolate the stimulus‐driven component, we adopted a “shifted‐repetition” matching strategy when computing inter‐subject representational similarity (IS‐RSA).
+Rather than pairing each repetition index $k\leftrightarrow k$ across participants (“unshifted”), we cyclically permute so that repetition $k$ in participant $p$ matches repetition $(k\bmod 3)+1$ in participant $q$ (i.e.
+$1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$; see Fig.
+12A).
+Because all shared trials occupy the same session slots, this shift preserves image identity while breaking any exact repetition‐locked confounds and also allows direct comparison with within‐subject analyses (which must compare different repetitions; see Supplementary Fig.
+3).
+Due to symmetry across participant‐pair comparisons, shifts of 1 and 2 yield identical group‐average maps.
+Recomputing the group‐average cortical RSA under both unshifted (Fig.
+12B) and shifted (Fig.
+12C) schemes recovers the same three hubs–early visual cortex, ventral, and LOTC.
+As expected, shifting is more conservative: alignment magnitudes are uniformly lower (Supplementary Fig.
+12D), yet the relative rank‐order of all parcels remains unchanged (Spearman’s $\rho=0.96$).
+Frontal parcels (e.g.
+FEF, inferior frontal gyrus) suffer proportionally larger attenuation under shifting than occipito‐temporal hubs, suggesting that these frontal regions encode task‐ or repetition‐related variance (e.g.
+decision or attentional control) that the shifted criterion minimizes.
+By contrast, the robust persistence of occipital and temporal alignments underscores their stable, stimulus‐driven representational geometry.
+To illustrate how these differences manifest in network structure, Fig.
+13 shows representational connectivity graphs for both matching strategies.
+We first identified parcels whose inter-subject RSA exceeded an absolute threshold of $r=0.05$ in either the shifted (shift = 1 or 2) or unshifted maps.
+We then constructed a reduced connectivity matrix among this common set of nodes and extracted a two-iteration minimum‐spanning‐tree backbone (using edge weights defined as $1-\mathrm{RSA}$; see main Methods).
+The core two-stream topology (Early Visual $\rightarrow$ Ventral hub and Early Visual $\rightarrow$ LOTC hub) is preserved under both schemes.
+However, only the unshifted graph highlights FEF as a high-centrality node, further suggesting that frontal regions carry shared task- or repetition-locked signals that are attenuated by the shifted criterion.
+While these frontal effects are exploratory–given their lower RSA magnitudes and sensitivity to thresholding–they point to important future work on how task and attention signals coexist with stimulus‐driven representations.
+[…]
+Supplementary Note 2: Parcel size and the reliability of IS-RSA estimates
+HCP–MMP parcels vary substantially in surface area and voxel count, raising the concern that IS-RSA estimates might be systematically less stable in smaller parcels due to reduced signal averaging.
+To evaluate this, we quantified the dispersion of IS-RSA estimates under two resampling schemes that isolate distinct sources of variability: changes in stimulus sampling and trial-level measurement noise.
+To assess stability with respect to stimulus sampling, we recomputed IS-RSA across 1,000 random half-splits of the image set (500 unique images per split, including all repetitions; Supplementary Fig.
+14a-d).
+For each parcel, we computed the coefficient of variation (CV; standard deviation of the distribution divided by the mean) as a scale-normalized index of estimation stability.
+To isolate trial-level measurement variability while holding stimulus content fixed, we performed 1,000 iterations in which we randomly selected a single repetition per image (from the 2–3 available) and recomputed IS-RSA (Supplementary Fig.
+14e-h).
+Dispersion was again summarized per parcel using the CV.
+Across both analyses, parcels with similar mean IS-RSA exhibited comparable dispersion despite large differences in voxel count (Supplementary Fig.14a-b, e-f).
+Critically, CV showed no systematic dependence on parcel size (Supplementary Fig.14c,g).
+Instead, dispersion decreased monotonically with the magnitude of IS-RSA (Supplementary Fig.
+14d,h), indicating that parcels with stronger shared geometry yield more stable IS-RSA estimates regardless of their size.
+These results suggest that, under the NSD preprocessing regime, the reliability of IS-RSA estimates is primarily governed by the strength of the shared representational signal rather than by the number of voxels available for averaging.
+Figure 14: Quantifying parcel reliability via resampling.
+(A–D) Stimulus generalization analysis.
+We computed IS-RSA on 1,000 random half-splits of the stimulus set (selecting 500 stimuli in each subsample).
+(A,B) Distribution of IS-RSA scores for a representative large parcel (V1) and small parcel (VMV2).
+Dashed vertical lines indicate the IS-RSA estimate from the full dataset.
+(C) IS-RSA coefficient of variation (CV; std/mean) plotted against parcel voxel count.
+(D) CV plotted against mean IS-RSA alignment.
+(E–H) Measurement reliability analysis based on randomly selecting one repetition per image across 1,000 iterations.
+(E,F) Corresponding distributions for V1 and VMV2.
+(G) CV versus voxel count.
+(H) CV versus mean IS-RSA alignment.
+In both analyses, the dispersion shows no dependence on voxel count but decreases as a function of the IS-RSA magnitude.
+Supplementary Note 3: Hemispheric Asymmetry and the Validity of Symmetric Analyses
+Figure 15: Hemispheric comparison of representational alignment.
+(A-C) Comparison of representational alignment (RSA, Pearson’s $r$; $N=8$, NSD) computed independently for the left (x-axis) and right (y-axis) hemispheres: (A) inter-subject alignment, (B) vision-model alignment, and (C) language-model alignment.
+Each point is a cortical parcel, colored by macro-anatomical group.
+The diagonal line indicates equal alignment.
+Squares denote parcels with a statistically significant hemispheric difference (paired $t$-test, $p<0.05$, FDR-corrected).
+(D) Boxplots of alignment within principal hubs.
+(E) Correlation between cross-hemispheric and bilateral inter-subject alignment across parcels.
+Significant hemisphere effects are marked.
+We quantified hemispheric asymmetry in representational geometry by computing alignment measures separately for the left and right hemispheres using the HCP-MMP1.0 atlas 34.
+For each cortical parcel, we assessed inter-subject alignment, vision-model alignment, and language-model alignment using the same NSD responses as in the main analyses (Fig.
+15).
+Across all modalities, a largely symmetric pattern was observed (Fig.
+15a-c).
+The rank-order of parcels was highly consistent between hemispheres (Spearman’s $\rho$: inter-subject =0.93=0.93, vision =0.92=0.92, language =0.82=0.82).
+However, specific asymmetries emerged: early visual cortex (V1–V4) consistently showed stronger alignment in the left hemisphere, while the ventral and LOTC hubs (e.g., hMT+, LO, TPOJ) showed stronger alignment in the right hemisphere (Fig.
+15d).
+These lateralization trends were stable across inter-subject and model-based analyses.
+To determine whether these asymmetries reflected fundamental differences in representational geometry or merely differences in signal strength, we compared cross-hemispheric RSA (e.g., subject $p$’s left vs subject $q$’s right) to bilateral RSA.
+These measures were near-perfectly correlated across parcels ($r=0.99$, Fig.
+15e), with no parcel deviating substantially from the regression line.
+This demonstrates that combining hemispheres increases overall alignment strength (likely via noise averaging) without distorting the underlying representational geometry; no region exhibited a lateralized geometry that was absent in the contralateral hemisphere.
+Together, these results validate our strategy of pooling hemispheres for the main analyses: aggregation maximizes statistical power for multivariate comparisons without introducing representational artifacts, while still allowing for targeted lateralization analyses where relevant.
+Supplementary Note 4: Power-Law Attenuation Model of RSA Metrics
+In the main text, we compared several RSA measures–including within-subject (WS), inter-subject (IS-1: shift=1 and IS-0: unshifted), and vision-model-brain (VM)–that probed a common cortical geometry but were subject to different levels of measurement noise.
+While all metrics revealed a similar cortical alignment pattern, their absolute magnitudes varied considerably.
+Parcel-wise scatter plots (e.g., Fig.
+2e; Fig.
+12) showed a systematic relationship between the metrics, which were tightly related by power-law curves.
+Although these curves differed in scale and curvature, their consistent functional form suggested that each metric captured the same latent similarity structure while imposing a unique, noise-dependent attenuation.
+Figure 16: Power-law attenuation model.
+(A) Goodness-of-fit.
+Parcel-wise observed RSA values (y-axis) plotted against predictions from the joint power-law model (Eq.
+1; x-axis).
+Colors denote metrics (WS: within-subject; IS-1: inter-subject shifted; IS-0: inter-subject unshifted; VM: vision model).
+Legend indicates explained variance ($R^{2}$).
+(B) Attenuation curves.
+Estimated latent correlation ($\hat{\rho}^{\mathrm{latent}}$, x-axis) vs.
+observed correlation (y-axis).
+Colored lines show the fitted metric-specific attenuation curves ($y=a_{j}x^{b_{j}}$).
+(C) Pairwise comparisons.
+Scatter plots for six representative metric pairs (gray points).
+Each plot includes an independent power-law fit (solid red line, 95% CI) and the prediction from the joint model (dashed black line, Eq.
+2), demonstrating high correspondence.
+(D) Latent similarity map.
+Cortical surface map of the estimated latent correlation values, $\hat{\rho}^{\mathrm{latent}}_{r}$ (HCP-MMP1 atlas, symmetric parcels).
+The model recovers the three major hubs (EVC, Ventral, LOTC) and suggests distinct signal in prefrontal regions after accounting for metric-specific attenuation.
+We formalize this observation by modeling the correlation measured for parcel $r$ with metric $j$ as a power-law transform of a latent correlation:
+$\rho^{\mathrm{obs}}_{rj}=a_{j}\bigl(\rho^{\mathrm{latent}}_{r}\bigr)^{\,b_{j}}+\epsilon_{rj},$ (1)
+where:
+  • $\rho^{\mathrm{latent}}_{r}\in[0,1]$ is the metric-independent similarity common to all measures;
+  • $a_{j}\in[0,1]$ is a linear attenuation factor for metric $j$;
+  • $b_{j}\geq 1$ captures non-linear signal compression.
+A value of $b_{j}=1$ corresponds to simple linear attenuation, whereas $b_{j}>1$ creates a convex curve that suppresses weak latent correlations more strongly than strong ones.
+This constraint is motivated by the observation that noisier metrics (e.g., IS-shift) systematically underestimate weakly aligned parcels relative to cleaner metrics (e.g., WS), consistent with a noise regime that disproportionately obscures weaker signals;
+  • $\epsilon_{rj}$ is a residual term absorbing parcel–metric–specific variability not captured by the deterministic power-law component.
+We jointly fitted the parameters of Eq.
+1 to four core metrics (WS, IS-0, IS-1, VM) using data from the symmetric-hemisphere ROIs.
+Parameter estimation was performed via the L-BFGS-B algorithm by minimizing the sum of squared residuals, with bounds: $a_{j}\in[0,1]$, $b_{j}\in[1,\infty)$, and $\rho^{\mathrm{latent}}_{r}\in[0,1]$.
+Figure 16A demonstrates the model’s high goodness-of-fit.
+In all cases, the model explained over 94% of the variance ($R^{2}>0.94$), confirming that the power-law formulation accurately captures the relationship between observed data and the latent construct.
+Figure 16B visualizes the distinct attenuation profile for each RSA strategy, illustrating how each metric compresses the estimated latent correlation.
+Because all metrics are modeled as functions of the same latent variable, $\rho^{\mathrm{latent}}_{r}$, the relationship between any two metrics, $j$ and $j^{\prime}$, can be derived analytically by eliminating this common term:
+$y=\frac{a_{j^{\prime}}}{a_{j}^{\,b_{j^{\prime}}/b_{j}}}\;x^{\,b_{j^{\prime}}/b_{j}},$ (2)
+where $x=\rho^{\mathrm{obs}}_{rj}$ and $y=\rho^{\mathrm{obs}}_{rj^{\prime}}$.
+Figure 16C validates this prediction by overlaying the derived curve (Eq.
+2) on the empirical parcel scatters.
+The tight alignment confirms that a single latent similarity, subject to metric-specific attenuation, explains the pairwise relationships.
+For completeness, Figure 17 extends this validation to a set of 12 metrics, demonstrating the model’s ability to unify all RSA strategies used in our analyses.
+The resulting cortical map of the latent variable, $\hat{\rho}^{\mathrm{latent}}_{r}$, is displayed in Figure 16D.
+This map represents a attenuation-corrected estimate of similarity that unifies the core RSA measures.
+It recovers the three-hub pattern observed in the main text, with peak correlations ($\rho^{\mathrm{latent}}_{r}\sim 0.50$) in the EVC, Ventral stream, and LOTC.
+Furthermore, the model suggests that frontal areas (e.g., FEF and PFC) become more prominent after correcting for the attenuation present in raw measurements.
+The fact that these hubs align with those highlighted by individual metrics confirms that Eq.
+1 successfully isolates a common signal while factoring out metric-specific noise.
+Note that because Eq.
+1 is scale-free, the absolute values of $\hat{\rho}^{\mathrm{latent}}$ and $a_{j}$ are defined only up to a common scaling factor.
+We fixed no anchor during optimization, yet repeated fits converged to the same scale, and all qualitative findings (relative parcel ranking, cross-metric relations) are invariant under rescaling.
+The power-law attenuation model offers a compact, quantitative account of how measurement differences degrade RSA scores.
+Its success supports the view that all metrics probe a shared latent representational geometry, differing only in the fidelity with which each metric samples that geometry.
+Figure 17: Global power-law model captures pairwise relations among all RSA metrics.
+Scatter plots (gray points) show parcel-wise observed RSA values for all metric combinations: inter-subject (shifted and unshifted), within-subject, subject-versus-group, and vision-model correlations, each additionally split by hemisphere sampling (left, right, both).
+Red curves depict an independent power-law fit for that specific scatter, whereas black dashed curves are the predictions from a single joint fit of Eq.
+2 applied to all metrics simultaneously (no panel-specific re-fitting).
+The accompanying table reports the fitted scale factors $a_{j}$, exponents $b_{j}$, and explained variance $R^{2}$ for every metric (model-predicted vs observed RSA).
+All metric trends are explained with $R^{2}>0.92$, confirming that one latent similarity map, together with metric-specific attenuation parameters, suffices to account for the entire measurement set.
+Supplementary Note 5: Token‐occurrence control for language–brain RSA
+To quantify how much of the language–brain alignment reflects initial tokenization, we constructed RDMs from token-occurrence vectors.
+Each caption was encoded as a count vector—each entry recording the number of times a given token appears—and pairwise Pearson dissimilarities among these vectors yielded a tokenizer-based RDM.
+We applied this procedure to six vocabularies (BLOOMZ, Gemma-2, LLaMA-1/2/3; see main text Table 1), covering all tokenizers used in our language models, and then computed parcel-wise RSA against NSD fMRI responses using the symmetric HCP atlas.
+Parcel-wise alignments were nearly identical across tokenizers (pairwise Spearman’s $\rho\geq 0.978$), so we report their average.
+The tokenizer–brain alignment map closely matches that of full language models: it peaks in the LOTC hub and shows minimal or negative alignment in early visual parcels (Suppl.
+Fig.
+18A,D).
+Token-count geometry correlates strongly with each model’s first-layer alignment (Pearson’s $\rho=0.78$; Suppl.
+Fig.
+18B) and with maximal alignment across layers ($\rho=0.68$; Suppl.
+Fig.
+18C), indicating that token frequencies account for a substantial portion of the language–brain correspondence.
+Repeating the analysis with a word tokenizer—splitting captions on whitespace rather than into subword units—again yielded pronounced LOTC alignment (Suppl.
+Fig.
+18E).
+These findings suggest that LOTC’s alignment to language models largely reflects sensitivity to discrete lexical tokens common across descriptions—functioning as a “bag-of-words” representation—rather than to compositional or syntactic structure 100.
+Figure 18: Tokenizer‐level contributions to language–brain alignment.
+We derived RDMs from token‐occurrence count vectors over each model’s vocabulary and computed parcel‐wise RSA against NSD fMRI responses (averaged across the six tokenizers used in our language models).
+(A) Scatter of tokenizer‐brain vs.
+inter‐subject alignment, with peak values in the LOTC hub.
+(B) Token‐count alignment against each model’s first‐layer alignment.
+(C) Token‐count alignment against each model’s maximal alignment across layers.
+(D) Cortical map of tokenizer‐brain alignment, showing pronounced LOTC localization.
+(E) Alignment map for a whitespace‐based word tokenizer, which likewise peaks in LOTC, consistent with a “bag‐of‐words” explanation rather than richer compositional structure.
+Supplementary Note 6: Robustness to Similarity Metric (Spearman‐RSA & CKA)
+Figure 19: Comparison of alignment metrics.
+(A–C) Spearman‐RSA.
+Replacing Pearson’s $r$ with Spearman’s rank correlation yields virtually identical maps: the inter‐subject alignment (A), the parcel‐wise scatter against vision‐ and language‐model alignment (B), and the representational connectivity matrix (C) all match the Pearson‐RSA findings, with parcel‐wise strengths correlating at $r=0.99$.
+(D–F) Unbiased CKA.
+Centered Kernel Alignment, which captures global dependence via an unbiased HSIC estimator, again highlights Early Visual, Ventral, and LOTC hubs (D).
+In the parcel‐wise comparison (E), CKA inflates LOTC alignment and attenuates early visual alignment–reflecting its sensitivity to low‐dimensional, dominant components–yet CKA and Pearson‐RSA remain strongly correlated ($r=0.91$).
+The CKA‐based connectivity matrix (F) preserves the two‐stream topology, though as a dependence measure it does not distinguish signed similarity.
+Recent work has cautioned that representational similarity findings can depend not only on stimulus design but also on the choice of similarity metric—some measures probe local, pairwise geometry while others emphasize global, space-wide dependence 25.
+To test the robustness of our results to these methodological choices, we repeated key analyses using two alternatives to Pearson-RSA: (i) rank-based Spearman-RSA, which preserves only the ordering of pairwise dissimilarities, and (ii) Centered Kernel Alignment (CKA), which quantifies global statistical dependence via an unbiased Hilbert-Schmidt Independence Criterion (HSIC) estimator 83.
+Replacing Pearson with Spearman correlation yields a purely rank-based, nonparametric measure agnostic to absolute distance magnitudes.
+As shown in Supplementary Fig.
+19A–C, the inter-subject map, the parcel-by-parcel scatter against model alignment, and the representational connectivity matrix all mirror our original Pearson-RSA results (parcel-wise Spearman’s $\rho=0.99$).
+This confirms that neither the linearity nor the distributional assumptions of Pearson correlation dictate our findings—our local pairwise geometry holds under a fully nonparametric test.
+CKA measures dependence across entire representational spaces, prioritizing alignment along principal axes rather than isolated pairwise distances.
+We computed linear-kernel CKA (using the unbiased HSIC estimator 83) under the same shifted-repetition protocol.
+Supplementary Fig.
+19D–F shows that CKA again recovers the three hubs but inflates LOTC alignment relative to Early Visual cortex in the parcel scatter.
+This reflects CKA’s known sensitivity to low-dimensional, dominant components (often stronger in higher-level cortex) over the high-dimensional, finer-grained structure of early vision.
+Despite this spectral re-weighting, parcel-wise CKA and Pearson-RSA correlate strongly ($r=0.91$), and the CKA connectivity matrix preserves the core two-stream topology, even though CKA is strictly non-negative and thus cannot index anticorrelated geometries.
+Supplementary Note 7: Shared Component Decomposition and Partial RSA Controls
+Figure 20: Extended analysis of representational dimensions.
+(A) KMCCA projections onto the top two canonical axes for six representative parcels: V1, V4, VMV2, PHA2, MT, and TPOJ2.
+V1 shows no clear clustering; V4 begins to separate by content; Ventral and LOTC parcels replicate the hub-level semantic gradients.
+(B) Words ranked by loading on the first KMCCA component (positive loadings in blue, negative in red) for each hub.
+Early Visual words share low-level shape features; Ventral words span scene vs.
+object terms; LOTC words distinguish animate vs.
+inanimate labels.
+(C) Partial RSA boxplots for the three hubs, showing inter-subject alignment under no control (baseline), controlling for the scene-object gradient, or biological presence.
+LOTC alignment is markedly reduced by each control, whereas Early Visual and Ventral hubs are minimally affected.
+This supplementary note details our use of Kernel Multi-view Canonical Correlation Analysis (KMCCA) 40 to isolate the axes driving our RSA findings.
+While RSA provides a single global similarity score, KMCCA decomposes this correspondence into a ranked set of orthogonal “representational channels,” revealing which latent axes contribute most strongly to the shared geometry.
+We performed KMCCA using the Python package mvlearn (v0.5.0) 68 with a correlation kernel and a regularization parameter of $\lambda=0.3$, utilizing the implementation’s default parameters.
+For each cortical hub, we formed a data matrix for each participant from the GLM-denoised single-trial $\beta$-estimates.
+We included only those trials for which each (image, repetition) pair was available for all eight participants.
+This procedure yielded a set of eight matrices ($X_{p,r}\in\mathbb{R}^{m\times v_{p,r}}$), one for each participant, all representing an identical sequence of stimulus trials.
+The KMCCA algorithm was then applied to this set of matrices, returning projections onto a common shared subspace.
+For visualization (Supplementary Fig.
+20A; Fig.
+5A-C), we projected each subject’s data onto the first two canonical axes and then averaged the projections for the repetitions of each unique image to produce a single, stable point per stimulus.
+The parcel-level KMCCA projections (Supplementary Fig.
+20A) confirmed that the organizational principles observed at the hub level were consistent within their constituent parcels.
+In Early Visual Cortex, the V1 parcel showed no clear semantic organization in its first two dimensions, whereas the V4 parcel already exhibited emergent clustering by image content.
+Within the Ventral hub, representative parcels such as VMV2 and PHA2 reproduced the hub-level scene-to-object gradient and showed clear organization by semantic category.
+Similarly, within the LOTC hub, parcels like MT and TPOJ2 showed the same primary separation between biological and non-biological stimuli.
+Within these main clusters, stimuli were further organized by fine-grained semantic categories, mirroring the main hub-level analysis (Fig.
+5C).
+To interpret these KMCCA axes semantically, we projected the image captions onto the shared subspace learned for each hub.
+We first created a vocabulary from the union of all words present in the captions ($s=11,090$ unique words), encoding each caption as a binary word-occurrence vector (a “bag-of-words” representation).
+We then projected this caption matrix onto the first KMCCA component derived for each hub, yielding a loading weight for each word in the vocabulary.
+By ranking these weights, we identified the words that contributed most positively and negatively to each dimension (Fig.
+20B).
+This word-loading analysis confirmed our interpretation of each hub’s organizing principle: in Early Visual Cortex, top-weighted words were not semantically coherent but appeared to share coarse visual features; in the Ventral hub, positive tokens index scene contexts (e.g., ‘kitchen’, ‘bathroom’) while negative tokens name isolated objects (e.g., ‘plate’, ‘frisbee’); and in the LOTC hub, tokens cleanly separated animate from inanimate labels.
+This analysis provides a semantic interpretation for the organization of each cluster, which can be visually confirmed in the corresponding image atlases (Supplementary Figs.
+911).
+Finally, we quantified the impact of the identified dimensions on the overall inter-subject alignment using partial RSA.
+Denoting $\rho(X,Y)$ as the standard Pearson correlation used to compute the alignment between two RDMs, and $Z$ as a control RDM (e.g., one built from the first KMCCA component or from a binary “biological presence” vector), the partial RSA was computed as:
+$\mathrm{pRSA}(X,Y\mid Z)=\frac{\rho(X,Y)\;-\;\rho(X,Z)\,\rho(Y,Z)}{\sqrt{(1-\rho(X,Z)^{2})\,(1-\rho(Y,Z)^{2})}}.$ (3)
+As shown in the main text (Fig.
+5d-f), controlling for the continuous, data-driven KMCCA axis produced only minimal drops in the Early Visual and Ventral hubs, but a substantial reduction in the LOTC hub.
+A complementary analysis confirmed this result: controlling for the discrete categorical variables yielded a drop in alignment of a similar magnitude to that produced by the continuous KMCCA components (Supplementary Fig.
+20C).
+Supplementary Note 8: THINGS-fMRI and BOLD5000 Processing
+THINGS-fMRI dataset
+We processed the THINGS-fMRI dataset 46 using a pipeline analogous to that for NSD.
+This dataset consists of 3T fMRI recordings from three participants performing an oddball detection task on 8,740 unique object images.
+A subset of 100 images was repeated once per session (12 presentations total); all other images appeared only once.
+For our primary computations, we used only the first presentation of each of the 8,740 images.
+Pre-processed, ICA-regressed single-trial estimates were obtained from the authors’ public repository. We extracted voxel responses into the HCP-MMP atlas parcels provided by the THINGS team.
+To control for within-session variability, inter-subject alignment was computed on a session-wise basis.
+For any pair of participants $(p,q)$ and parcels $(r,r^{\prime})$, we computed IS-RSA by correlating the RDMs generated from the images within a given session $s$:
+$\mathcal{A}_{p,q,r,r^{\prime},s}^{\mathrm{IS}}=\mathrm{RSA}\bigl(X_{p,r,s},X_{q,r^{\prime},s}\bigr).$ (4)
+Group-level maps were then obtained by averaging these scores across sessions and all participant pairs.
+Model-brain RSA followed the same session-wise procedure.
+As a control, we confirmed that computing IS-RSA over the full 8,740-trial set (using $8,740\times 8,740$ RDMs) yielded parcel-wise results that were almost perfectly correlated with our session-wise computation (Pearson’s $r=0.995$).
+While the full-set computation produced systematically lower alignment values (linear slope $\approx 0.53$; $R^{2}=0.99$), this correlation validates that our session-wise approach preserves the relative spatial pattern of the findings.
+BOLD5000 dataset
+We processed the BOLD5000 dataset 11 using the same session-wise pipeline as for THINGS-fMRI.
+This dataset consists of 3T recordings from four participants performing a valence rating task on 4,916 unique images drawn from the MS-COCO, ImageNet, and SUN databases.
+We used the pre-processed, single-trial $\beta$-estimates provided by the authors, selecting only the first presentation of each image.
+Inter-subject and model-brain RSA were computed following the same session-wise procedure described above.
+For voxel extraction, we registered the HCP-MMP1.0 atlas from MNI152 standard space to each subject’s native T1-weighted scan and then to their mean functional volume using FSL’s FLIRT and FNIRT.
+Qualitative inspection confirmed consistent parcel definitions across participants.
+All other preprocessing and computational steps paralleled the main NSD pipeline.
+Supplementary Note 9: Implementation details for large-scale RSA comparisons
+To efficiently handle the computational demands of large-scale RSA analyses–driven by dataset size and extensive permutation testing–we developed a GPU-accelerated implementation using PyTorch.
+Specifically, we reformulated representational dissimilarity matrix (RDM) calculations as optimized batch matrix operations.
+For each participant pair $p,q\in\mathcal{P}$ and cortical region $r\in\mathcal{R}$, we constructed stacked matrices containing flattened RDMs for all regions simultaneously, significantly reducing memory overhead and runtime compared to iterative loop-based implementations.
+Given data matrices for matched trials $X_{p,r}[\boldsymbol{i}_{pq},:]$ and $X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]$ (where the number of matched trials is $N_{\text{trials}}$), we first computed the upper-triangular vectors of dissimilarities:
+$\mathbf{x}^{(p)}_{pq,r}=\operatorname{vec}_{u}D\left(X_{p,r}[\boldsymbol{i}_{pq},:]\right),\quad\mathbf{x}^{(q)}_{pq,r^{\prime}}=\operatorname{vec}_{u}D\left(X_{q,r^{\prime}}[\boldsymbol{j}_{pq},:]\right),$ (5)
+where $D(\cdot)$ denotes the dissimilarity function (e.g., 1Pearson r1-$\text${Pearson }r) and $\operatorname{vec}_{u}$ denotes vectorization of the upper triangle.
+Each RDM vector was centered and $l_{2}$-normalized before being stacked into a single matrix containing all ROIs:
+$\mathbf{X}^{(p)}_{pq}=\begin{bmatrix}\tilde{\mathbf{x}}^{(p)}_{pq,1}\\ \tilde{\mathbf{x}}^{(p)}_{pq,2}\\ \vdots\\ \tilde{\mathbf{x}}^{(p)}_{pq,|\mathcal{R}|}\end{bmatrix}\quad\in\mathbb{R}^{|\mathcal{R}|\times\frac{N_{\text{trials}}(N_{\text{trials}}-1)}{2}},$ (6)
+and similarly for $\mathbf{X}^{(q)}_{pq}$.
+The full inter-subject connectivity matrix was then obtained efficiently via a single matrix multiplication:
+$\mathcal{A}^{\mathrm{IS}}_{p,q}=\mathbf{X}^{(p)}_{pq}\left(\mathbf{X}^{(q)}_{pq}\right)^{\top}\in\mathbb{R}^{|\mathcal{R}|\times|\mathcal{R}|}$. (7)
+Analogous formulations were applied for Spearman-based RSA, with rank-transformation applied to RDM vectors prior to normalization and dot products.
+For brain–model alignment, we constructed similar stacked tensors for brain and model RDMs across sessions. The brain tensor was organized as:
+$\mathbf{X}^{(\text{brain})}_{p}\in\mathbb{R}^{|\mathcal{R}|\times S\times\frac{N_{\text{trials}}(N_{\text{trials}}-1)}{2}},$ (8)
+and the model tensor as:
+$\mathbf{X}^{(\text{model})}_{m}\in\mathbb{R}^{L\times S\times\frac{N_{\text{trials}}(N_{\text{trials}}-1)}{2}},$ (9)
+where $S$ is the number of sessions and $L$ is the number of model layers.
+Brain–model RSA alignments across all regions and layers were computed via a batched contraction of the normalized RDM vectors:
+$\mathcal{C}_{p,m}=\langle\mathbf{X}^{(\text{brain})}_{p},\;\mathbf{X}^{(\text{model})}_{m}\rangle$. (10)
+This operation was implemented as an Einstein summation corresponding to the signature $r\,s\,d,l\,s\,d\to r\,l$ (summing over sessions $s$ and dissimilarity pairs $d$).
+For permutation testing, we permuted stimulus labels according to $\sigma\in S_{N_{\text{trials}}}$ prior to RSA computation.
+Rather than recomputing dissimilarity matrices from raw data after each shuffle (an $O(N_{\text{trials}}^{2})$ operation), we pre-calculated the equivalent permutation $\sigma^{\prime}$ over the vectorized upper-triangular indices that matches the induced pairwise index shifts.
+This allowed us to permute the flattened RDMs directly via:
+$\mathbf{X}^{(p)}_{pq}[:,\sigma^{\prime}],$ (11)
+preserving the correct stimulus-pair structure with minimal computational overhead.